### v2

# Riesz Representer Fitting under Bregman Divergence: A Unified Framework for Debiased Machine Learning

Abstract
Estimating the Riesz representer is central to debiased machine learning for causal and structural parameter estimation. We propose generalized Riesz regression, a unified framework that estimates the Riesz representer by fitting a representer model via Bregman divergence minimization. This framework includes the squared loss and the Kullback–Leibler (KL) divergence as special cases: the former recovers Riesz regression, while the latter recovers tailored loss minimization. Under suitable model specifications, the dual problems correspond to covariate balancing, which we call automatic covariate balancing. Moreover, under the same specifications, outcome averages weighted by the estimated Riesz representer satisfy Neyman orthogonality even without estimating the regression function, a property we call automatic Neyman orthogonalization. This property not only reduces the estimation error of Neyman orthogonal scores but also clarifies a key distinction between debiased machine learning and targeted maximum likelihood estimation. Our framework can also be viewed as a generalization of density ratio fitting under Bregman divergences to Riesz representer estimation, and it applies beyond density ratio estimation. We provide convergence analyses for both reproducing kernel Hilbert space (RKHS) and neural network model classes. A Python package for generalized Riesz regression is available at https://github.com/MasaKat0/grr.
1 Introduction
The Riesz representer plays a crucial role in debiased machine learning for a variety of causal and structural parameter estimation problems (Chernozhukov et al., 2022b), such as Average Treatment Effect (ATE) estimation (Imbens & Rubin, 2015), Average Marginal Effect (AME) estimation, Average Policy Effect (APE) estimation, and covariate shift adaptation (Shimodaira, 2000; Uehara et al., 2020; Chernozhukov et al., 2025). The Riesz representer arises from the Riesz representation theorem for the parameter functional, and it also has a close connection to semiparametric efficiency bounds (Newey, 1994). In particular, by using the Riesz representer appropriately, we can obtain semiparametric estimators that are asymptotically linear with the efficient influence function, which is also referred to as the Neyman orthogonal score (Chernozhukov et al., 2018).
Straightforward approaches to approximating the Riesz representer often rely on intermediate steps. For example, in ATE estimation, the Riesz representer can be written in terms of the inverse propensity score. A naive approach is to estimate the propensity score and then construct the Riesz representer by taking its inverse. In covariate shift adaptation, the Riesz representer is given by the density ratio, the ratio of two probability density functions (pdfs). A naive approach is to estimate the two pdfs and then take their ratio. However, it is unclear whether such approaches perform well for the task of Riesz representer estimation because they are not designed to minimize the estimation error of the Riesz representer itself.
To address this issue, end-to-end approaches for Riesz representer estimation have been explored, including Riesz regression (Chernozhukov et al., 2021). In particular, in ATE estimation, entropy balancing weights (Hainmueller, 2012), stable balancing weights (Zubizarreta, 2015), and tailored loss minimization (Zhao, 2019) have been proposed. In covariate shift adaptation, direct density ratio estimation methods have been proposed (Sugiyama et al., 2012).
This study provides a general framework for Riesz representer estimation that accommodates these methods. We formulate Riesz representer estimation as a problem of fitting a Riesz representer model to the true Riesz representer under a Bregman divergence (Bregman, 1967). The Bregman divergence includes various discrepancy measures, such as squared distance and Kullback–Leibler (KL) divergence, as special cases. We measure the discrepancy between a Riesz representer model and the true Riesz representer using a Bregman divergence and train the model by minimizing this divergence, where the discrepancy is interpreted as a loss function. Although the true Riesz representer is unknown, we derive an objective function that does not involve the true Riesz representer and can be approximated using only observations. Therefore, we can train the Riesz representer model within an empirical risk minimization framework.
Notably, with the squared loss, the Bregman divergence minimization problem aligns with Riesz regression Chernozhukov et al. (2021). With the KL divergence loss, the Bregman divergence minimization problem aligns with tailored loss minimization Zhao (2019). We note that Bruns-Smith et al. (2025) shows that stable balancing weights are dual solutions of Riesz regression, while Zhao (2019) shows that entropy balancing weights are dual solutions of tailored loss minimization. Thus, our Bregman divergence minimization formulation unifies these existing methods within a single framework.
In the following sections, we introduce our general setup (Section 2) and then propose our Riesz representer estimation method (Section 3). We refer to Riesz representer fitting under a Bregman divergence as generalized Riesz regression. We may also refer to it as Bregman–Riesz regression, direct bias correction term estimation, generalized tailored loss minimization, or generalized covariate balancing (Remark Remark). However, we adopt the term generalized Riesz regression because the choice of loss function is closely related to the choice of link function. For example, the duality relationship between Riesz regression and stable balancing weights holds when we use a linear link (a linear model) for the Riesz representer. In contrast, the duality relationship between tailored loss minimization and entropy balancing weights holds when we use a logistic link (a logistic model) for the Riesz representer. We make this correspondence explicit in Section 4. We refer to this property as automatic covariate balancing.
For our proposed framework, this paper provides convergence rate analyses, major applications, and several extensions. In Section 6, we show the convergence rate of the Riesz representer model to the true Riesz representer when using reproducing kernel Hilbert space (RKHS) regression and neural networks. In particular, we establish minimax optimality. In Section 7, we provide applications of our framework to ATE, AME, and APE estimation, as well as covariate shift adaptation. In Appendices G and H, we also introduce extensions of our framework that are developed in subsequent works. Kato (2025c) points out that nearest neighbor matching-based density ratio estimation proposed in Lin et al. (2023) is a special case of least-squares importance fitting (LSIF) for density ratio estimation (Kanamori et al., 2009). Moreover, since LSIF can be interpreted as Riesz regression, nearest neighbor matching-based ATE estimation can also be interpreted as ATE estimation via Riesz regression. Kato (2025e) proposes a Riesz representer estimation method based on score matching in diffusion models (Song et al., 2021).
Our framework is primarily built on results from Riesz regression, covariate balancing weights, and density ratio estimation. In particular, Bregman divergences have already been applied to density ratio estimation in Sugiyama et al. (2011b), where the authors also report the duality between empirical risk minimization and covariate balancing weights. Our framework generalizes these results to a broader class of applications and bridges the literature on density ratio estimation with causal inference, where Riesz regression and covariate balancing have, in parallel, studied estimation of the Riesz representer via empirical risk minimization and its dual formulation in terms of covariate balancing weights.
Figure 1: A unified framework for debiased machine learning via Riesz representer estimation and Bregman divergence minimization.
2 Setup
We first describe our general setup, which includes various tasks, such as ATE estimation, as special cases. We denote the pair $W\coloneqq(X,Y)$, where $Y\in{\mathcal{Y}}$ is an outcome and $X\in{\mathcal{X}}$ is a regressor vector. Here, $\mathcal{Y}$$\subseteq{\mathbb{R}}$ and ${\mathcal{X}}\subseteq{\mathbb{R}}^{d}$ are outcome and ($d$-dimensional) regressor spaces, respectively. Let $P_{0}$ be the distribution that generates $W$. We assume that we can observe $n$ i.i.d. copies of $W$, denoted as
${\mathcal{D}}\coloneqq\big\{W_{i}\big\}^{n}_{i=1}=\big\{(X_{i},Y_{i})\big\}^{n}_{i=1}$
We denote the regression function by $\gamma_{0}(x)\coloneqq{\mathbb{E}}_{P_{0}}\big[Y\mid X=x\big]$, where ${\mathbb{E}}_{P_{0}}$ denotes the expectation over $P_{0}$. We drop $P_{0}$ when the dependence is obvious.
Our goal is to estimate a parameter of interest of the form
$\theta_{0}\coloneqq{\mathbb{E}}\big[m(W,\gamma_{0})\big],$
where $m(W,\gamma)$ is a functional that depends on data $W$ and a regression function $\gamma\colon{\mathcal{X}}\to{\mathcal{Y}}$. Note that $m(W,\gamma)$ can receive any function $\gamma\colon{\mathcal{X}}\to{\mathcal{Y}}$, not limited to the “true” $\gamma_{0}$. In the following sections, when we first introduce a function (or functional) that depends on parameters of the data-generating process (DGP), we can replace it with an estimator that has the same mapping. Here, $\theta_{0}$ depends on the functional $m$, and by changing $m$, we can derive various parameters as special cases, such as ATE, AME, and APE.
2.1 Riesz representer
For simplicity, we assume that the expected functional $\gamma\mapsto{\mathbb{E}}\big[m(W,\gamma)\big]$ is linear and continuous in $\gamma$, which implies that there is a constant $C>0$ such that ${\mathbb{E}}\big[m(W,\gamma)\big]^{2}\leq C\mathbb{E}\left[\gamma(X)^{2}\right]$ holds for all $\gamma$ with $\mathbb{E}\left[\gamma(X)^{2}\right]<\infty$. From the Riesz representation theorem, there exists a function $v_{m}$ with $\mathbb{E}\left[v_{m}(X)^{2}\right]<\infty$ such that
$\mathbb{E}\left[m(W,\gamma)\right]=\mathbb{E}\left[v_{m}(X)\gamma(X)\right]$
for all $\gamma$ with $\mathbb{E}\left[\gamma(X)^{2}\right]<\infty$. The function $v_{m}$ is referred to as the Riesz representer in Chernozhukov et al. (2025). In ATE estimation, the Riesz representer has also been referred to as the bias-correction term or the clever covariates (van der Laan, 2006; Schuler & van der Laan, 2024).
Remark.
This formulation follows Chernozhukov et al. (2022b) and can be generalized to non-linear maps $\gamma\mapsto\mathbb{E}\left[m(W,\gamma)\right]$. However, we do not present this generalization because it is not our main focus, and the linear case is sufficient for presenting our results.
2.2 Neyman Orthogonal Scores
Let $\eta_{0}=(\gamma_{0},\alpha_{0})$ be a pair of the nuisance parameters, where $\alpha_{0}$ is the Riesz representer associated with the parameter functional $m$. The Neyman orthogonal score is defined as
$\psi(W;\eta,\theta)\coloneqq m(W,\gamma)+\alpha(X)\big(Y-\gamma(X)\big)-\theta.$
Here, it holds that
$\mathbb{E}\left[\psi(W;\eta_{0},\theta_{0})\right]=0,$
which serves as the estimation equation (or moment condition) for estimating $\theta_{0}$. Note that, by Neyman orthogonality, the Gateaux derivative with respect to $\eta$ at $\eta_{0}$ vanishes as
$\partial_{\eta}\mathbb{E}\left[\psi(W;\eta,\theta_{0})\right]\big|_{\eta=\eta_{0}}=0.$
Thus, orthogonality ensures that first-order errors from estimating $\eta_{0}$ do not affect the asymptotic distribution of the final estimator $\widehat{\theta}$ of $\theta_{0}$, provided cross fitting (or a Donsker condition) and mild convergence rate conditions on $\widehat{\eta}$ hold (Chernozhukov et al., 2018, 2022b).
By replacing the moment condition with its empirical analogue, we obtain an estimator $\widehat{\theta}$ of $\theta_{0}$ as the value satisfying
$\frac{1}{n}\sum^{n}_{i=1}\psi\left(W_{i};\widehat{\eta},\widehat{\theta}\right)=0,$
where $\widehat{\eta}$ denotes estimates of $\eta_{0}$ plugged into the Neyman orthogonal estimating equations.
This estimation approach is traditionally called one-step bias correction, or the estimating equation approach (van der Vaart, 2002; Schuler & van der Laan, 2024). Recent work reformulates it as (automatic) debiased machine learning (Chernozhukov et al., 2018, 2022b, 2022c, 2024). This estimator generalizes the augmented inverse probability weighting (AIPW) estimator in causal inference (Bang & Robins, 2005; Tsiatis, 2007).
2.3 Examples
We provide examples of the problems, along with the corresponding Riesz representers and Neyman orthogonal scores.
ATE estimation.
Let the regressor $X$ be $X\coloneqq(D,Z)$, where $D\in\{1,0\}$ is a treatment indicator, and $Z\in{\mathcal{Z}}$ is a covariate vector, where $\mathcal{Z}$} denotes its support. Following the Neyman–Rubin framework, let $Y(1),Y(0)\in{\mathcal{Y}}$ be the potential outcomes for treated and control units. In ATE estimation, using the observations $\{(X_{i},Y_{i})\}_{i=1}^{n}$, our goal is to estimate the ATE, defined as
$\theta_{0}^{\text{ATE}}\coloneqq\mathbb{E}\big[m^{\text{ATE}}\big(X,\gamma_{0}\big)\big],\quad m^{\text{ATE}}(X,\gamma_{0})\coloneqq\gamma_{0}(1,Z)-\gamma_{0}(0,Z).$
To identify the ATE, we assume standard conditions such as unconfoundedness, positivity, and boundedness of the random variables, that is, $Y(1)$ and $Y(0)$ are independent of $D$ given $Z$, there exists a universal constant $\epsilon\in(0,1/2)$ such that $\epsilon<e_{0}(X)<1-\epsilon$, and $X$, $Y(1)$, and $Y(0)$ are bounded.
In ATE estimation, the Riesz representer is given by
$\alpha_{0}^{\text{ATE}}(X)\coloneqq\frac{D}{e_{0}(Z)}-\frac{1-D}{1-e_{0}(Z)}.$
This term is referred to by various names across different methods. In the classical semiparametric inference literature, it is called the bias-correction term (Schuler & van der Laan, 2024). In TMLE, it is called the clever covariates (van der Laan, 2006). In the debiased machine learning (DML) literature, it is called the Riesz representer (Chernozhukov et al., 2022b). The component $\frac{1}{e_{0}(Z)}$ may also be referred to as balancing weights (Imai & Ratkovic, 2013b; Hainmueller, 2012), the inverse propensity score (Horvitz & Thompson, 1952), or a density ratio (Sugiyama et al., 2012).
The Neyman orthogonal score is given as
$\psi^{\text{ATE}}\left(W;\eta,\theta\right)\coloneqq m^{\text{ATE}}(W,\gamma)+\alpha^{\text{ATE}}(X)\Big(Y-\gamma(X)\Big)-\theta.$
Plugging in estimates of $\eta^{\text{ATE}}_{0}\coloneqq\left(\gamma_{0},\alpha^{\text{ATE}}_{0}\right)$ and solving $\frac{1}{n}\sum_{i=1}^{n}\psi^{\text{ATE}}\left(W_{i};\widehat{\eta},\theta\right)=0$ for $\theta$, we can obtain an ATE estimator as
$).\widehat{\theta}^{\text{ATE}}\coloneqq\frac{1}{n}\sum_{i=1}^{n}\Big(\widehat{\alpha}^{\text{ATE}}(X_{i})\left(Y_{i}-\widehat{\gamma}(X_{i})\right)+m^{\text{ATE}}\left(X_{i},\widehat{\gamma}\right$$\Big)$.
This estimator is called the augmented inverse probability weighting (AIPW) estimator, or the doubly robust estimator.
AME estimation.
Let the regressor $X$ be $X=(D,Z)$ with a (scalar) continuous treatment $D$. We define the AME as
$\theta^{\text{AME}}_{0}\coloneqq\mathbb{E}\Big[\partial_{d}\gamma_{0}(D,Z)\Big].$
Here, the linear functional is given by
$m^{\text{AME}}(W,\gamma)\coloneqq\partial_{d}\gamma(D,Z).$
The Riesz representer that satisfies $\mathbb{E}\left[m^{\text{AME}}(W,\gamma)\right]=\mathbb{E}\left[\alpha^{\text{AME}}_{0}(X)\gamma(X)\right]$ is the (negative) score of the joint density of $X=(D,Z)$ with respect to $d$:
$\alpha^{\text{AME}}_{0}(X)=-\partial_{d}\log f_{0}(D,Z),$
where $f_{0}(X)$ is the joint probability density of $X$.
The Neyman orthogonal score is given as
$\psi^{\text{AME}}(W;\eta,\theta)\coloneqq m^{\text{AME}}(W,\gamma)+\alpha^{\text{AME}}(X)\big(Y-\gamma(X)\big)-\theta.$
Plugging in estimates of $\eta^{\text{AME}}_{0}\coloneqq\left(\gamma_{0},\alpha^{\text{AME}}_{0}\right)$ and solving $\frac{1}{n}\sum_{i=1}^{n}\psi^{\text{AME}}\left(W_{i};\widehat{\eta},\theta\right)=0$ for $\theta$, we can obtain an AME estimator as
$).\widehat{\theta}^{\text{AME}}\coloneqq\frac{1}{n}\sum_{i=1}^{n}\Big(\widehat{\alpha}^{\text{AME}}(X_{i})\left(Y_{i}-\widehat{\gamma}(X_{i})\right)+m^{\text{AME}}\left(X_{i},\widehat{\gamma}\right$$\Big)$.
APE estimation.
We consider the average effect of a counterfactual shift in the distribution of the regressors from a known $P_{-1}$ to another $P_{1}$, when $\gamma_{0}$ is invariant to the distribution of $X$. We refer to this average effect as the APE and define it as
$\theta^{\text{APE}}_{0}\coloneqq\int\gamma_{0}(x){\mathrm{d}}\mu(x),$
where $\mu(x)\coloneqq P_{1}(x)-P_{-1}(x)$. Here, the linear functional is given by
$m^{\text{APE}}(W,\gamma)\coloneqq\int\gamma(x){\mathrm{d}}\mu(x).$
For simplicity, let us assume that the distributions $P_{1}$ and $P_{-1}$ have pdfs, which we denote by $p_{1}(x)$ and $p_{-1}(x)$. We also assume that there exist common supports among the marginal covariate pdf $p_{0}(x)$ of the DGP and the pdfs $p_{1}(x)$ and $p_{-1}(x)$. Then, the Riesz representer is given as
$\alpha^{\text{APE}}_{0}(X)\coloneqq\frac{p_{1}(X)-p_{-1}(X)}{p_{0}(X)}$
The Neyman orthogonal score is
$\psi^{\text{APE}}(W;\eta,\theta)=m^{\text{APE}}(W,\gamma)+\alpha^{\text{APE}}_{0}(X)\big(Y-\gamma(X)\big)-\theta.$
Plugging in estimates of $\eta^{\text{APE}}_{0}\coloneqq\left(\gamma_{0},\alpha^{\text{APE}}_{0}\right)$ and solving $\frac{1}{n}\sum_{i=1}^{n}\psi^{\text{APE}}\left(W_{i};\widehat{\eta},\theta\right)=0$ for $\theta$, we can obtain an APE estimator as
$).\widehat{\theta}^{\text{APE}}\coloneqq\frac{1}{n}\sum_{i=1}^{n}\Big(\widehat{\alpha}^{\text{APE}}(X_{i})\left(Y_{i}-\widehat{\gamma}(X_{i})\right)+m^{\text{APE}}\left(X_{i},\widehat{\gamma}\right$$\Big)$.
Covariate shift adaptation.
Let $X$ denote a source covariate distribution that generates labeled data $(X,Y)$ under $P_{0}$, where the pdf of $X$ is $p_{0}(x)$. Let $\widetilde{X}$ denote a target covariate distribution $P_{X,1}$ whose pdf is $p_{1}(x)$. Let $\{(X_{i},Y_{i})\}_{i\in{\mathcal{I}}_{S}}$ be i.i.d. from $P_{0}$ (source) and $\{\widetilde{X}_{j}\}_{j\in{\mathcal{I}}_{T}}$ be i.i.d. from $P_{X,1}$ (target), independent.
Suppose we train $\gamma_{0}(x)=\mathbb{E}\left[Y\mid X=x\right]$ on a source population with covariates $(X,Y)\sim P_{0}$, but the target parameter averages $\gamma_{0}$ over a shifted covariate distribution $\widetilde{X}\sim P_{X,1}$:
$\theta^{\text{CS}}_{0}\coloneqq\mathbb{E}\big[\gamma_{0}(\widetilde{X})\big]=\mathbb{E}\big[m^{\text{CS}}(\widetilde{X},\gamma_{0})\big],\qquad m^{\text{CS}}(x,\gamma)\coloneqq\gamma(x).$
Let $P_{X,0}$ be the marginal distribution of $X$ under $P_{0}$, and $p_{0}(x)$ and $p_{1}(x)$ be the pdfs. Assume that if $p_{0}(x)>0$, then $p_{1}(x)>0$ holds. Then, the Riesz representer is
$\alpha^{\text{CS}}_{0}(X)=r_{0}(X)=\frac{p_{1}(X)}{p_{0}(X)}.$
The Neyman orthogonal score is given as
$\psi^{\text{CS}}(W;\eta,\theta)=\gamma(X)+\alpha^{\text{CS}}(X)\big(Y-\gamma(X)\big)-\theta.$
Plugging in estimates of $\eta^{\text{CS}}_{0}\coloneqq\left(\gamma_{0},\alpha^{\text{CS}}_{0}\right)$ and solving $\frac{1}{n}\sum_{i=1}^{n}\psi^{\text{CS}}\left(W_{i};\widehat{\eta},\theta\right)=0$ for $\theta$, we can obtain an estimator as
$).\widehat{\theta}^{\text{CS}}\coloneqq\frac{1}{n}\sum_{i=1}^{n}\Big(\widehat{\alpha}^{\text{CS}}(X_{i})\left(Y_{i}-\widehat{\gamma}(X_{i})\right)+m^{\text{CS}}\left(X_{i},\widehat{\gamma}\right$$\Big)$.
Density ratio estimation
We also note that the density ratio itself is important in various tasks such as learning with noisy labels (Liu & Tao, 2016), anomaly detection (Smola et al., 2009; Hido et al., 2008; Abe & Sugiyama, 2019; Nam & Sugiyama, 2015; Kato & Teshima, 2021), two-sample testing (Keziou & Leoni-Aubin, 2005; Sugiyama et al., 2011a), and change point detection (Kawahara & Sugiyama, 2009). Learning from positive and unlabeled data (PU learning) can also be interpreted as an application of density ratio estimation (Kato et al., 2019). Thus, density ratio estimation has been studied as an independent task in machine learning (Sugiyama et al., 2012).
Sugiyama et al. (2008) consider covariate shift adaptation using importance weighting estimated by LSIF, which are equivalent to Riesz regression in density ratio estimation (Kato, 2025d). Chernozhukov et al. (2025) and Kato et al. (2024a) investigate efficient estimation of parameters under a covariate shift from some different perspectives.
Notations and assumptions.
If there are double parentheses $((\cdot))$, we omit one of them. For example, in ATE estimation, since $X=(D,Z)$, we often encounter $f(X)=f((D,Z))$ for some function $f$ of $X$. In such a case, we write $f(X)=f(D,Z)$. Let $\mathbb{E}$} be the expectation over $P_{0}$ if there are no other explanations. We use the subscript 0 to denote parameters under $P_{0}$.
Table 1: Correspondence among Bregman divergence losses, density ratio (DR) estimation methods, and Riesz representer ($R$) estimation in ATE estimation. RR estimation for ATE estimation includes covariate balancing weights. In the table, $C\in{\mathbb{R}}$ denotes a constant that is determined by the problem and the loss function.
$g(\alpha)$ DR estimation RR estimation for ATE estimation
$(\alpha-C)^{2}$ LSIF Riesz regression (Riesz net and Riesz forest)
(Kanamori et al., 2009) (Chernozhukov et al., 2021, 2022a)
Stable balancing weights
(Zubizarreta, 2015)
Nearest neighbor matching
(Lin et al., 2023)
Causal tree/ causal forest
(Wager & Athey, 2018)
Dual solution with linear links
Kernel mean matching Stable balancing weights
(Gretton et al., 2009) (Zubizarreta, 2015; Bruns-Smith et al., 2025)
Covariate balancing by SVM
(Tarr & Imai, 2025)
$\big(|\alpha|-C\big)\log\big(|\alpha|-C\big)-|\alpha|$ UKL divergence minimization Tailored loss minimization ($\alpha=\beta=-1$)
(Nguyen et al., 2010) (Zhao, 2019)
Dual solution with logistic links
KLIEP Entropy balancing weights
(Sugiyama et al., 2008) (Hainmueller, 2012)
$(|\alpha|-C)\log\big(|\alpha|-C\big)$ BKL divergence minimization MLE of the propensity score
$\qquad-(|\alpha|+C)\log(|\alpha|+C)$ (Qin, 1998) (Standard approach)
Tailored loss minimization ($\alpha=\beta=0$)
(Zhao, 2019)
$C\log\left(1-|\alpha|\right)$ PU learning
$C|\alpha|\left(\log\left(|\alpha|\right)-\log\left(1-|\alpha|\right)\right)$ (du Plessis et al., 2015)
     for $\alpha\in(0,1)$ (Kato & Teshima, 2021)
3 Generalized Riesz Regression
Various methods for estimating the Riesz representer have been proposed. In ATE estimation, a standard approach is to estimate the propensity score $e(x)=P(D=1\mid X=x)$ via maximum likelihood estimation (MLE) in a logistic model and then plug the estimate into the Riesz representer $\alpha_{0}^{\text{ATE}}(X)$. Beyond MLE, covariate balancing approaches have been proposed, where we estimate the propensity score or balancing weights, which implicitly or explicitly correspond to the inverse propensity score, by matching covariate moments. Chernozhukov et al. (2021) proposes Riesz regression as a general method for Riesz representer estimation. In density ratio estimation, related approaches include moment matching (Huang et al., 2007; Gretton et al., 2009), probabilistic classification (Qin, 1998; Cheng & Chu, 2004), density matching (Nguyen et al., 2010), and density ratio fitting (Kanamori et al., 2009). For details on related work, see Section A.
In this study, we generalize existing methods through the lens of Bregman divergence minimization. The Bregman divergence is a general discrepancy measure that includes the squared loss and the KL divergence as special cases. In this section, we propose Riesz representer fitting under a Bregman divergence, which we also refer to as generalized Riesz regression. The term generalized Riesz regression reflects the fact that the choice of loss function is closely connected to the choice of link function from the viewpoint of covariate balancing. We explain this viewpoint in Section 4 and refer to it as automatic covariate balancing. This section provides a general formulation, and we introduce applications of generalized Riesz regression in Section 7.
Remark.
We also refer to our method as direct bias correction term estimation, Bregman Riesz regression, or generalized tailored loss minimization. In an earlier draft, we used the term direct bias correction term estimation because the Riesz representer is almost equivalent to the bias correction term in one step bias correction. Bregman Riesz regression highlights that the method combines the Bregman divergence with Riesz regression. Generalized tailored loss minimization emphasizes that our generalized Riesz regression extends tailored loss minimization (Zhao, 2019) and covers a broader class of methods, including Riesz regression. As discussed in Section 5, standard covariate balancing methods implicitly assume a constant (homogeneous) ATE across $x$, whereas the covariate balancing property under generalized Riesz regression allows for heterogeneity. From this viewpoint, we call the method generalized covariate balancing.
3.1 Bregman Divergence
This study fits a Riesz representer model $\alpha(X)$ to the true Riesz representer $\alpha_{0}(X)$ under a Bregman divergence. Let $g\colon{\mathbb{R}}\to{\mathbb{R}}$ be a differentiable and strictly convex function. As discussed in Section 4, this function $g$ corresponds to the objective function in covariate balancing.
Given $x\in{\mathcal{X}}$, the Bregman divergence between the scalar values $\alpha_{0}(x)$ and $\alpha(x)$ is defined as
$\text{BD}^{\dagger}_{g}\big(\alpha_{0}(x)\mid\alpha(x)\big)\coloneqq g(\alpha_{0}(x))-g(\alpha(x))-\partial g(\alpha(x))\big(\alpha_{0}(x)-\alpha(x)\big),$
where $\partial g$ denotes the derivative of $g$. We then define the average Bregman divergence as
$\text{BD}^{\dagger}_{g}\big(\alpha_{0}\mid\alpha\big)\coloneqq\mathbb{E}\Big[g(\alpha_{0}(X))-g(\alpha(X))-\partial g(\alpha(X))\big(\alpha_{0}(X)-\alpha(X)\big)\Big].$
We define the population target as
$\alpha^{*}=\operatorname*{arg\,min}_{\alpha\in{\mathcal{A}}}\text{BD}^{\dagger}_{g}\big(\alpha_{0}\mid\alpha\big),$
where $\mathcal{A}$} is a set of candidates for $\alpha_{0}$. If $\alpha_{0}\in{\mathcal{A}}$, then $\alpha^{*}=\alpha_{0}$ holds.
Although $\alpha_{0}$ is unknown, we can define an equivalent optimization problem that does not involve $\alpha_{0}$:
$\alpha^{*}=\operatorname*{arg\,min}_{\alpha\in{\mathcal{A}}}\text{BD}_{g}\big(\alpha\big),$
where
$\text{BD}_{g}\big(\alpha\big)\coloneqq\mathbb{E}\Big[-g(\alpha(X))+\partial g(\alpha(X))\alpha(X)-m\big(W,(\partial g)\circ\alpha\big)\Big].$
Here, we use the linearity of $m$ and the Riesz representation theorem, which imply that
${\mathbb{E}}\Big[\partial g(\alpha(X))\alpha_{0}(X)\Big]={\mathbb{E}}\Big[m\big(W,(\partial g)\circ\alpha\big)\Big].$
We estimate the Riesz representer $\alpha_{0}$ by minimizing an empirical Bregman divergence:
$\widehat{\alpha}\coloneqq\operatorname*{arg\,min}_{\alpha\in{\mathcal{A}}}\widehat{\text{BD}}_{g}\big(\alpha\big)+\lambda J(\alpha),$ (1)
where $J(\alpha)$ is a regularization function, and
$\widehat{\text{BD}}_{g}(\alpha)\coloneqq\frac{1}{n}\sum^{n}_{i=1}\Big(-g(\alpha(X_{i}))+\partial g(\alpha(X_{i}))\alpha(X_{i})-m\big(W_{i},(\partial g)\circ\alpha\big)\Big).$
The choice of the regularization function is important because Riesz representer estimation is known to exhibit a characteristic overfitting phenomenon, often described as train-loss hacking or the density chasm. For details, see Section 8.
3.2 Special Cases of the Bregman Divergence
By choosing different $g$, we obtain various objectives for Riesz representer estimation, including Riesz regression. Specifically, we obtain the following divergences (loss functions) as special cases of the Bregman divergence:
  • Squared distance (squared loss): $g^{\text{SQ}}(\alpha)\coloneqq(\alpha-C)^{2}$ for some constant $C\in{\mathbb{R}}$.
  • Unnormalized KL (UKL) divergence: $g^{\text{UKL}}(\alpha)\coloneqq\big(|\alpha|-C\big)\log\big(|\alpha|-C\big)-|\alpha|$ for some constant $C<\inf_{x}|\alpha(x)|$.
  • Binary KL (BKL) divergence: $g^{\text{BKL}}(\alpha)\coloneqq(|\alpha|-C)\log\big(|\alpha|-C\big)-(|\alpha|+C)\log(|\alpha|+C)$ for some constant $C<\inf_{x}|\alpha(x)|$.
  • Basu’s power (BP) divergence (BP-Riesz): $g^{\text{BP}}(\alpha)\coloneqq\frac{\big(|\alpha|-C\big)^{1+\delta}-\big(|\alpha|-C\big)}{\delta}-\big(|\alpha|-C\big)$ for some $\delta\in(0,\infty)$ and $C<\inf_{x}|\alpha(x)|$.
  • PU learning loss: $g^{\text{PU}}(\alpha)\coloneqq\widetilde{C}\log\left(1-|\alpha|\right)+\widetilde{C}|\alpha|\left(\log\left(|\alpha|\right)-\log\left(1-|\alpha|\right)\right)$ for some $\widetilde{C}\in{\mathbb{R}}$, where $\alpha$ takes values in $(0,1)$.
See also Table 1 for a summary.
We refer to our method as SQ-Riesz when using the squared loss, UKL-Riesz when using the UKL divergence, BKL-Riesz when using the BKL divergence, BP-Riesz when using the BP divergence, and PU-Riesz when using the PU learning loss. We explain these special cases in detail in the following subsections.
3.3 SQ-Riesz Regression
Let $C\in{\mathbb{R}}$ be a constant. We consider the following convex function:
$g^{\text{SQ}}(\alpha)=(\alpha-C)^{2}.$
This choice of convex function is motivated by the squared loss. The choice of $C$ depends on the researcher. We propose choosing $C$ so that the automatic covariate balancing property holds, see Section 4. The derivative of $g^{\text{SQ}}(\alpha)$ with respect to $\alpha$ is given as
$\partial g^{\text{SQ}}(\alpha)$ $=2(\alpha-C).$
Under this choice of $g$, the Bregman divergence objective is given as
$\text{BD}_{g}\big(\alpha\big)\coloneqq\mathbb{E}\Big[\alpha(X)^{2}-2m\big(W,\big(\alpha(\cdot)-C\big)\big)\Big].$
Then, the estimation problem can be written as
$\widehat{\alpha}\coloneqq\operatorname*{arg\,min}_{\alpha\in{\mathcal{A}}}\widehat{\text{BD}}_{g^{\text{SQ}}}\big(\alpha\big)+\lambda J(\alpha),$ (2)
where
$\widehat{\text{BD}}_{g^{\text{SQ}}}\big(\alpha\big)\coloneqq\frac{1}{n}\sum^{n}_{i=1}\left(\alpha(X_{i})^{2}-2m\big(W_{i},\big(\alpha(\cdot)\big)\big)\right).$
Here, for simplicity, we drop constant terms that are irrelevant for the optimization and use the linearity of $m$ for $2(\alpha(\cdot)-C)$ [^1] The original empirical Bregman divergence objective using $g(\alpha)=(\alpha-C)^{2}$ in (1) is given as $\widehat{\text{BD}}_{g^{\text{SQ}}}\big(\alpha\big)$ $={\mathbb{E}}\left[-(\alpha(X)-C)^{2}+2(\alpha(X)-C)\alpha(X)-2m\big(W,(\alpha(\cdot)-C)\big)\right]$ $={\mathbb{E}}\left[\alpha(X)^{2}-C^{2}-2m\big(W,\big(\alpha(\cdot)-C\big)\big)\right].$ . This estimation method corresponds to Riesz regression in debiased machine learning (Chernozhukov et al., 2021) and least-squares importance fitting (LSIF) in density ratio estimation (Kanamori et al., 2009). Moreover, if we define $\mathcal{A}$} appropriately, we can recover nearest neighbor matching, as pointed out in Kato (2025c), which extends the argument in Lin et al. (2023).
3.4 UKL-Riesz Regression
Next, we consider a KL-divergence-motivated convex function. Let $C<\inf_{x}|\alpha(x)|$ be a constant. We define
$g^{\text{UKL}}(\alpha)=(|\alpha|-C)\log\left(|\alpha|-C\right)-|\alpha|.$
The choice of $C$ depends on the researcher. We propose choosing $C$ so that the automatic covariate balancing property holds, see Section 4. The derivative of $g^{\text{UKL}}(\alpha)$ with respect to $\alpha$ is given as
$\partial g^{\text{UKL}}(\alpha)$ $=\operatorname{sign}(\alpha)\log\left(|\alpha|-C\right).$
Under this choice of $g$, the Bregman divergence objective is given as follows [^2] This Bregman divergence objective is derived as follows: $\widehat{\text{BD}}_{g^{\text{UKL}}}\big(\alpha\big)$ $={\mathbb{E}}\Big[-\left(|\alpha(X)|-C\right))\log\left(|\alpha(X)|-C\right)+|\alpha(X)|+\operatorname{sign}\left(\alpha(X)\right)\alpha(X)\log\left(|\alpha(X)|-C\right)$ $\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ -m\Big(W,\operatorname{sign}\big(\alpha(\cdot)\big)\log\left(|\alpha(\cdot)|-C\right)\Big)\Big].$ :
$\text{BD}_{g^{\text{UKL}}}\big(\alpha\big)\coloneqq\mathbb{E}\Big[C\log\left(|\alpha(X)|-C\right)+|\alpha(X)|-m\Big(W,\operatorname{sign}\big(\alpha(\cdot)\big)\log\left(|\alpha(\cdot)|-C\right)\Big)\Big].$
We estimate $\alpha_{0}$ by minimizing the empirical objective:
$\widehat{\alpha}\coloneqq\operatorname*{arg\,min}_{\alpha\in{\mathcal{A}}}\widehat{\text{BD}}_{g^{\text{UKL}}}\big(\alpha\big)+\lambda J(\alpha),$
where
$\widehat{\text{BD}}_{g^{\text{UKL}}}\big(\alpha\big)$ $=\frac{1}{n}\sum^{n}_{i=1}\left(C\log\left(|\alpha(X_{i})|-C\right)+|\alpha(X_{i})|-m\Big(W_{i},\operatorname{sign}\big(\alpha(\cdot)\big)\log\left(|\alpha(\cdot)|-C\right)\Big)\right).$
In the next subsection, we also introduce the BKL divergence as a KL-divergence-motivated divergence, but the UKL divergence more closely corresponds to the standard KL divergence. The equivalent formulation is known as KLIEP in density ratio estimation. Note that KLIEP is a constrained formulation that is equivalent to UKL divergence minimization, and this equivalence is also known as Silverman’s trick (Silverman, 1978; Kato et al., 2023). This constrained formulation can also be interpreted as a dual formulation. In ATE estimation, tailored loss minimization corresponds to UKL minimization, whose dual yields entropy balancing weights (Hainmueller, 2012).
3.5 BKL-Riesz Regression
We introduce the BKL divergence and BKL-Riesz, which are motivated by the KL divergence and logistic regression. Let $C<\inf_{x}|\alpha(x)|$ be a constant. We define
$g^{\text{BKL}}(\alpha)\coloneqq(|\alpha|-C)\log\big(|\alpha|-C\big)-(|\alpha|+C)\log(|\alpha|+C).$
The choice of $C$ depends on the researcher. We propose choosing $C$ so that the automatic covariate balancing property holds, see Section 4.
Under this choice of $g$, the Bregman divergence objective is given as follows [^3] This Bregman divergence objective is derived as follows: $\text{BD}_{g^{\text{BKL}}}\big(\alpha\big)$ $={\mathbb{E}}\Big[-(|\alpha(X)|-C)\log\big(|\alpha(X)|-C\big)-(|\alpha(X)|+C)\log(|\alpha(X)|+C)+\operatorname{sign}\big(\alpha(X)\big)\alpha(X)\log\left(\frac{|\alpha(X)|-C}{|\alpha(\cdot)|+C}\right)$ $\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ -m\left(W,\operatorname{sign}\big(\alpha(\cdot)\big)\log\left(\frac{|\alpha(\cdot)|-C}{|\alpha(\cdot)|+C}\right)\right)\Big].$ :
$\text{BD}_{g^{\text{BKL}}}\big(\alpha\big)\coloneqq\mathbb{E}\left[C\log\left(\frac{|\alpha(X)|-C}{|\alpha(X)|+C}\right)-m\left(W,\operatorname{sign}\big(\alpha(\cdot)\big)\log\left(\frac{|\alpha(\cdot)|-C}{|\alpha(\cdot)|+C}\right)\right)\right].$
We estimate $\alpha_{0}$ by minimizing the empirical objective:
$\widehat{\alpha}\coloneqq\operatorname*{arg\,min}_{\alpha\in{\mathcal{A}}}\widehat{\text{BD}}_{g^{\text{BKL}}}\big(\alpha\big)+\lambda J(\alpha),$
where
$\widehat{\text{BD}}_{g^{\text{BKL}}}\big(\alpha\big)$ $=\frac{1}{n}\sum^{n}_{i=1}\left(C\log\left(\frac{|\alpha(X_{i})|-C}{|\alpha(X_{i})|+C}\right)-m\left(W_{i},\operatorname{sign}\big(\alpha(\cdot)\big)\log\left(\frac{|\alpha(\cdot)|-C}{|\alpha(\cdot)|+C}\right)\right)\right).$
In ATE estimation, this formulation corresponds to MLE for a logistic model of the propensity score. In density ratio estimation, this formulation corresponds to a logistic regression approach, where we classify two datasets using a logistic model and then take the ratio to obtain a density ratio estimator. For details, see Section 7.
3.6 BP-Riesz Regression
Basu’s power (BP) divergence bridges the squared loss and KL divergence (Basu et al., 1998). Let $C<\inf_{x}|\alpha(x)|$ be a constant. Based on the BP divergence, we introduce the following function:
$g^{\text{BP}}(\alpha)\coloneqq\frac{\big(|\alpha|-C\big)^{1+\delta}-\big(|\alpha|-C\big)}{\delta}-|\alpha|.$
The derivative is given as
$\partial g^{\text{BP}}(\alpha)=\left(1+\frac{1}{\delta}\right)\operatorname{sign}(\alpha)\Big(\big(|\alpha|-C\big)^{\delta}-1\Big).$
Using this function in the Bregman divergence yields a BP-motivated loss and the corresponding objective for BP-Riesz regression. The choice of $C$ depends on the researcher. We propose choosing $C$ so that the automatic covariate balancing property holds, see Section 4.
Under this choice of $g$, the Bregman divergence objective is given as follows [^4] This Bregman divergence objective is derived from $\text{BD}_{g^{\text{BP}}}\big(\alpha\big)$ $\coloneqq{\mathbb{E}}\Bigg[-\frac{\big(|\alpha(X)|-C\big)^{1+\delta}-\big(|\alpha(X)|-C\big)}{\delta}+|\alpha(X)|+(1+1/\delta)|\alpha(X)|\Big(\big(|\alpha(X)|-C\big)^{\delta}-1\Big)$ $\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ -m\left(W,(1+1/\delta)\operatorname{sign}(\alpha)\Big(\big(|\alpha(\cdot)|-C\big)^{\delta}-1\Big)\right)\Bigg].$ :
$\text{BD}_{g^{\text{BP}}}\big(\alpha\big)\coloneqq{\mathbb{E}}\Bigg[\frac{C\big(\big(|\alpha(X)|-C\big)^{\delta}-1\big)}{\delta}+C|\alpha(X)|\big(|\alpha(X)|-C\big)^{\delta}$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ -m\left(W,\left(1+\frac{1}{\delta}\right)\operatorname{sign}(\alpha(\cdot))\Big(\big(|\alpha(\cdot)|-C\big)^{\delta}-1\Big)\right)\Bigg].$
We estimate $\alpha_{0}$ by minimizing the empirical objective:
$\widehat{\alpha}\coloneqq\operatorname*{arg\,min}_{\alpha\in{\mathcal{A}}}\widehat{\text{BD}}_{g^{\text{BP}}}\big(\alpha\big)+\lambda J(\alpha),$
where
$\widehat{\text{BD}}_{g^{\text{BP}}}\big(\alpha\big)\coloneqq\frac{1}{n}\sum^{n}_{i=1}\Bigg(\frac{C\big(\big(|\alpha(X_{i})|-C\big)^{\delta}-1\big)}{\delta}+C\big(|\alpha(X_{i})|-C\big)^{\delta}+\big(|\alpha(X_{i})|-C\big)^{1+\delta}$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ -m\left(W_{i},\left(1+\frac{1}{\delta}\right)\operatorname{sign}(\alpha(\cdot))\Big(\big(|\alpha(\cdot)|-C\big)^{\delta}-1\Big)\right)\Bigg).$
Basu’s power divergence bridges the squared loss and the (U)KL divergence. When $\delta=1$, BP-Riesz regression reduces to SQ-Riesz regression, while when $\delta\to 0$, BP-Riesz regression reduces to UKL-Riesz regression. This follows because
$\lim_{\delta\to 0}\frac{\big(|\alpha|-C\big)^{\delta}-1}{\delta}=\log\big(|\alpha|-C\big).$
BP-Riesz regression plays an important role in robust estimation of the Riesz representer. UKL-Riesz regression implicitly assumes exponential or sigmoid models for the Riesz representer. If the model is misspecified, the estimation accuracy can deteriorate. As Sugiyama et al. (2012) notes, SQ-Riesz regression is more robust to outliers, while UKL-Riesz regression can perform well under correct specification. BP-Riesz regression provides an intermediate objective between these two extremes. In addition, BP-Riesz regression is useful for understanding the automatic covariate balancing property.
3.7 PU-Riesz Regression
We introduce PU learning loss and PU-Riesz, which are motivated by PU learning. Let $C<\inf_{x}|\alpha(x)|$ be some constant. We define $g^{\text{PU}}$ as
$g^{\text{PU}}(\alpha)\coloneqq\widetilde{C}\log\left(1-|\alpha|\right)+\widetilde{C}|\alpha|\Big(\log\left(|\alpha|\right)-\log\left(1-|\alpha|\right)\Big)$
for some $\widetilde{C}\in{\mathbb{R}}$, and we restrict $\alpha$ to take values in $(0,1)$. The choice of $\widetilde{C}$ depends on the researcher. It corresponds to the class prior in PU learning and plays a role that differs from the parameter $C$ in the other loss functions. The derivative of $g^{\text{PU}}(\alpha)$ with respect to $\alpha$ is given as
$\partial g^{\text{PU}}(\alpha)$ $=-\frac{\widetilde{C}\operatorname{sign}(\alpha)}{1-|\alpha|}+\widetilde{C}\operatorname{sign}(\alpha)\left(\log\left(|\alpha|\right)-\log\left(1-|\alpha|\right)+\frac{1}{1-|\alpha|}\right)$
$=\widetilde{C}\operatorname{sign}(\alpha)\Big(\log\left(|\alpha|\right)-\log\left(1-|\alpha|\right)\Big).$
Under this choice of $g$, the Bregman divergence objective is given as follows:
$\text{BD}_{g^{\text{PU}}}\big(\alpha\big)\coloneqq\mathbb{E}\left[-\widetilde{C}\log\left(1-|\alpha(X)|\right)-m\left(W,\widetilde{C}\operatorname{sign}(\alpha)\Big(\log\left(|\alpha(\cdot)|\right)-\log\left(1-|\alpha(\cdot)|\right)\Big)\right)\right].$
Then, we estimate $\alpha_{0}$ by minimizing the empirical objective:
$\widehat{\alpha}\coloneqq\operatorname*{arg\,min}_{\alpha\in{\mathcal{A}}}\widehat{\text{BD}}_{g^{\text{PU}}}\big(\alpha\big)+\lambda J(\alpha),$
where
$\widehat{\text{BD}}_{g^{\text{PU}}}\big(\alpha\big)$ $=\frac{1}{n}\sum^{n}_{i=1}\left(-\widetilde{C}\log\left(1-|\alpha(X_{i})|\right)-m\left(W_{i},\widetilde{C}\operatorname{sign}(\alpha)\Big(\log\left(|\alpha(\cdot)|\right)-\log\left(1-|\alpha(\cdot)|\right)\Big)\right)\right).$
PU learning is a classical problem. For example, Lancaster & Imbens (1996) studies this problem under a stratified sampling scheme (Wooldridge, 2001). du Plessis et al. (2015) rediscovers this formulation and calls it unbiased PU learning. Kato et al. (2019) points out the relationship between PU learning and density ratio estimation, and Kato & Teshima (2021) shows that PU learning is a special case of density ratio model fitting under a Bregman divergence. Our results further generalize these results. Note that PU learning in these settings and our setting is called case-control PU learning. There is also another formulation called censoring PU learning (Elkan & Noto, 2008). Kato et al. (2025) considers ATE estimation in a PU learning setup and applies our method in their study.
4 Automatic Covariate Balancing
Under specific choices of the Riesz representer model and the Bregman divergence, we can guarantee an automatic covariate balancing property. The key tool is the duality between Bregman divergence minimization and covariate balancing methods. This automatic covariate balancing property yields the automatic Neyman orthogonalization property discussed in Section 5, which in turn automatically guarantees Neyman orthogonality for IPW-type estimators. We note that the covariate balancing property does not hold if we use cross fitting.
4.1 Generalized Linear Models
Throughout this section, we consider a Riesz representer model of the form
$\alpha(X)=\zeta^{-1}\Big(X,{\bm{\phi}}(X)^{\top}{\bm{\beta}}\Big),$
where $\zeta^{-1}$ is the inverse of a link function and ${\bm{\phi}}\colon{\mathcal{X}}\to{\mathbb{R}}^{p}$ is a basis function. A link function $\zeta$ connects the Riesz representer $\alpha$ to a linear, or linear in parameters, index,
${\bm{\phi}}(X)^{\top}{\bm{\beta}}.$
This introduction of a linear index is motivated by generalized linear models. While standard generalized linear models assume a link of the form $\alpha(X)=\zeta^{-1}\big({\bm{\phi}}(X)^{\top}{\bm{\beta}}\big)$, we allow the transformation to depend on $X$ as $\alpha(X)=\zeta^{-1}\big(X,{\bm{\phi}}(X)^{\top}{\bm{\beta}}\big)$, as described below. Given $X$, we define $\zeta$ so that
$\zeta\big(X,\alpha(X)\big)={\bm{\phi}}(X)^{\top}{\bm{\beta}}.$
Examples.
For example, we can approximate the Riesz representer by
$\alpha_{{\bm{\beta}}}(X)\coloneqq{\bm{\phi}}(X)^{\top}{\bm{\beta}},$
which corresponds to using a linear link function for $\zeta^{-1}$. This linear specification can be applied in many settings, including ATE estimation and density ratio estimation.
We can improve estimation accuracy by incorporating additional modeling assumptions. For example, in ATE estimation, we can approximate the propensity score $e_{0}(Z)=P(D=1\mid Z)$ by a logistic model,
$e_{{\bm{\beta}}}(Z)\coloneqq\frac{1}{1+\exp\Big(-{\bm{\phi}}(Z)^{\top}{\bm{\beta}}\Big)},$
where ${\bm{\phi}}\colon{\mathcal{Z}}\to{\mathbb{R}}^{p}$ is a basis function, and $\bm{\beta}$} is the corresponding parameter. Note that in this case, we consider a basis function that receives $Z$ not $X$, or we can interpret that ${\bm{\phi}}(D,Z)$ only depends on $Z$ and is independent of $D$. Plugging this propensity score model into the Riesz representer for ATE, we can approximate the Riesz representer by
$\alpha^{\text{ATE}}_{{\bm{\beta}}}(X)\coloneqq\frac{D}{e_{{\bm{\beta}}}(Z)}+\frac{1-D}{1-e_{{\bm{\beta}}}(Z)}.$
In such cases, we define $\zeta^{-1}$ so that
$\alpha_{{\bm{\beta}}}(X)$ $=\zeta^{-1}\left(X,{\bm{\phi}}(X)^{\top}{\bm{\beta}}\right)$
$=\frac{D}{e_{{\bm{\beta}}}(Z)}-\frac{1-D}{1-e_{{\bm{\beta}}}(Z)}$
$=D\Big(1+\exp\Big(-{\bm{\phi}}(Z)^{\top}{\bm{\beta}}\Big)\Big)-(1-D)\Big(1+\exp\Big({\bm{\phi}}(Z)^{\top}{\bm{\beta}}\Big)\Big).$ (3)
Note that we can also model the Riesz representer as
$\alpha_{{\bm{\beta}}}(X)$ $=\zeta^{-1}\left(X,{\bm{\phi}}(X)^{\top}{\bm{\beta}}\right)$
$=\frac{D}{e_{{\bm{\beta}}}(Z)}-\frac{1-D}{1-e_{{\bm{\beta}}}(Z)}$
$=D\Big(1+\exp\Big(-{\bm{\phi}}(Z)^{\top}{\bm{\beta}}\Big)\Big)-(1-D)\Big(1+\exp\Big({\bm{\phi}}(Z)^{\top}{\bm{\beta}}\Big)\Big),$ (4)
with including $D$ in the basis function. The choice of basis functions depend on the heterogeneity of $\gamma_{0}(X)$. As we discuss in Section 5, if $\gamma_{0}(x)$ is constant for all $x$, (4.1) may be more appropriate. In contrast, if $\gamma_{0}(x)$ varies across $x$, (4.1) may be more appropriate.
Similarly, in covariate shift adaptation, we can model the density ratio as
$\alpha^{\text{CS}}_{{\bm{\beta}}}(X)=\exp\Big(-{\bm{\phi}}(X)^{\top}{\bm{\beta}}\Big).$
Here, the link function is given by
$\zeta\left(X,\alpha^{\text{CS}}_{{\bm{\beta}}}(X)\right)=-{\bm{\phi}}(X)^{\top}{\bm{\beta}}.$
4.2 Automatic Covariate Balancing
This section presents a covariate balancing property for the basis function ${\bm{\phi}}(X)$ when the Riesz representer model has the form $\alpha(X)=\zeta^{-1}\big(X,{\bm{\phi}}(X)^{\top}{\bm{\beta}}\big)$. Related properties have been reported in Ben-Michael et al. (2021). More specifically, for Riesz regression, Bruns-Smith & Feller (2022) and Bruns-Smith et al. (2025) discuss the duality between Riesz regression and stable balancing weights (Zubizarreta, 2015). Zhao (2019) also reports the duality between tailored loss minimization and entropy balancing weights. Tarr & Imai (2025) investigates such a duality in covariate balancing method using the support vector machine (SVM). Thus, the duality has garnered attention in covariate balancing (Tan, 2019; Orihara et al., 2025). In density ratio estimation, such a duality is also well known, as discussed in Sugiyama et al. (2007) and Sugiyama et al. (2011b). In the density ratio estimation literature, the resulting balancing property is typically interpreted as moment matching. This study generalizes these results and provides new findings on the balancing property. For the proof, see Appendix B.
Theorem 4.1 (Automatic covariate balancing).
Assume that there exists a function $\widetilde{g}\colon{\mathcal{X}}\times{\mathbb{R}}\to{\mathbb{R}}$ such that
$\partial g\big(\alpha_{{\bm{\beta}}}(X_{i})\big)=\sum^{p}_{j=1}\beta_{j}\widetilde{g}\left(X_{i},\phi_{j}(X_{i})\right);$
that is, $\partial g\big(\alpha_{{\bm{\beta}}}(X_{i})\big)$ is linear in $\widetilde{g}\left(X_{i},\phi_{1}(X_{i})\right),\dots,\widetilde{g}\left(X_{i},\phi_{p}(X_{i})\right)$. Consider a Riesz representer estimator $\widehat{\alpha}=\alpha_{\widehat{{\bm{\beta}}}}$ trained by generalized Riesz regression (empirical risk minimization) with $\ell_{a}$-penalty as
$\widehat{{\bm{\beta}}}\coloneqq\operatorname*{arg\,min}_{{\bm{\beta}}\in{\mathbb{R}}^{p}}$ $\ \ \ \left\{\frac{1}{n}\sum^{n}_{i=1}\Big(-g(\alpha_{{\bm{\beta}}}(X_{i}))+\alpha_{{\bm{\beta}}}(X_{i})\partial g\big(\alpha_{{\bm{\beta}}}(X_{i})\big)-m\big(W_{i},(\partial g)\circ\alpha_{{\bm{\beta}}}\big)\Big)+\frac{1}{a}\lambda\|{\bm{\beta}}\|^{a}_{a}\right\}.$
In this case $\widehat{\alpha}$ satisfies
$\left|\frac{1}{n}\sum^{n}_{i=1}\Big(\widehat{\alpha}(X_{i})\widetilde{g}\big(X_{i},\phi_{j}(X_{i})\big)-m(W_{i},\widetilde{g}\left(X_{i},\phi_{j}(X_{i})\right))\Big)\right|\leq\lambda\left|\widehat{\beta}_{j}\right|^{a-1}.$
Generalized Riesz regression returns a Riesz representer model that minimizes the given loss among models satisfying the above inequality.
In addition, if $\partial g\big(\alpha_{{\bm{\beta}}}(X_{i})\big)=\sum^{p}_{j=1}\beta_{j}\phi_{j}(X_{i})$ holds, the following result holds.
Corollary 4.2.
Assume the conditions in Theorem 4.1. Then, if $\partial g\big(\alpha_{{\bm{\beta}}}(X_{i})\big)=\sum^{p}_{j=1}\beta_{j}\phi_{j}(X_{i})$ holds, then the inequality can be written as
$\left|\frac{1}{n}\sum^{n}_{i=1}\Big(\widehat{\alpha}(X_{i})\phi_{j}(X_{i})-m(W_{i},\phi_{j})\Big)\right|\leq\lambda\left|\widehat{\beta}_{j}\right|^{a-1}.$
These results correspond to covariate balancing in ATE estimation and moment matching in density ratio estimation.
Examples.
For simplicity, we consider $\ell_{1}$-penalty in generalized Riesz regression. Then, in ATE estimation, if we use a linear link and SQ-Riesz regression, we have
$\left|\frac{1}{n}\sum^{n}_{i=1}\widehat{\alpha}(X_{i})\phi_{j}(D_{i},Z_{i})-\frac{1}{n}\sum^{n}_{i=1}\Big(\phi_{j}(1,Z_{i})-\phi_{j}(0,Z_{i})\Big)\right|<\lambda\quad(\text{for}\ j=1,2,\dots,p).$
If we use a logistic link and UKL-Riesz regression, we have
$\left|\frac{1}{n}\sum^{n}_{i=1}\Big(\mathbbm{1}[D_{i}=1]\widehat{\alpha}(X_{i})\phi_{j}(Z_{i})-\mathbbm{1}[D_{i}=0]\widehat{\alpha}(X_{i})\phi_{j}(Z_{i})\Big)\right|\leq\lambda\quad(\text{for}\ j=1,2,\dots,p).$
In density ratio estimation, if we use a linear link and SQ-Riesz regression, we have
$\left|\frac{1}{n}\sum^{n}_{i=1}\widehat{\alpha}(X_{i})\phi_{j}(X_{i})-\frac{1}{n}\sum^{n}_{i=1}\phi_{j}(X_{i})\right|\leq\lambda\quad(\text{for}\ j=1,2,\dots,p).$
If we use a logistic link and UKL-Riesz regression, we have
$\left|\frac{1}{n}\sum^{n}_{i=1}\widehat{\alpha}(X_{i})\phi_{j}(X_{i})-1\right|\leq\lambda\quad(\text{for}\ j=1,2,\dots,p).$
We explain the details of the automatic covariate balancing property in the following subsections.
4.3 Choice of Loss and Link Functions
For the automatic covariate balancing property, the linearity assumption
$\partial g\big(\alpha_{{\bm{\beta}}}(X_{i})\big)=\sum^{p}_{j=1}\beta_{j}\phi_{j}(X_{i})$
or $\partial g\big(\alpha_{{\bm{\beta}}}(X_{i})\big)=\sum^{p}_{j=1}\beta_{j}\widetilde{g}\left(X_{i},\phi_{j}(X_{i})\right)$ are crucial. These assumptions depend on the choice of the function $g$ in the Bregman divergence, which corresponds to the loss function, and on the choice of the link function.
For example, for SQ-Riesz regression, UKL-Riesz regression, and BP-Riesz regression, the following choice of Riesz representer models (link functions) guarantee the linearity assumption.
  • Squared distance (SQ-Riesz): We use the identity link function,
    $\alpha_{{\bm{\beta}}}(X)$ $=\zeta^{-1}\Big(X,{\bm{\phi}}(X)^{\top}{\bm{\beta}}\Big)$
    $=({\bm{\phi}}(X)^{\top}{\bm{\beta}}+C)/2.$
  • UKL divergence loss (UKL-Riesz): We use a log or logistic link function,
    $\alpha_{{\bm{\beta}}}(X)$ $=\zeta^{-1}\Big(X,{\bm{\phi}}(X)^{\top}{\bm{\beta}}\Big)$
    $=\xi(X)\Big(C+\exp\big({\bm{\phi}}(X)^{\top}{\bm{\beta}}\big)\Big)-(1-\xi(X))\Big(C+\exp\big(-{\bm{\phi}}(X)^{\top}{\bm{\beta}}\big)\Big).$
  • BP divergence loss (BP-Riesz):
    $\alpha_{{\bm{\beta}}}(X)$ $=\zeta^{-1}\Big(X,{\bm{\phi}}(X)^{\top}{\bm{\beta}}\Big)$
    $=\xi(X)\left(C+\Big(1+\frac{{\bm{\phi}}(X)^{\top}{\bm{\beta}}}{k}\Big)^{1/\delta}\right)-(1-\xi(X))\left(C+\Big(1-\frac{{\bm{\phi}}(X)^{\top}{\bm{\beta}}}{k}\Big)^{1/\delta}\right),$
    for $k\coloneqq 1+1/\delta$.
Here, $\xi\colon{\mathcal{X}}\to\{1,0\}$ is chosen by the researcher. Note that these are not the only possible choices, and we can still choose link functions that satisfy the linearity assumption. For example, we can multiply $\alpha(X)={\bm{\phi}}(X)^{\top}{\bm{\beta}}+1$ by a function $\kappa\colon{\mathcal{X}}\to{\mathbb{R}}$. Such extensions and other cases are straightforward, so we omit them.
Justification.
We justify the above choices of the link functions by showing that under them, the automatic covariate balancing property holds. The derivative of $g$ is given as
  • Squared distance (SQ-Riesz): $\partial g^{\text{SQ}}(\alpha)=2(\alpha-C)$.
  • UKL divergence loss (UKL-Riesz): $\partial g^{\text{UKL}}(\alpha)=\operatorname{sign}(\alpha)\log\big(|\alpha|-C\big)$ for $\alpha<-C$ and $\alpha>C$. We introduce the following branchwise maps:
    $\partial g^{\text{UKL}}_{+}(\alpha)\coloneqq\log(\alpha-C),\qquad\alpha\in(C,\infty),$
    $\partial g^{\text{UKL}}_{-}(\alpha)\coloneqq-\log(-\alpha-C),\qquad\alpha\in(-\infty,-C),$
    so that $\partial g^{\text{UKL}}(\alpha)=\partial g^{\text{UKL}}_{+}(\alpha)$ for $\alpha>C$ and $\partial g^{\text{UKL}}(\alpha)=\partial g^{\text{UKL}}_{-}(\alpha)$ for $\alpha<-C$.
  • BP divergence loss (BP-Riesz): $\partial g^{\text{BP}}(\alpha)=\operatorname{sign}(\alpha)\Big(1+\frac{1}{\delta}\Big)\Big(\big(|\alpha|-C\big)^{\delta}-1\Big),\qquad\mathrm{dom}\left(\partial g^{\text{BP}}\right)=\{\alpha\in{\mathbb{R}}:|\alpha|\geq C\}$ for $\delta>0$, $\alpha<-C$, and $\alpha>C$. We introduce the following branchwise maps:
    $\partial g^{\text{BP}}_{+}(\alpha)\coloneqq\Big(1+\frac{1}{\delta}\Big)\Big((\alpha-C)^{\delta}-1\Big),\qquad\alpha\in[C,\infty),$
    $\partial g^{\text{BP}}_{-}(\alpha)\coloneqq-\Big(1+\frac{1}{\delta}\Big)\Big((-\alpha-C)^{\delta}-1\Big),\qquad\alpha\in(-\infty,-C],$
    so that $\partial g^{\text{BP}}(\alpha)=\partial g^{\text{BP}}_{+}(\alpha)$ for $\alpha\geq C$ and $\partial g^{\text{BP}}(\alpha)=\partial g^{\text{BP}}_{-}(\alpha)$ for $\alpha\leq-C$.
We introduce $\partial g^{\text{UKL}}_{\pm}$ and $\partial g^{\text{BP}}_{\pm}$ (and related functions) to make the inverse mapping one-to-one on each branch.
The inverse of these derivatives gives intuition for how to choose loss and link functions. The inverse functions are
  • Squared distance (SQ-Riesz):
    $\left(\partial g^{\text{SQ}}\right)^{-1}(v)\coloneqq(v+C)/2\qquad(v\in{\mathbb{R}}).$
  • UKL divergence loss (UKL-Riesz):
    $\left(\partial g^{\text{UKL}}_{+}\right)^{-1}(v)$ $=C+\exp(v),\qquad(v\in{\mathbb{R}}),$
    $\left(\partial g^{\text{UKL}}_{-}\right)^{-1}(v)$ $=-C-\exp(-v),\qquad(v\in{\mathbb{R}}).$
  • BP divergence loss (BP-Riesz):
    $\left(\partial g^{\text{BP}}_{+}\right)^{-1}(v)$ $=C+\Big(1+\frac{v}{k}\Big)^{1/\delta},\qquad(v\geq-k),$
    $\left(\partial g^{\text{BP}}_{-}\right)^{-1}(v)$ $=-C-\Big(1-\frac{v}{k}\Big)^{1/\delta},\qquad(v\leq k),$
    for $k\coloneqq 1+1/\delta$.
These inverse functions suggest the above link functions.
4.4 SQ-Riesz regression with a Linear Link Function
We first introduce the combination of the squared loss and a linear model. Consider the linear model
$\alpha_{\bm{\beta}}(X)={\bm{\phi}}(X)^{\top}{\bm{\beta}},$
where ${\bm{\phi}}\colon{\mathcal{X}}\to{\mathbb{R}}^{p}$ is a basis function.
For this model, using the squared distance (SQ-Riesz regression, or standard Riesz regression) yields automatic covariate balancing. Specifically, under a linear model, if we use $\ell_{1}$-penalty in SQ-Riesz regression, the dual formulation implies that SQ-Riesz regression is equivalent to solving
$\min_{\bm{\alpha}\in{\mathbb{R}}^{n}}$ $\ \ \ \frac{1}{n}\sum^{n}_{i=1}(\alpha_{i}-C)^{2}$
subjectto $\ \ \ \left|\frac{1}{n}\sum^{n}_{i=1}\Big(\alpha_{i}\phi_{j}(X_{i})-m(W_{i},\phi_{j})\Big)\right|\leq\lambda\qquad j=1,\dots,p.$
This optimization problem matches that used to obtain stable balancing weights (Zubizarreta, 2015). When $\lambda=0$, it enforces the covariate balancing condition
$\sum^{n}_{i=1}\widehat{\alpha}_{i}\phi(D_{i},Z_{i})-\left(\sum^{n}_{i=1}\Big(\phi(1,Z_{i})-\phi(0,Z_{i})\Big)\right)=0,$
where $\widehat{\alpha}_{i}={\bm{\phi}}(X_{i})^{\top}\widehat{{\bm{\beta}}}$.
An advantage of linear models is that we can express the entire ATE estimation problem using a single linear model, as shown by Bruns-Smith et al. (2025).
4.5 UKL-Riesz Regression with a Log Link Function
We next introduce the combination of the UKL divergence loss and a log link function. Consider the log link model
$\alpha_{\bm{\beta}}(X)=\xi(X)\Big(C+\exp\big({\bm{\phi}}(X)^{\top}{\bm{\beta}}\big)\Big)-(1-\xi(X))\Big(C+\exp\big(-{\bm{\phi}}(X)^{\top}{\bm{\beta}}\big)\Big).$
We call this a log link function because $\alpha_{\bm{\beta}}(X)=\zeta^{-1}(X,{\bm{\phi}}(X)^{\top}{\bm{\beta}})$ can be inverted in ${\bm{\phi}}(X)^{\top}{\bm{\beta}}$ by taking logarithms on each branch given $\xi(X)$. For example, in ATE estimation, we set $\xi(X)=D$ and $C=1$. In density ratio estimation, we set $\xi(X)=1$ and $C=0$, as discussed below.
For this model, using the UKL divergence (UKL-Riesz regression) yields automatic covariate balancing. Specifically, under this specification, if we use $\ell_{1}$-penalty in UKL-Riesz regression, the dual formulation implies that UKL-Riesz regression is equivalent to solving
$\min_{\bm{\alpha}\in{\mathbb{R}}^{n}}$ $\ \ \ \frac{1}{n}\sum^{n}_{i=1}\operatorname{sign}(\alpha_{i})\log\big(|\alpha_{i}|-C\big)$
subjectto $\ \ \ \left|\frac{1}{n}\sum^{n}_{i=1}\left(\alpha_{i}\Big(\xi(X_{i})\phi(X_{i})-(1-\xi(X_{i}))\phi(X_{i})\Big)-m\Big(W_{i},\phi_{j}(\cdot)\Big)\right)\right|\leq\lambda\qquad j=1,\dots,p.$
This optimization problem matches that used to obtain entropy balancing weights (Hainmueller, 2012). When $\lambda=0$, it enforces the covariate balancing condition
$\frac{1}{n}\sum^{n}_{i=1}\left(\alpha_{i}\Big(\xi(X_{i})\phi(X_{i})-(1-\xi(X_{i}))\phi(X_{i})\Big)-m\Big(W_{i},\phi_{j}(\cdot)\Big)\right)=0,$
where $\widehat{\alpha}_{i}=\alpha_{\widehat{{\bm{\beta}}}}(X_{i})$.
An advantage of a log link function is that it naturally imposes modeling assumptions. For example, when we assume a logistic model for the propensity score, the induced Riesz representer takes this log link form.
Special case of the log link function.
As a special case, we can model the Riesz representer as
$\alpha_{\bm{\beta}}(X)=\exp\left({\bm{\phi}}(X)^{\top}{\bm{\beta}}\right),$
which corresponds to $\xi(X)=1$ and $C=0$. Such a specification appears in density ratio estimation, as discussed below. For this model, using the UKL divergence (UKL-Riesz regression) yields automatic covariate balancing. Specifically, under this specification, the dual formulation implies that UKL-Riesz regression is equivalent to solving
$\min_{\bm{\alpha}\in{\mathbb{R}}^{n}}$ $\ \ \ \frac{1}{n}\sum^{n}_{i=1}\log\big(\alpha_{i}\big)$
subjectto $\ \ \ \left|\frac{1}{n}\sum^{n}_{i=1}\left(\alpha_{i}\phi_{j}(X_{i})-m\Big(W_{i},\phi_{j}(\cdot)\Big)\right)\right|\leq\lambda\qquad\text{for}\ \ j=1,\dots,p.$
This optimization problem matches that used to obtain entropy balancing weights in the density ratio setting. When $\lambda=0$, it enforces the covariate balancing condition
$\frac{1}{n}\sum^{n}_{i=1}\left(\alpha_{i}\phi_{j}(X_{i})-m\Big(W_{i},\phi_{j}(\cdot)\Big)\right)=0,$
where $\widehat{\alpha}_{i}=\alpha_{\widehat{{\bm{\beta}}}}(X_{i})$.
Exponential models, or log link functions, are closely related to density ratio modeling. When two probability densities $p(x)$ and $q(x)$ belong to exponential families, the density ratio can also be expressed in an exponential form. Exponential models also impose nonnegativity of the density ratio without sacrificing smoothness. For example, if we model a density ratio $r(x)$ by a linear model, the linear model can violate the nonnegativity condition, since $r(x)>0$ must hold.
4.6 BP-Riesz Regression and a Power Link Function
We next introduce a specification that pairs the BP divergence loss with a link function that interpolates between the linear link used for SQ-Riesz regression and the log link used for UKL-Riesz regression. This specification is useful both as a robustness device and as a way to understand how automatic covariate balancing varies continuously with the choice of loss and link functions.
Let $\delta\in(0,\infty)$ and define $k\coloneqq 1+1/\delta$. Consider the following model for the Riesz representer:
$\alpha_{\bm{\beta}}(X)=\xi(X)\left(C+\left(1+\frac{{\bm{\phi}}(X)^{\top}{\bm{\beta}}}{k}\right)^{1/\delta}\right)-(1-\xi(X))\left(C+\left(1-\frac{{\bm{\phi}}(X)^{\top}{\bm{\beta}}}{k}\right)^{1/\delta}\right),$ (5)
where ${\bm{\phi}}\colon{\mathcal{X}}\to{\mathbb{R}}^{p}$ is a basis function and $\xi\colon{\mathcal{X}}\to\{0,1\}$ selects the branch. We call a link function in (5) a power link function.
The choice of $(\xi,C)$ is application dependent. For example, in ATE estimation we typically set $\xi(X)=D$ and $C=1$, while in density ratio estimation we often set $\xi(X)=1$ and $C=0$ so that $\alpha_{\bm{\beta}}(X)$ is nonnegative by construction.
Under (5), the dual characterization implies that BP-Riesz regression returns the minimum BP-loss solution among approximately balancing models. In particular, if we use $\ell_{1}$-penalty in BP-Riesz regression, BP-Riesz regression is equivalent to solving a constrained problem of the form
$\min_{\bm{\alpha}\in{\mathbb{R}}^{n}}$ $\ \ \ \frac{1}{n}\sum^{n}_{i=1}g^{\text{BP}}(\alpha_{i})$
subjectto $\ \ \ \left|\frac{1}{n}\sum^{n}_{i=1}\Big(\alpha_{i}\phi_{j}(X_{i})-m(W_{i},\phi_{j})\Big)\right|\leq\lambda\qquad\text{for}\ \ j=1,\dots,p,$ (6)
with $\alpha_{i}$ restricted to the domain of $g^{\text{BP}}$, that is, $|\alpha_{i}|\geq C$. When $\lambda=0$, the constraint (6) enforces exact balancing:
$\frac{1}{n}\sum^{n}_{i=1}\widehat{\alpha}_{i}\phi_{j}(X_{i})=\frac{1}{n}\sum^{n}_{i=1}m(W_{i},\phi_{j}),\qquad\text{for}\ \ j=1,\dots,p,$
where $\widehat{\alpha}_{i}=\alpha_{\widehat{{\bm{\beta}}}}(X_{i})$.
Relationship to the linear and log links.
The power link (5) provides a continuous bridge between the specifications in the previous subsections. As $\delta\to 0$, we have $k=1+1/\delta\to\infty$ and
$\Big(1+\frac{t}{k}\Big)^{1/\delta}=\Big(1+\delta t+o(\delta)\Big)^{1/\delta}\to\exp(t),$
so (5) reduces to the log link form used for UKL-Riesz regression. At $\delta=1$, the BP loss reduces to the squared loss, and the link becomes an affine transformation of the linear index around the origin, which connects BP-Riesz regression to SQ-Riesz regression up to reparameterization.
This interpolation perspective is also consistent with the robustness interpretation of the BP divergence (Basu et al., 1998; Sugiyama et al., 2012). Smaller $\delta$ makes the objective closer to a KL-type criterion, which can be efficient under correct specification, while larger $\delta$ yields behavior closer to squared loss and is typically more robust to misspecification and extreme weights.
5 Automatic Neyman Orthogonalization
The advantage of (automatic) covariate balancing lies in the fact that it automatically implies Neyman orthogonality. Specifically, we consider an estimator given by
$\widehat{\theta}\coloneqq\frac{1}{n}\sum_{i=1}^{n}\widehat{\alpha}(X_{i})Y_{i},$
where $\widehat{\alpha}$ is an estimator of the Riesz representer obtained from generalized Riesz regression. This estimator corresponds to the IPW estimator in ATE estimation (Horvitz & Thompson, 1952) and covariate shift adaptation with importance weighting (Shimodaira, 2000). We show that this estimator satisfies Neyman orthogonality if the regression function $\gamma_{0}$ belongs to the linear space spanned by ${\bm{\phi}}(X)$, which is used for Riesz representer estimation. We refer to this property as automatic Neyman orthogonalization.
5.1 Automatic Neyman Orthogonalization
The following result holds.
Theorem 5.1.
Consider generalized Riesz regression with the model
$\alpha(X)=\zeta^{-1}\Big(X,{\bm{\phi}}(X)^{\top}{\bm{\beta}}\Big),$
where $\zeta^{-1}$ is the inverse of a link function and ${\bm{\phi}}\colon{\mathcal{X}}\to{\mathbb{R}}^{p}$ is a vector of basis functions. Suppose that $\partial g\big(\alpha_{{\bm{\beta}}}(X_{i})\big)=\sum^{p}_{j=1}\beta_{j}\phi_{j}(X_{i})$ holds. If $\gamma_{0}$ belongs to a linear space of ${\bm{\phi}}(X)$, then it holds that
$\widehat{\theta}$ $\coloneqq\frac{1}{n}\sum^{n}_{i=1}\widehat{\alpha}(X_{i})Y_{i}$
$=\frac{1}{n}\sum^{n}_{i=1}\Big(\widehat{\alpha}(X_{i})\big(Y_{i}-\gamma_{0}(X_{i})\big)+m(W_{i},\gamma_{0})\Big).$
If $\widehat{\alpha}$ satisfies the Donsker condition and is consistent, then $\widehat{\theta}$ is asymptotically efficient.
The derivation is as follows. Recall that the sample average of the Neyman orthogonal score is given by
$\frac{1}{n}\sum^{n}_{i=1}\psi(W;\widetilde{\eta},\theta)=\frac{1}{n}\sum^{n}_{i=1}\Big(\widehat{\alpha}(X_{i})\big(Y_{i}-\gamma_{0}(X_{i})\big)+m(W_{i},\gamma_{0})-\theta\Big),$
[…]
In Section 4, we model the Riesz representer as $\alpha(X)=\zeta^{-1}\Big(X,{\bm{\phi}}(X)^{\top}{\bm{\beta}}\Big)$. Then, a Riesz representer estimator $\widehat{\alpha}$ trained by generalized Riesz regression with $\lambda=0$ satisfies
$\frac{1}{n}\sum^{n}_{i=1}\Big(\widehat{\alpha}(X_{i})\phi_{j}(X_{i})-m(W_{i},\phi_{j})\Big)=0$
for all $j=1,2,\dots,p$ by the automatic covariate balancing property (Corollary 4.2).
If $\gamma_{0}(X)$ is given as $\gamma_{0}(X)={\bm{\phi}}(X)^{\top}{\bm{\rho}}_{0}$ for some ${\bm{\rho}}_{0}\in{\mathbb{R}}^{p}$, then we have
$\frac{1}{n}\sum^{n}_{i=1}\psi(W;\widetilde{\eta},\theta)$
$=\frac{1}{n}\sum^{n}_{i=1}\Big(\widehat{\alpha}(X_{i})\big(Y_{i}-\gamma_{0}(X_{i})\big)+m(W_{i},\gamma_{0})-\theta\Big),$
$=\frac{1}{n}\sum^{n}_{i=1}\Big(\widehat{\alpha}(X_{i})\big(Y_{i}-{\bm{\rho}}_{0}^{\top}{\bm{\phi}}(X_{i})\big)+{\bm{\rho}}_{0}^{\top}m(W_{i},{\bm{\phi}}(X))-\theta\Big)$
$=\frac{1}{n}\sum^{n}_{i=1}\Big(\widehat{\alpha}(X_{i})Y_{i}-\theta\Big),$
where we used
$\sum^{p}_{j=1}\rho_{0,j}\frac{1}{n}\sum^{n}_{i=1}\Big(\widehat{\alpha}(X_{i})\phi_{j}(X_{i})-m(W_{i},\phi_{j})\Big)=0$
from the automatic covariate balancing property.
Such a property has been reported by Wong & Chan (2017) and Zhao (2019) in ATE estimation. Although they require a constant treatment effect, our results generalize them by allowing heterogeneous treatment effects. We emphasize that (i) the Riesz representer estimator $\widehat{\alpha}$ does not require a specific convergence rate; (ii) we cannot use cross fitting, and we assume the Donsker condition (if we use cross fitting, the covariate balancing collapses).
5.2 Automatic Neyman Error Minimization
Covariate balancing can also be interpreted as minimizing the estimation error of the Neyman orthogonal score. Given a nuisance parameter estimator $\widehat{\eta}$ and the true nuisance parameter $\eta_{0}$, the estimation error is given by
$\psi(W;\widehat{\eta},\theta)-\psi(W;\eta_{0},\theta)$
$=\Big(m(W,\widehat{\gamma})+\widehat{\alpha}(X)\big(Y-\widehat{\gamma}(X)\big)\Big)-\Big(m(W,\gamma_{0})+\alpha_{0}(X)\big(Y-\gamma_{0}(X)\big)\Big).$
Here, $\alpha_{0}(X)\big(Y-\gamma_{0}(X)\big)$ has mean zero and is ignorable in expectation. Therefore, we consider
$\text{NeymanError}\coloneqq\frac{1}{n}\sum^{n}_{i=1}\left(\widehat{\alpha}(X_{i})\big(Y_{i}-\widehat{\gamma}(X_{i})\big)+m(W_{i},\widehat{\gamma})-m(W_{i},\gamma_{0})\right).$
Suppose that we estimate $\alpha_{0}$ by generalized Riesz regression using the basis function ${\bm{\phi}}(X)$, and that $\gamma_{0}$ belongs to the linear space spanned by ${\bm{\phi}}(X)$. Recall that we have
$\sum^{p}_{j=1}\rho_{0,j}\frac{1}{n}\sum^{n}_{i=1}\Big(\widehat{\alpha}(X_{i})\phi_{j}(X_{i})-m(W_{i},\phi_{j})\Big)=0.$
If $\widehat{\gamma}$ equals $\gamma_{0}$ or $\widehat{\gamma}$ is estimated as ${\bm{\phi}}(X)^{\top}{\bm{\rho}}$, then by Neyman orthogonality, we have
$\text{NeymanError}=\frac{1}{n}\sum^{n}_{i=1}\left(\widehat{\alpha}(X_{i})Y_{i}-m(W_{i},\gamma_{0})\right).$
Furthermore, since $Y=\gamma_{0}(X)+\varepsilon$ for an error term $\varepsilon$ such that ${\mathbb{E}}[\varepsilon\mid X]=0$, we have
$\text{NeymanError}=\frac{1}{n}\sum^{n}_{i=1}\left(\widehat{\alpha}(X_{i})\gamma_{0}(X_{i})-m(W_{i},\gamma_{0})\right)+U,$
where $U$ is a term whose expectation is zero. Here, we again have
$\frac{1}{n}\sum^{n}_{i=1}\left(\widehat{\alpha}(X_{i})\gamma_{0}(X_{i})-m(W_{i},\gamma_{0})\right)=0$
from the automatic covariate balancing property. Thus, automatic covariate balancing and automatic Neyman orthogonalization also imply minimization of the Neyman error.
This property is discussed in Zhao (2019) for the case where UKL-Riesz regression with a basis function ${\bm{\phi}}\colon{\mathcal{Z}}\to{\mathbb{R}}^{p}$ is used in ATE estimation. Our result generalizes the finding in that work to more general models, losses, and parameters of interest.
Cross fitting and the Neyman error
If we use cross fitting, the exact automatic covariate balancing property may not hold. On the other hand, exact automatic covariate balancing typically requires the Donsker condition to obtain an efficient estimator of the parameter of interest. We need to carefully address this trade-off.
Basis functions depending only on $Z$
In ATE estimation with standard propensity score modeling, we typically use basis functions depending on $Z$, that is, ${\bm{\phi}}\colon{\mathcal{Z}}\to{\mathbb{R}}^{p}$. Under this choice, $m(W,\phi_{j}(\cdot))=0$. Therefore, automatic covariate balancing guarantees only $\frac{1}{n}\sum_{i=1}^{n}\left(\widehat{\alpha}(X_{i})\gamma_{0}(X_{i})\right)=0$. This result eliminates part of the Neyman error only when the conditional ATE given $x$ is invariant across $x$, that is, when the treatment effect is homogeneous. From the regression-function point of view, this requires that $\gamma_{0}(X)$ lies in the linear space spanned by ${\bm{\phi}}(Z)$. If we aim to guarantee automatic Neyman orthogonalization under treatment effect heterogeneity, we should use ${\bm{\phi}}\colon{\mathcal{X}}\to{\mathbb{R}}^{p}$, even though the Riesz representer corresponds to inverse propensity scores $1/e_{0}(Z)$ and $1/(1-e_{0}(Z))$.
5.3 Comparison with TMLE
TMLE is another promising approach in debiased machine learning (van der Laan, 2006). TMLE adds a perturbation to an initial estimate of $\gamma_{0}$. TMLE works to eliminate the sample average of the $(\star)$ part in the following Neyman error:
$\text{NeymanError}\coloneqq\frac{1}{n}\sum^{n}_{i=1}\left(\underbrace{\widehat{\alpha}(X_{i})(Y_{i}-\widehat{\gamma}(X_{i}))}_{\text{= $(\star)$}}+m(W_{i},\widehat{\gamma})-m(W_{i},\gamma_{0})\right).$
In contrast, generalized Riesz regression works to eliminate the sample average of the $(\star\star)$ part in the following Neyman error:
$\text{NeymanError}\coloneqq\frac{1}{n}\sum^{n}_{i=1}\left(\widehat{\alpha}(X_{i})(Y_{i}-\underbrace{\widehat{\gamma}(X_{i}))+m(W_{i},\widehat{\gamma})}_{\text{= $(\star\star)$}}-m(W_{i},\gamma_{0})\right).$
That is, TMLE and generalized Riesz regression differ as follows:
  • TMLE puts the difficulty of efficient estimation of $\theta_{0}$ into the estimation of $\gamma_{0}$ by eliminating the influence of $\widehat{\alpha}$.
  • Generalized Riesz regression puts the difficulty of efficient estimation of $\theta_{0}$ into the estimation of $\alpha_{0}$ by eliminating the influence of $\widehat{\gamma}$.
Remark (TMLE).
We introduce TMLE in more detail. Let $\widehat{\gamma}^{(0)}$ be an initial estimate of $\gamma_{0}$. Then, given estimates $\widehat{\gamma}^{(0)}$ and $\widehat{\alpha}$, TMLE updates the regression function as
$\widehat{\gamma}^{(1)}(x)\coloneqq\widehat{\gamma}^{(0)}(x)+\frac{\sum_{i=1}^{n}\widehat{\alpha}(X_{i})\big(Y_{i}-\widehat{\gamma}^{(0)}(X_{i})\big)}{\sum_{i=1}^{n}\widehat{\alpha}(X_{i})^{2}}\widehat{\alpha}(x).$
It then estimates the parameter of interest as
$\widehat{\theta}^{\text{TMLE}}\coloneqq\frac{1}{n}\sum^{n}_{i=1}m(W_{i},\widehat{\gamma}^{(1)}).$
Note that this update is derived as the solution in $\epsilon$ of
$\sum^{n}_{i=1}\widehat{\alpha}(X_{i})\left(Y_{i}-\left(\widehat{\gamma}^{(0)}(X_{i})+\epsilon\widehat{\alpha}(X_{i})\right)\right)=0,$
which is given by
$\widehat{\epsilon}\coloneqq\frac{\sum_{i=1}^{n}\widehat{\alpha}(X_{i})\{Y_{i}-\widehat{\gamma}^{(0)}(X_{i})\}}{\sum_{i=1}^{n}\widehat{\alpha}(X_{i})^{2}}.$
Then, we derive $\widehat{\gamma}^{(1)}(x)$ as
$\widehat{\gamma}^{(1)}(x)=\widehat{\gamma}^{(0)}(x)+\widehat{\epsilon}\widehat{\alpha}(x).$
Under this update, the sample average of $(\star)$ in the Neyman error becomes zero.
6 Convergence Rate Analysis
This section provides an estimation error analysis for generalized Riesz regression. We model the Riesz representer $\alpha_{0}$ by
$\alpha_{f}(X)=\zeta^{-1}\Big(X,f(X)\Big),$
where $\zeta$ is a continuously differentiable and globally Lipschitz link function, and $f$ is a base model. Note that unlike Section 4, we do not restrict $f$ to be a linear model. For example, in addition to linear models ${\bm{\phi}}(X)^{\top}{\bm{\beta}}$, we can use random forests, neural networks, and other models for $f$. In this section, we consider the case where we use RKHS methods and neural networks for $f$.
Throughout this section, we assume that the Riesz representer is bounded.
Assumption 6.1.
There exists a constant $C>0$ independent of $n$ such that $|\alpha(x)|<C$ for all $x\in{\mathcal{X}}$.
This boundedness assumption holds in the standard ATE setting, which assumes common support of the treated and control groups and bounded outcomes. In many other applications, this assumption also holds. If we wish to allow unbounded support, we can develop such an extension by imposing appropriate tail conditions. For example, the density ratio between two Normal distributions may violate this assumption. In such cases, Zheng et al. (2022) presents a convergence rate analysis, and we can follow their approach. In practical data analysis, it is often reasonable to treat the Riesz representer as bounded.
6.1 RKHS
First, we study the case with RKHS regression. Let ${\mathcal{F}}^{\text{RKHS}}$ be a class of RKHS functions, and define
$\widehat{f}^{\text{RKHS}}\coloneqq\operatorname*{arg\,min}_{f\in{\mathcal{F}}^{\text{RKHS}}}\widehat{\text{BD}}_{g}(\alpha_{f})+\lambda\|f\|^{2}_{{\mathcal{F}}},$
where $\|\cdot\|^{2}_{{\mathcal{F}}}$ is the RKHS norm. Then, we define an estimator as
$\widehat{\alpha}^{\text{RKHS}}(x)\coloneqq\alpha_{\widehat{f}^{\text{RKHS}}}(x)\coloneqq\zeta^{-1}\left(x,\widehat{f}^{\text{RKHS}}(x)\right).$
We analyze the estimation error by employing the results in Kanamori et al. (2012), which studies RKHS based LSIF for DRE. We define the following localized class of RKHS functions as a technical device: ${\mathcal{F}}^{\text{RKHS}}_{M}\coloneqq\big\{f\in{\mathcal{F}}^{\text{RKHS}}\colon I(f)\leq M\big\}$ for some norm $I(f)$ of $f$. We also define ${\mathcal{A}}^{\text{RKHS}}\coloneqq\left\{\zeta^{-1}(\cdot,f(\cdot))\colon f\in{\mathcal{F}}^{\text{RKHS}}\right\}$. We then impose the following assumption on this localized class.
Assumption 6.2.
There exist constants $0<\tau<2$, $0\leq{\bm{\beta}}\leq 1$, $c_{0}>0$, and $A>0$ such that for all $M\geq 1$, it holds that $H_{B}(\delta,{\mathcal{F}}^{\text{RKHS}}_{M},P_{0})\leq A\left(\frac{M}{\delta}\right)^{\tau}$, where $H_{B}(\delta,{\mathcal{F}}^{\text{RKHS}}_{M},P_{0})$ is the bracketing entropy with radius $\delta>0$ for the function class ${\mathcal{F}}^{\text{RKHS}}_{M}$ and the distribution $P_{0}$.
For details on bracketing entropy, see Appendix E and Definition 2.2 in van de Geer (2000).
Under these preparations, we establish an estimation error bound.
Theorem 6.1 ($L_{2}$-norm estimation error bound).
Suppose that $g$ is $\mu$-strongly convex and there exists a constant $C>0$ such that $|g^{\prime\prime}(t)|\leq C\quad\forall t\in{\mathbb{R}}$. Assume also that $\zeta^{-1}(0)$ is finite. Suppose that Assumptions 6.1 and 6.2 hold. Set the regularization parameter $\lambda=\lambda_{n}$ so that $\lim_{n\to\infty}\lambda_{n}=0$ and $\lambda^{-1}_{n}=O(n^{1-\delta})$ ($n\to\infty$). If $\alpha_{0}\in{\mathcal{A}}^{\text{RKHS}}$, then we have
$\left\|\widehat{\alpha}^{\text{RKHS}}(X)-\alpha_{0}(X)\right\|^{2}_{L_{2}(P_{0})}=O_{P_{0}}\left(\lambda^{1/2}\right).$
The proof is provided in Appendix E, following the approach of Kanamori et al. (2012). The parameter $\tau$ is determined by the function class to which $f_{0}$ belongs.
6.2 Neural Networks
Second, we provide an estimation error analysis when we use neural networks for $\mathcal{A}$}. Our analysis is mostly based on Kato & Teshima (2021) and Zheng et al. (2022). We define Feedforward neural networks (FNNs) as follows:
Definition 6.1 (FNNs. From Zheng et al. (2022)).
Let $\mathcal{D}$}, $\mathcal{W}$}, $\mathcal{U}$}, and ${\mathcal{S}}\in(0,\infty)$ be parameters that can depend on $n$. Let ${\mathcal{F}}^{\text{FNN}}\coloneqq{\mathcal{F}}^{\text{FNN}}_{M,{\mathcal{D}},{\mathcal{W}},{\mathcal{U}},{\mathcal{S}}}$ be a class of ReLU activated FNNs with parameter $\bm{\beta}$}, depth $\mathcal{D}$}, width $\mathcal{W}$}, size $\mathcal{S}$}, and number of neurons $\mathcal{U}$}, and satisfying the following conditions: (i) the number of hidden layers is $\mathcal{D}$}; (ii) the maximum width of the hidden layers is $\mathcal{W}$}; (iii) the number of neurons in $e_{\bm{\beta}}$ is $\mathcal{U}$}; (iv) the total number of parameters in $e_{\bm{\beta}}$ is $\mathcal{S}$}.
For the model ${\mathcal{F}}^{\text{FNN}}$, we define $\widehat{f}^{\text{FNN}}\coloneqq\operatorname*{arg\,min}_{f\in{\mathcal{F}}^{\text{FNN}}}\widehat{\text{BD}}_{g}(\alpha_{f})$. Then, we define an estimator as
$\widehat{\alpha}^{\text{FNN}}(x)\coloneqq\zeta^{-1}\left(x,\widehat{f}^{\text{FNN}}(x)\right).$
For this estimator, we can prove an estimation error bound. We make the following assumption.
Assumption 6.3.
There exists a constant $0<M<\infty$ such that $\|f_{0}\|_{\infty}<M$, and $\|f\|_{\infty}\leq M$ for any $f\in{\mathcal{F}}^{\text{FNN}}$.
Let $\text{Pdim}({\mathcal{F}}^{\text{FNN}})$ be the pseudodimension of ${\mathcal{F}}^{\text{FNN}}$. For the definition, see Anthony & Bartlett (1999) and Definition 3 in Zheng et al. (2022). Then, we prove the following estimation error bound:
Theorem 6.2 (Estimation error bound for neural networks).
Suppose that $g$ is $\mu$-strongly convex and there exists a constant $C>0$ such that $|g^{\prime\prime}(t)|\leq C\quad\forall t\in{\mathbb{R}}$. Assume also that $\zeta^{-1}(0)$ is finite. Suppose that Assumption 6.3 holds. For $f_{0}$ such that
$\alpha_{0}(x)=\zeta^{-1}(x,f_{0}(x)),$
also assume $f_{0}\in\Sigma({\bm{\beta}},M,[0,1]^{d})$ with ${\bm{\beta}}=k+a$, where $k\in{\mathbb{N}}^{+}$ and $a\in(0,1]$, and ${\mathcal{F}}^{\text{FNN}}$ has width $\mathcal{W}$} and depth $\mathcal{D}$} such that ${\mathcal{W}}=38\big(\lfloor{\bm{\beta}}\rfloor+1\big)^{2}d^{\lfloor{\bm{\beta}}\rfloor+1}$ and ${\mathcal{D}}=21\big(\lfloor{\bm{\beta}}\rfloor+1\big)^{2}\lceil n^{\frac{d}{2(d+2{\bm{\beta}})}}\log_{2}\left(8n^{\frac{d}{2(d+2{\bm{\beta}})}}\right)\rceil$. Then, for $M\geq 1$ and $n\leq\text{Pdim}({\mathcal{F}}^{\text{FNN}})$, it holds that
$\left\|\widehat{\alpha}^{\text{FNN}}(X)-\alpha_{0}(X)\right\|^{2}_{L_{2}(P_{0})}=C_{0}\big(\lfloor{\bm{\beta}}\rfloor+1\big)^{9}d^{2\lfloor{\bm{\beta}}\rfloor+({\bm{\beta}}\land 3)}n^{-\frac{2{\bm{\beta}}}{d+2{\bm{\beta}}}}\log^{3}n,$
where $C_{0}>0$ is a constant independent of $n$.
The proof is provided in Appendix F, following the approach of Zheng et al. (2022). This result directly implies the minimax optimality of the proposed method when $f_{0}$ belongs to a Hölder class.
6.3 Construction of an Efficient Estimator
This section provides how we construct an efficient estimator for the parameter of interest $\theta_{0}$ using generalized Riesz regression. As we discussed in Section 2, we construct an estimator $\widehat{\theta}$ of $\theta_{0}$ as
$\frac{1}{n}\sum^{n}_{i=1}\psi\left(W_{i};\widehat{\eta},\widehat{\theta}\right)=0,$
where $\psi(W;\eta,\theta)$ is the Neyman orthogonal score is defined as
$\psi(W;\eta,\theta)\coloneqq m(W,\gamma)+\alpha(X)\big(Y-\gamma(X)\big)-\theta_{0}$
for $\eta=(\alpha,\gamma)$ ($\alpha,\gamma\colon{\mathcal{X}}\to{\mathbb{R}}$). As introduced in Section 2, we refer to this estimator as the AIPW estimator. We prove that under certain conditions, the proposed estimator is asymptotically normal.
We first make the following assumption.
Assumption 6.4 (Donsker condition or cross fitting).
Either of the followings holds: (i) the hypothesis classes $\mathcal{A}$} and $\mathcal{M}$} belong to the Donsker class, or (ii) $\widehat{\gamma}$ and $\widehat{\alpha}$ are estimated via cross fitting.
For example, the Donsker condition holds when the bracketing entropy of $\mathcal{A}$} is finite. In contrast, it is violated in high-dimensional regression or series regression settings where the model complexity diverges as $n\to\infty$. For neural networks, the assumption holds if both the number of layers and the width are finite. However, if these quantities grow with the sample size, the assumption is no longer valid.
Even if the Donsker condition does not hold, we can still establish asymptotic normality by employing sample splitting (Klaassen, 1987). There are various ways to implement sample splitting, and one of the most well-known is cross fitting, used in debiased machine learning (Chernozhukov et al., 2018). In debiased machine learning, the dataset is split into several folds, and the nuisance parameters are estimated using only a subset of the folds. This ensures that in $m\big(W_{i},\widehat{\gamma}\big)+\widehat{\alpha}(X_{i})\left(Y_{i}-\widehat{\gamma}(X_{i})\right)$, the observations $(X_{i},Y_{i})$ are not used to construct $\widehat{\gamma}$ and $\widehat{\alpha}$. For more details, see Chernozhukov et al. (2018).
Assumption 6.5 (Convergence rate).
$\big\|\widehat{\alpha}-\alpha_{0}\big\|_{2}=o_{p}(1)$, $\big\|\widehat{\gamma}-\gamma_{0}\big\|_{2}=o_{p}(1)$, and $\big\|\widehat{\alpha}-\alpha_{0}\big\|_{2}\big\|\widehat{\gamma}-\gamma_{0}\big\|_{2}=o_{p}(1/\sqrt{n})$.
Under these assumptions, we show the asymptotic normality of $\widehat{\theta}$. We omit the proof. For details, see Chernozhukov et al. (2018) or Schuler & van der Laan (2024), for example.
Theorem 6.3 (Asymptotic normality).
Suppose that Assumptions 6.1, and 6.46.5 hold. Then, the AIPW estimator converges in distribution to a normal distribution as
$\sqrt{n}\Big(\widehat{\theta}-\theta_{0}\Big)\xrightarrow{{\mathrm{d}}}{\mathcal{N}}(0,V^{*}),$
where $V^{*}$ is the efficiency bound defined as
$V^{*}\coloneqq{\mathbb{E}}\left[\psi(W;\eta_{0},\theta_{0})^{2}\right].$
Here, $V^{*}$ matches the efficiency bound given as the variance of the efficient influence function (van der Vaart, 1998; Hahn, 1998). Thus, this estimator is efficient.
7 Applications
This section provides applications of generalized Riesz regression: ATE estimation, AME estimation, and covariate shift adaptation (density ratio estimation).
7.1 ATE Estimation.
In ATE estimation, the linear functional is
$m^{\text{ATE}}(X,\gamma)\coloneqq\gamma(1,Z)-\gamma(0,Z),$
and the Riesz representer is
$\alpha^{\text{ATE}}_{0}(X)=\frac{D}{e_{0}(Z)}-\frac{1-D}{1-e_{0}(Z)},$
where $e_{0}(Z)=P(D=1\mid Z)$ is the propensity score. Let $r_{0}(1,Z)\coloneqq\frac{1}{e_{0}(Z)}$ and $r_{0}(0,Z)\coloneqq\frac{1}{1-e_{0}(Z)}$ be the inverse propensity score, also called the density ratio. We estimate $\alpha^{\text{ATE}}_{0}$ by minimizing the empirical Bregman divergence objective $\widehat{\text{BD}}_{g}(\alpha)$ introduced in Section 3, with $m=m^{\text{ATE}}$, an application-specific choice of $g$, and a model class for $\alpha$.
SQ-Riesz Regression.
We take the squared loss,
$g^{\text{SQ}}(\alpha)=\alpha^{2},$
and minimize the corresponding empirical Bregman objective. By substituting $g^{\text{SQ}}$ into (1) and using $m^{\text{ATE}}(X,\gamma)=\gamma(1,Z)-\gamma(0,Z)$, we obtain, up to an additive constant that does not depend on $\alpha$,
$\text{BD}_{g^{\text{SQ}}}(\alpha)={\mathbb{E}}\left[\alpha(D,Z)^{2}-2\big(\alpha(1,Z)-\alpha(0,Z)\big)\right].$
Thus, SQ-Riesz regression estimates $\alpha^{\text{ATE}}_{0}$ by
$\widehat{\alpha}\coloneqq\operatorname*{arg\,min}_{\alpha\in{\mathcal{A}}}\widehat{\text{BD}}_{g^{\text{SQ}}}(\alpha)+\lambda J(\alpha),$
where
$\widehat{\text{BD}}_{g^{\text{SQ}}}(\alpha)\coloneqq\frac{1}{n}\sum_{i=1}^{n}\left(\alpha(D_{i},Z_{i})^{2}-2\big(\alpha(1,Z_{i})-\alpha(0,Z_{i})\big)\right).$
This coincides with Riesz regression in Chernozhukov et al. (2021) and corresponds to LSIF in density ratio estimation (Kanamori et al., 2009). With appropriate choices of $\mathcal{A}$}, it also recovers nearest neighbor matching-based constructions, as discussed in Kato (2025c).
Remark (Linear link function).
A recommended Riesz representer modeling is
$\alpha_{\bm{\beta}}(X)={\bm{\phi}}(X)^{\top}{\bm{\beta}},$
where ${\bm{\phi}}\colon{\mathcal{X}}\to{\mathbb{R}}^{p}$ is a basis function. Under this model, minimizing the SQ-Riesz regression yields an estimator that satisfies an automatic covariate balancing property, as discussed in Section 4 and Zhao (2019).
Concretely, letting $\widehat{\alpha}=\alpha_{\widehat{{\bm{\beta}}}}$, we estimate $\bm{\beta}$} by
$\widehat{{\bm{\beta}}}$ $\coloneqq\operatorname*{arg\,min}_{{\bm{\beta}}}\frac{1}{n}\sum_{i=1}^{n}\Bigg(\Big({\bm{\phi}}(X)^{\top}{\bm{\beta}}\Big)^{2}-2\Big({\bm{\phi}}(1,Z)^{\top}-{\bm{\phi}}(0,Z)^{\top}\Big){\bm{\beta}}\Bigg)+\frac{1}{a}\lambda\|{\bm{\beta}}\|^{a}_{a}.$
By duality, if $a=1$, SQ-Riesz regression is equivalent to the following covariate balancing problem:
$\min_{\alpha\in{\mathbb{R}}^{n}}\$$\sum_{i=1}^{n}\alpha_{i}^{2}$
s.t. $\ \ \left|\sum_{i=1}^{n}\alpha_{i}\phi_{j}(X_{i})+\phi_{j}(1,Z_{i})-\phi_{j}(0,Z_{i})\right|\leq\lambda\ \ \ \text{for}\ j=1,2,\dots,p,$
where the solution $\widehat{w}_{i}$ corresponds to the estimator of $\alpha_{0}(X_{i})$ if $D_{i}=1$ and that of $\alpha_{0}(X_{i})$ if $D_{i}=0$; that is,
$\widehat{w}_{i}=\begin{cases}\widehat{\alpha}(1,Z_{i})&\text{if }D_{i}=1,\\ -\widehat{\alpha}(0,Z_{i})&\text{if }D_{i}=0.\end{cases}.$
UKL-Riesz Regression.
Consider Riesz representer models $\alpha$ such that $\alpha(1,x)>1$ and $\alpha(0,x)<-1$ for all $x$. We next use the UKL divergence loss with $C=1$,
$g^{\text{UKL}}(\alpha)=(|\alpha|-1)\log\left(|\alpha|-1\right)-|\alpha|.$
By substituting $g^{\text{UKL}}$ into (1) and using $m^{\text{ATE}}(X,\gamma)=\gamma(1,Z)-\gamma(0,Z)$, we obtain
$\text{BD}_{g^{\text{UKL}}}\big(\alpha\big)$ $\coloneqq\mathbb{E}\Big[\log\left(|\alpha(X)|-1\right)+|\alpha(X)|$
$\ \ \ \ \ -\Big(\operatorname{sign}\big(\alpha(1,Z)\big)\log\left(|\alpha(1,Z)|-1\right)-\operatorname{sign}\big(\alpha(0,Z)\big)\log\left(|\alpha(0,Z)|-1\right)\Big)\Big].$
Thus, UKL-Riesz regression estimates $\alpha^{\text{ATE}}_{0}$ by
$\widehat{\alpha}\coloneqq\operatorname*{arg\,min}_{\alpha\in{\mathcal{A}}}\widehat{\text{BD}}_{g^{\text{UKL}}}(\alpha)+\lambda J(\alpha),$
where
$\text{BD}_{g^{\text{UKL}}}\big(\alpha\big)$ $\coloneqq\frac{1}{n}\sum^{n}_{i=1}\Big(\log\left(|\alpha(X_{i})|-1\right)+|\alpha(X_{i})|$
$\ \ \ \ \ -\Big(\operatorname{sign}\big(\alpha(1,Z_{i})\big)\log\left(|\alpha(1,Z_{i})|-1\right)-\operatorname{sign}\big(\alpha(0,Z_{i})\big)\log\left(|\alpha(0,Z_{i})|-1\right)\Big)\Big).$
This coincides with the tailored loss minimization with $\alpha=\beta=-1$ in Zhao (2019) and corresponds to KLIEP in density ratio estimation (Sugiyama et al., 2008).
Remark (Log link function).
A recommended Riesz representer modeling is
$\alpha_{\bm{\beta}}(X)=\mathbbm{1}[D=1]\Big(1+\exp\big(-{\bm{\phi}}(X)^{\top}{\bm{\beta}}\big)\Big)-\mathbbm{1}[D=0]\Big(1+\exp\big({\bm{\phi}}(X)^{\top}{\bm{\beta}}\big)\Big),$
where ${\bm{\phi}}\colon{\mathcal{X}}\to{\mathbb{R}}^{p}$ is a basis function. Under this model, minimizing the UKL-Riesz regression yields an estimator that satisfies an automatic covariate balancing property, as discussed in Section 4 and Zhao (2019).
Concretely, letting $\widehat{\alpha}=\alpha_{\widehat{{\bm{\beta}}}}$, we estimate $\bm{\beta}$} by
$\widehat{{\bm{\beta}}}$ $\coloneqq\operatorname*{arg\,min}_{{\bm{\beta}}}\frac{1}{n}\sum_{i=1}^{n}\Bigg(\mathbbm{1}[D_{i}=1]\left(-{\bm{\phi}}(1,Z)^{\top}{\bm{\beta}}+1+\exp\big(-{\bm{\phi}}(1,Z)^{\top}{\bm{\beta}}\big)\right)$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\mathbbm{1}[D_{i}=0]\left({\bm{\phi}}(0,Z)^{\top}{\bm{\beta}}+1+\exp\big({\bm{\phi}}(0,Z)^{\top}{\bm{\beta}}\big)\right)$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ -\Big({\bm{\phi}}(1,Z)^{\top}-{\bm{\phi}}(0,Z)^{\top}\Big){\bm{\beta}}\Bigg)+\frac{1}{a}\lambda\|{\bm{\beta}}\|^{a}_{a}.$
By duality, if $a=1$, UKL-Riesz regression is equivalent to the following covariate balancing problem:
$\min_{{\bm{w}}\in(1,\infty)^{n}}\$$\sum_{i=1}^{n}(w_{i}-1)\log(w_{i}-1)$
s.t. $\ \ \left|\sum_{i=1}^{n}\Big(\mathbbm{1}[D_{i}=1]w_{i}\phi_{j}(X_{i})-\mathbbm{1}[D_{i}=0]w_{i}\phi_{j}(X_{i})\Big)-\Big(\phi_{j}(1,Z_{i})-\phi_{j}(0,Z_{i})\Big)\right|\leq\lambda$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \text{for}\ j=1,2,\dots,p,$
where the solution $\widehat{w}_{i}$ corresponds to the estimator of $\alpha_{0}(X_{i})$ if $D_{i}=1$ and that of $\alpha_{0}(X_{i})$ if $D_{i}=0$; that is,
$\widehat{w}_{i}=\begin{cases}\widehat{\alpha}(1,Z_{i})&\text{if }D_{i}=1,\\ -\widehat{\alpha}(0,Z_{i})&\text{if }D_{i}=0.\end{cases}.$
This modeling enforces the correct signs and nonnegativity of the Riesz representer.
Remark (Propensity score modeling).
We can interpret that the Riesz representer model is based on a propensity score model:
$\alpha_{\bm{\beta}}(X)=\mathbbm{1}[D=1]r_{\bm{\beta}}(1,Z)-\mathbbm{1}[D=0]r_{\bm{\beta}}(0,Z),$
where
$r_{\bm{\beta}}(1,Z)$ $=\frac{1}{e_{\bm{\beta}}(Z)},\qquad r_{\bm{\beta}}(0,Z)=\frac{1}{1-e_{\bm{\beta}}(Z)},$
$e_{\bm{\beta}}(Z)$ $\coloneqq\frac{1}{1+\exp\big(-{\bm{\phi}}(Z)^{\top}{\bm{\beta}}\big)},$
and ${\bm{\phi}}\colon{\mathcal{Z}}\to{\mathbb{R}}^{p}$ is a basis function. Under this model, minimizing the UKL flavored empirical Bregman objective yields an estimator that satisfies an automatic covariate balancing property, as discussed in Section 4 and Zhao (2019).
Concretely, letting $\widehat{\alpha}=\alpha_{\widehat{{\bm{\beta}}}}$, we estimate $\bm{\beta}$} by
$\widehat{{\bm{\beta}}}$ $\coloneqq\operatorname*{arg\,min}_{{\bm{\beta}}}\frac{1}{n}\sum_{i=1}^{n}\Bigg(\mathbbm{1}[D_{i}=1]\left(-\log\left(\frac{1}{r_{\bm{\beta}}(1,Z_{i})-1}\right)+r_{\bm{\beta}}(1,Z_{i})\right)$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\mathbbm{1}[D_{i}=0]\left(-\log\left(\frac{1}{r_{\bm{\beta}}(0,Z_{i})-1}\right)+r_{\bm{\beta}}(0,Z_{i})\right)\Bigg)+\frac{1}{a}\lambda\|{\bm{\beta}}\|^{a}_{a}.$
By duality, if $a=1$, this KL divergence objective is equivalent to a covariate balancing program:
$\min_{{\bm{w}}\in(1,\infty)^{n}}\$$\sum_{i=1}^{n}(w_{i}-1)\log(w_{i}-1)$
s.t. $\ \ \left|\sum_{i=1}^{n}\Big(\mathbbm{1}[D_{i}=1]w_{i}\phi_{j}(Z_{i})-\mathbbm{1}[D_{i}=0]w_{i}\phi_{j}(Z_{i})\Big)\right|\leq\lambda\quad\text{for}\ j=1,2,\dots,p,$
where the solution $\widehat{w}_{i}$ corresponds to the estimator of $\alpha_{0}(X_{i})$ if $D_{i}=1$ and that of $\alpha_{0}(X_{i})$ if $D_{i}=0$; that is,
$\widehat{w}_{i}=\begin{cases}\widehat{r}(1,Z_{i})&\text{if }D_{i}=1,\\ \widehat{r}(0,Z_{i})&\text{if }D_{i}=0.\end{cases},$
and $\widehat{r}$ is an estimator of the density ratio $r_{0}$. This constrained optimization matches entropy balancing (Hainmueller, 2012). In particular, when $\lambda=0$ we obtain exact balance,
$\sum_{i=1}^{n}\Big(\mathbbm{1}[D_{i}=1]\widehat{w}_{i}\phi_{j}(Z_{i})-\mathbbm{1}[D_{i}=0]\widehat{w}_{i}\phi_{j}(Z_{i})\Big)=0\quad\text{for}\ j=1,2,\dots,p.$
This specification has the practical advantage that $\phi(Z)$ can be chosen independently of $D$.
BP-Riesz Regression.
BP-Riesz regression uses Basu’s power divergence with $C=1$ and $\delta\in(0,\infty)$:
$g^{\text{BP}}(\alpha)\coloneqq\frac{\big(|\alpha|-1\big)^{1+\delta}-\big(|\alpha|-1\big)}{\delta}-|\alpha|.$
Plugging $g^{\text{BP}}$ into (1) and using $m^{\text{ATE}}$ yields the empirical objective
$\widehat{\text{BD}}_{g^{\text{BP}}}(\alpha)$ $\coloneqq\frac{1}{n}\sum_{i=1}^{n}\Bigg(\frac{\big(|\alpha(D_{i},Z_{i})|-1\big)^{\delta}-1}{\delta}+|\alpha(D_{i},Z_{i})|\big(|\alpha(D_{i},Z_{i})|-1\big)^{\delta}$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ -\upsilon\Big(\big(|\alpha(1,X)|-1\big)^{\delta}-\big(|\alpha(0,X)|-1\big)^{\delta}\Big)\Bigg),$
where $\upsilon\coloneqq 1+1/\delta$. We then estimate $\alpha^{\text{ATE}}_{0}$ by
$\widehat{\alpha}\coloneqq\operatorname*{arg\,min}_{\alpha\in{\mathcal{A}}}\widehat{\text{BD}}_{g^{\text{BP}}}(\alpha)+\lambda J(\alpha).$
Remark (Power link function).
A convenient parametric specification that is consistent with Section 4 models $\alpha$ through a power link function for the inverse propensity components:
$\alpha_{\bm{\beta}}(X)=\mathbbm{1}[D=1]r_{\bm{\beta}}(1,Z)-\mathbbm{1}[D=0]r_{\bm{\beta}}(0,Z),$
with
$r_{\bm{\beta}}(1,Z)\coloneqq 1+\left(1+\frac{{\bm{\phi}}(1,Z)^{\top}{\bm{\beta}}}{\upsilon}\right)^{1/\delta},\qquad r_{\bm{\beta}}(0,Z)\coloneqq 1+\left(1-\frac{{\bm{\phi}}(0,Z)^{\top}{\bm{\beta}}}{\upsilon}\right)^{1/\delta},$
on the domain where the above powers are well defined. This specification interpolates between the squared distance and UKL divergence: $\delta=1$ recovers SQ-Riesz regression, and the limit $\delta\to 0$ recovers UKL-Riesz regression, as discussed in Section 3. In applications, BP-Riesz regression can mitigate sensitivity to extreme inverse propensity weights while retaining the covariate balancing behavior implied by the dual characterization.
BKL-Riesz Regression.
Consider Riesz representer models $\alpha$ such that $\alpha(1,x)>1$ and $\alpha(0,x)<-1$ for all $x$. BKL-Riesz regression uses BKL divergence with $C=1$:
$g^{\text{BKL}}(\alpha)\coloneqq(|\alpha|-1)\log\big(|\alpha|-1\big)-(|\alpha|+1)\log(|\alpha|+1).$
By plugging $g^{\text{BKL}}$ into (1) and using $m^{\text{ATE}}$, we have the following empirical objective function:
$\widehat{\text{BD}}_{g^{\text{BKL}}}(\alpha)$ $\coloneqq\frac{1}{n}\sum^{n}_{i=1}\Big(\log\left(\frac{|\alpha(X_{i})|-1}{|\alpha(X_{i})|+1}\right)$
$\ \ \ \ \ \ -\Big(\log\left(\frac{\alpha(1,Z_{i})-1}{\alpha(1,Z_{i})+1}\right)+\log\left(\frac{-\alpha(0,Z_{i})-1}{-\alpha(0,Z_{i})+1}\right)\Big)\Big).$
We then estimate $\alpha^{\text{ATE}}_{0}$ by
$\widehat{\alpha}\coloneqq\operatorname*{arg\,min}_{\alpha\in{\mathcal{A}}}\widehat{\text{BD}}_{g^{\text{BKL}}}(\alpha)+\lambda J(\alpha).$
Remark (MLE of the propensity score).
BKL-Riesz regression corresponds to estimating the propensity score by regularized logistic likelihood, which is the standard MLE approach in ATE estimation. Let
$e_{\bm{\beta}}(Z)\coloneqq\frac{1}{1+\exp\big(-{\bm{\phi}}(Z)^{\top}{\bm{\beta}}\big)},$
and define the Riesz representer model obtained by plugging in $e_{\bm{\beta}}$,
$\alpha_{\bm{\beta}}(X)\coloneqq\frac{D}{e_{\bm{\beta}}(Z)}-\frac{1-D}{1-e_{\bm{\beta}}(Z)}.$
Under the BKL choice in Section 3, minimizing the corresponding empirical Bregman divergence specializes to minimizing the Bernoulli negative log-likelihood:
$\widehat{{\bm{\beta}}}\coloneqq\operatorname*{arg\,min}_{{\bm{\beta}}}-\frac{1}{n}\sum_{i=1}^{n}\Big(D_{i}\log e_{\bm{\beta}}(Z_{i})+(1-D_{i})\log\big(1-e_{\bm{\beta}}(Z_{i})\big)\Big)+\lambda\|{\bm{\beta}}\|_{2}^{2},$
and we set $\widehat{e}(Z)\coloneqq e_{\widehat{{\bm{\beta}}}}(Z)$ and
$\widehat{\alpha}(X)\coloneqq\frac{D}{\widehat{e}(Z)}-\frac{1-D}{1-\widehat{e}(Z)}.$
This viewpoint aligns with the interpretation of BKL-Riesz as a probabilistic classification approach to density ratio estimation (Qin, 1998; Cheng & Chu, 2004), here applied to treatment assignment modeling. It also provides a baseline for comparison with the direct objectives in SQ-Riesz, UKL-Riesz, and BP-Riesz regression.
7.2 AME Estimation
We consider the AME setup described in Section 2. Let $X=(D,Z)$, where $D$ is a scalar continuous regressor and $Z$ is a vector of covariates. The target parameter is
$\theta^{\text{AME}}_{0}\coloneqq\mathbb{E}\Big[\partial_{d}\gamma_{0}(D,Z)\Big],\qquad m^{\text{AME}}(W,\gamma)\coloneqq\partial_{d}\gamma(D,Z).$
Assume that $X$ admits a density $f_{0}$ that is continuously differentiable and that an integration by parts argument is valid, for example, $\gamma(x)f_{0}(x)$ vanishes on the boundary of the support in the $d$ direction. Then
$\mathbb{E}\Big[\partial_{d}\gamma(X)\Big]=\mathbb{E}\Big[\alpha^{\text{AME}}_{0}(X)\gamma(X)\Big],\qquad\alpha^{\text{AME}}_{0}(X)=-\partial_{d}\log f_{0}(X),$
so the AME Riesz representer is the negative score of the marginal density of $X$ with respect to $d$. Since $\partial_{d}\log f_{0}(D,Z)=\partial_{d}\log f_{0}(D\mid Z)$, we can equivalently view $\alpha^{\text{AME}}_{0}$ as the negative score of the conditional density of $D$ given $Z$.
To estimate $\alpha^{\text{AME}}_{0}$, we apply generalized Riesz regression with $m=m^{\text{AME}}$. The population objective in Section 3 becomes
$\text{BD}^{\text{AME}}_{g}(\alpha)\coloneqq\mathbb{E}\Big[-g\big(\alpha(X)\big)+\partial g\big(\alpha(X)\big)\alpha(X)-\partial_{d}\Big(\partial g\big(\alpha(X)\big)\Big)\Big],$
and we minimize its empirical analogue over a differentiable model class $\mathcal{A}$} (so that $\partial_{d}\alpha(X)$ and $\partial_{d}\{\partial g(\alpha(X))\}$ are well defined), possibly with regularization.
SQ-Riesz Regression.
Let $g^{\text{SQ}}(\alpha)=(\alpha-C)^{2}$ for an arbitrary constant $C\in{\mathbb{R}}$, so that $\partial g^{\text{SQ}}(\alpha)=2(\alpha-C)$. Substituting into $\text{BD}^{\text{AME}}_{g}$ yields
$\text{BD}^{\text{AME}}_{g^{\text{SQ}}}(\alpha)$ $=\mathbb{E}\Big[-\big(\alpha(X)-C\big)^{2}+2\big(\alpha(X)-C\big)\alpha(X)-\partial_{d}\Big(2\big(\alpha(X)-C\big)\Big)\Big]$
$=\mathbb{E}\Big[\alpha(X)^{2}-2\partial_{d}\alpha(X)\Big]+\text{const},$
where the constant does not depend on $\alpha$. Hence SQ-Riesz regression targets $\alpha^{\text{AME}}_{0}$ in $L_{2}$. This objective is also a score matching style criterion for estimating the score, written here in terms of the negative score $\alpha^{\text{AME}}_{0}$. The empirical estimator is
$\widehat{\alpha}\coloneqq\operatorname*{arg\,min}_{\alpha\in{\mathcal{A}}}\frac{1}{n}\sum_{i=1}^{n}\Big(\alpha(X_{i})^{2}-2\partial_{d}\alpha(X_{i})\Big)+\lambda J(\alpha).$
This method corresponds to Riesz regression for AME, as discussed in Chernozhukov et al. (2021).
UKL-Riesz Regression.
To obtain a KL motivated loss that allows signed $\alpha$, we use the signed KL type convex function
$g^{\text{UKL}}(\alpha)=|\alpha|\log|\alpha|-|\alpha|,\qquad\partial g^{\text{UKL}}(\alpha)=\operatorname{sign}(\alpha)\log|\alpha|,$
on a domain that excludes $\alpha=0$. Plugging into $\text{BD}^{\text{AME}}_{g}$ gives
$\text{BD}^{\text{AME}}_{g^{\text{UKL}}}(\alpha)$ $=\mathbb{E}\Big[-|\alpha(X)|\log|\alpha(X)|+|\alpha(X)|+\operatorname{sign}\big(\alpha(X)\big)\log|\alpha(X)|\alpha(X)-\partial_{d}\Big(\operatorname{sign}\big(\alpha(X)\big)\log|\alpha(X)|\Big)\Big]$
$=\mathbb{E}\Big[|\alpha(X)|-\partial_{d}\Big(\operatorname{sign}\big(\alpha(X)\big)\log|\alpha(X)|\Big)\Big]+\text{const}.$
Accordingly, we estimate $\alpha^{\text{AME}}_{0}$ by minimizing the empirical version over $\mathcal{A}$}:
$\widehat{\alpha}\coloneqq\operatorname*{arg\,min}_{\alpha\in{\mathcal{A}}}\frac{1}{n}\sum_{i=1}^{n}\Big(|\alpha(X_{i})|-\partial_{d}\Big(\operatorname{sign}\big(\alpha(X_{i})\big)\log|\alpha(X_{i})|\Big)\Big)+\lambda J(\alpha).$
In practice, one can use the shifted UKL loss in Section 3 to avoid the singularity at zero and combine it with a branchwise link specification as in Section 4.
BP-Riesz Regression.
BP-Riesz regression interpolates between squared distance and UKL divergence. For simplicity, we present the unshifted form with $C=0$:
$g^{\text{BP}}(\alpha)\coloneqq\frac{|\alpha|^{1+\gamma}-|\alpha|}{\gamma}-|\alpha|,\qquad\partial g^{\text{BP}}(\alpha)=\left(1+\frac{1}{\gamma}\right)\operatorname{sign}(\alpha)\Big(|\alpha|^{\gamma}-1\Big),\qquad\gamma\in(0,\infty).$
Let $k\coloneqq 1+1/\gamma$. Then $\text{BD}^{\text{AME}}_{g}$ simplifies to
$\text{BD}^{\text{AME}}_{g^{\text{BP}}}(\alpha)=\mathbb{E}\Big[|\alpha(X)|^{1+\gamma}-\partial_{d}\Big(k\operatorname{sign}\big(\alpha(X)\big)\Big(|\alpha(X)|^{\gamma}-1\Big)\Big)\Big]+\text{const}.$
We estimate $\alpha^{\text{AME}}_{0}$ by minimizing the empirical objective:
$\widehat{\alpha}\coloneqq\operatorname*{arg\,min}_{\alpha\in{\mathcal{A}}}\frac{1}{n}\sum_{i=1}^{n}\Big(|\alpha(X_{i})|^{1+\gamma}-\partial_{d}\Big(k\operatorname{sign}\big(\alpha(X_{i})\big)\Big(|\alpha(X_{i})|^{\gamma}-1\Big)\Big)\Big)+\lambda J(\alpha).$
As in Section 3, $\gamma=1$ recovers the squared loss behavior (up to scaling), while $\gamma\to 0$ approaches a KL type criterion through the identity $\lim_{\gamma\to 0}(|\alpha|^{\gamma}-1)/\gamma=\log|\alpha|$.
BKL-Riesz Regression.
Finally, we can use the BKL loss from Section 3 to obtain a logistic motivated criterion. Let $C>0$ and define
$g^{\text{BKL}}(\alpha)\coloneqq(|\alpha|-C)\log\big(|\alpha|-C\big)-(|\alpha|+C)\log\big(|\alpha|+C\big),\qquad\partial g^{\text{BKL}}(\alpha)=\operatorname{sign}(\alpha)\log\left(\frac{|\alpha|-C}{|\alpha|+C}\right).$
Then the AME objective is
$\text{BD}^{\text{AME}}_{g^{\text{BKL}}}(\alpha)={\mathbb{E}}\left[C\log\left(\frac{|\alpha(X)|-C}{|\alpha(X)|+C}\right)-\partial_{d}\left(\operatorname{sign}\big(\alpha(X)\big)\log\left(\frac{|\alpha(X)|-C}{|\alpha(X)|+C}\right)\right)\right]+\text{const},$
and the estimator minimizes its empirical counterpart over $\mathcal{A}$} with regularization. As in the ATE case, this loss is naturally paired with a logistic style link for the magnitude of $\alpha$, while sign changes can be handled via the branchwise constructions in Section 4.
Once we obtain $\widehat{\alpha}^{\text{AME}}$ and an outcome regression estimator $\widehat{\gamma}$, we plug them into the Neyman orthogonal score in Section 2 to form an estimator of $\theta_{0}^{\text{AME}}$.
7.3 Covariate Shift Adaptation (Density Ratio Estimation)
We consider the covariate shift setting in Section 2. Let $X$ be the source covariate distribution that generates labeled observations $\{(X_{i},Y_{i})\}^{n}_{i=1}$, and let $\widetilde{X}$ be the target covariate distribution that generates unlabeled observations $\{\widetilde{X}_{j}\}^{m}_{j=1}$, independent of the source sample. Let $p_{0}(x)$ and $p_{1}(x)$ be the pdfs of $X$ and $\widetilde{X}$, respectively. We assume that $p_{0}(x),p_{1}(x)>0$ for all $x\in{\mathcal{X}}$. The Riesz representer for covariate shift adaptation is the density ratio
$\alpha^{\text{CS}}_{0}(X)=r_{0}(X)\coloneqq\frac{p_{1}(X)}{p_{0}(X)}.$
We estimate $r_{0}$ directly by density ratio fitting under a Bregman divergence, avoiding separate density estimation for $p_{0}(X)$ and $p_{1}(X)$.
Let $g\colon{\mathbb{R}}_{+}\to{\mathbb{R}}$ be differentiable and strictly convex. The Bregman divergence between $r_{0}$ and a candidate ratio model $\alpha$ is
$\text{BD}^{\dagger}_{g}\big(r_{0}\mid\alpha\big)\coloneqq{\mathbb{E}}_{X}\Big(g\big(r_{0}(X)\big)-g\big(\alpha(X)\big)-\partial g\big(\alpha(X)\big)\big(r_{0}(X)-\alpha(X)\big)\Big).$
Dropping the constant ${\mathbb{E}}_{X}\big[g(r_{0}(X))\big]$ and using the identity ${\mathbb{E}}_{X}\big[r_{0}(X)h(X)\big]={\mathbb{E}}_{\widetilde{X}}\big[h(X)\big]$, we obtain the equivalent population objective
$\text{BD}^{\text{CS}}_{g}(\alpha)$ $\coloneqq{\mathbb{E}}_{X}\big[\partial g(\alpha(X))\alpha(X)-g(\alpha(X))\big]-{\mathbb{E}}_{\widetilde{X}}\big[\partial g(\alpha(X))\big].$
Given samples $\{X_{i}\}_{i\in{\mathcal{I}}_{S}}$ and $\{\widetilde{X}_{j}\}_{j\in{\mathcal{I}}_{T}}$, the empirical objective is
$\widehat{\text{BD}}^{\text{CS}}_{g}(\alpha)$ $\coloneqq\frac{1}{n}\sum^{n}_{i=1}\Big(\partial g\big(\alpha(X_{i})\big)\alpha(X_{i})-g\big(\alpha(X_{i})\big)\Big)-\frac{1}{m}\sum^{m}_{j=1}\partial g\big(\alpha(\widetilde{X}_{j})\big).$ (7)
We estimate the density ratio by
$\widehat{\alpha}\coloneqq\operatorname*{arg\,min}_{\alpha\in{\mathcal{A}}}\widehat{\text{BD}}^{\text{CS}}_{g}(\alpha)+\lambda J(\alpha),$
where $\mathcal{A}$} is a model class and $J$ is a regularizer. A convenient way to enforce $\alpha(x)\geq 0$ is to use a link specification such as $\alpha(x)=\exp\big(f(x)\big)$ with a flexible regression model $f$.
SQ-Riesz Regression.
For the squared loss, take
$g^{\text{SQ}}(\alpha)=(\alpha-1)^{2},\qquad\partial g^{\text{SQ}}(\alpha)=2(\alpha-1).$
Substituting into (7) and dropping constants that do not depend on $\alpha$, we obtain
$\widehat{\text{BD}}^{\text{CS}}_{g^{\text{SQ}}}(\alpha)=\frac{1}{n}\sum^{n}_{i=1}\alpha(X_{i})^{2}-\frac{2}{m}\sum^{m}_{j=1}\alpha(\widetilde{X}_{j}).$
This is the classical least-squares importance fitting (LSIF) objective in density ratio estimation (Kanamori et al., 2009). In the debiased machine learning literature, the same squared loss criterion is also used as Riesz regression for covariate shift adaptation (Chernozhukov et al., 2025). Related extensions include doubly robust covariate shift adaptation schemes that combine density ratio estimation with regression adjustment (Kato et al., 2024a).
UKL-Riesz Regression.
For a KL motivated objective on $\mathbb{R}$$_{+}$, take
$g^{\text{UKL}}(\alpha)=\alpha\log\alpha-\alpha,\qquad\partial g^{\text{UKL}}(\alpha)=\log\alpha.$
Then (7) becomes, up to an additive constant that does not depend on $\alpha$,
$\widehat{\text{BD}}^{\text{CS}}_{g^{\text{UKL}}}(\alpha)=\frac{1}{n}\sum^{n}_{i=1}\alpha(X_{i})-\frac{1}{m}\sum^{m}_{j=1}\log\alpha(\widetilde{X}_{j}).$
A standard implementation imposes the normalization constraint $\frac{1}{n}\sum^{n}_{i=1}\alpha(X_{i})=1$, in which case minimizing $\widehat{\text{BD}}^{\text{CS}}_{g^{\text{UKL}}}$ is equivalent to maximizing the target log-likelihood $\frac{1}{m}\sum^{m}_{j=1}\log\alpha(\widetilde{X}_{j})$ subject to normalization and nonnegativity, which yields KLIEP style procedures (Sugiyama et al., 2008). This constrained view is also useful for understanding the dual characterization and the associated moment matching property.
BP-Riesz Regression.
BP-Riesz regression interpolates between squared loss and KL type objectives. For $\gamma\in(0,\infty)$, consider the BP choice on $\mathbb{R}$$_{+}$ with $C=0$,
$g^{\text{BP}}(\alpha)\coloneqq\frac{\alpha^{1+\gamma}-\alpha}{\gamma}-\alpha,\qquad\partial g^{\text{BP}}(\alpha)=\left(1+\frac{1}{\gamma}\right)\Big(\alpha^{\gamma}-1\Big).$
A useful simplification is that $\partial g^{\text{BP}}(\alpha)\alpha-g^{\text{BP}}(\alpha)=\alpha^{1+\gamma}$, so (7) reduces, up to constants, to
$\widehat{\text{BD}}^{\text{CS}}_{g^{\text{BP}}}(\alpha)=\frac{1}{n}\sum^{n}_{i=1}\alpha(X_{i})^{1+\gamma}-\left(1+\frac{1}{\gamma}\right)\frac{1}{m}\sum^{m}_{j=1}\alpha(\widetilde{X}_{j})^{\gamma}.$
When $\gamma=1$, this objective coincides with the SQ-Riesz objective, up to scaling and constants. As $\gamma\to 0$, it approaches a KL flavored criterion via the expansion $\alpha^{\gamma}=1+\gamma\log\alpha+o(\gamma)$, providing a continuous bridge between LSIF and KLIEP, and offering a robustness device against extreme ratios (Basu et al., 1998; Sugiyama et al., 2012).
BKL-Riesz Regression.
BKL-Riesz regression corresponds to probabilistic classification based density ratio estimation, which estimates the log density ratio by fitting a classifier that discriminates target covariates from source covariates (Qin, 1998; Cheng & Chu, 2004). Let $S\in\{0,1\}$ denote a domain indicator, where $S=1$ for target and $S=0$ for source, and let $\pi\coloneqq P(S=1)$ denote the mixture class prior. Under Bayes’ rule, the density ratio satisfies
$r_{0}(x)=\frac{p_{1}(x)}{p_{0}(x)}=\frac{1-\pi}{\pi}\frac{P(S=1\mid X=x)}{P(S=0\mid X=x)}.$
We model $P(S=1\mid X=x)$ by a logistic specification
$p_{\bm{\beta}}(S=1\mid X=x)\coloneqq\frac{1}{1+\exp\big(-{\bm{\phi}}(x)^{\top}{\bm{\beta}}\big)},$
and estimate $\bm{\beta}$} by regularized Bernoulli likelihood on the pooled sample:
$\widehat{{\bm{\beta}}}\coloneqq\operatorname*{arg\,min}_{{\bm{\beta}}}-\frac{1}{n+m}\left(\sum^{n}_{i=1}\log\big(1-p_{\bm{\beta}}(S=1\mid X_{i})\big)+\sum^{m}_{j=1}\log p_{\bm{\beta}}(S=1\mid\widetilde{X}_{j})\right)+\lambda\|{\bm{\beta}}\|_{2}^{2}.$
With $\widehat{\pi}\coloneqq\frac{m}{n+m}$, we then set
[…]
This construction enforces nonnegativity by design and connects density ratio estimation to standard classification tools.
Remark (From density ratio estimation to covariate shift adaptation).
Once we obtain $\widehat{\alpha}$ and an outcome regression estimator $\widehat{\gamma}$, we plug them into the covariate shift Neyman score in Section 2. In particular, a doubly robust estimator that accommodates separate source and target samples is
$\widehat{\theta}^{\text{CS}}\coloneqq\frac{1}{m}\sum^{m}_{j=1}\widehat{\gamma}(\widetilde{X}_{j})+\frac{1}{n}\sum^{n}_{i=1}\widehat{\alpha}(X_{i})\Big(Y_{i}-\widehat{\gamma}(X_{i})\Big).$
The corresponding IPW estimator is obtained by dropping the regression adjustment term and using $\widehat{\theta}^{\text{CS}}_{\text{IPW}}=\frac{1}{n}\sum^{n}_{i=1}\widehat{\alpha}(X_{i})Y_{i}$. Cross fitting can be applied by estimating $\widehat{\alpha}$ and $\widehat{\gamma}$ on auxiliary folds and evaluating the above scores on held out folds.
8 Overfitting Problems
Density ratio estimation often suffers from a characteristic form of overfitting. Kato & Teshima (2021) refers to this issue as train-loss hacking and shows that the empirical objective can be artificially reduced by inflating $r(X^{(\text{nu})})$ at the training points. Intuitively, the term $\hat{{\mathbb{E}}}_{\text{nu}}[\partial f(r(X))]$ decreases as $r$ increases on $\{X^{(\text{nu})}_{j}\}$, so an optimizer can exploit this monotonicity, which can lead to divergence or saturation at the output bounds. Rhodes et al. (2020) highlights a related mechanism: when $p_{\text{nu}}$ and $p_{\text{de}}$ are far apart, for example when $\mathrm{KL}(p_{\text{nu}}\|p_{\text{de}})$ is on the order of tens of nats, the estimation problem enters a large-gap regime that exacerbates overfitting. They refer to this phenomenon as the density chasm. Although the two papers emphasize different viewpoints, both point to the same underlying difficulty, finite samples provide weak control of the ratio in regions where the two distributions have little overlap.
Non-negative Bregman divergence
Kato & Teshima (2021) proposes a modification of the Bregman divergence objective that isolates the problematic component and applies a non-negative correction under a mild boundedness condition on $r_{0}$. Specifically, choose $0<C<1/R$ with $R:=\sup r_{0}$. The population objective decomposes, up to an additive constant, into a non-negative term plus a bounded residual. At the sample level, the method replaces the non-negative component with its positive part $\cdot$$_{+}$. This yields an objective that curbs train-loss hacking while remaining within the Bregman-divergence framework (Kiryo et al., 2017; Kato & Teshima, 2021).
Telescoping density ratio estimation
Rhodes et al. (2020) proposes telescoping density ratio estimation, which targets overfitting in large-gap regimes by introducing intermediate waymark distributions $p_{0}=p_{\text{nu}},p_{1},\dots,p_{m}=p_{\text{de}}$. The method estimates local ratios $p_{k}/p_{k+1}$ and combines them through the identity
$\frac{p_{0}(x)}{p_{m}(x)}=\prod_{k=0}^{m-1}\frac{p_{k}(x)}{p_{k+1}(x)}.$
Each local ratio corresponds to a smaller distributional gap, which makes perfect classification harder and typically makes the ratio estimation problem more stable at finite sample sizes. As a result, telescoping can improve robustness and generalization in practice.
Telescoping density ratio estimation is also closely connected to score matching. When the number of intermediate ratios tends to infinity, the log density ratio can be expressed as an integral of time scores along a continuum of bridge distributions, and can be approximated by aggregating these score functions (Choi et al., 2022). Building on this idea, Choi et al. (2022) proposes density ratio estimation via infinitesimal classification. See Appendix H.4 for details.
Table 2: Experimental results using the synthetic dataset. We report the mean squared error (MSE) of the ATE estimator and the empirical coverage ratio (CR) of nominal 9595% Wald-type confidence intervals over 100 Monte Carlo replications. The column block “True” uses oracle nuisance functions (the true propensity score and true outcome regressions) and is therefore infeasible. “SQ-Riesz” and “UKL-Riesz” estimate the ATE Riesz representer by generalized Riesz regression with the squared-loss and unnormalized-KL objectives, respectively; “BKL-Riesz = MLE” corresponds to estimating the propensity score by BKL objective (logistic MLE) and plugging it into the ATE Riesz representer. For SQ-Riesz we compare two link specifications (Linear and Logit). For UKL-Riesz we compare two feature sets: ${\bm{\phi}}(Z)$ uses only covariates $Z$, while ${\bm{\phi}}(X)$ uses the full regressor $X=(D,Z)$ (allowing treatment-dependent features). For each Riesz-representer fit we report three estimators: the direct method (DM), inverse probability weighting (IPW), and augmented IPW (AIPW). DM depends only on the outcome regression; small differences across DM columns arise from recomputing the outcome regression within each run of the AIPW construction. Values of CR close to $0.95$ indicate well-calibrated uncertainty quantification.
True SQ-Riesz (Linear) SQ-Riesz (Logit) UKL-Riesz (${\bm{\phi}}(Z)$) UKL-Riesz (${\bm{\phi}}(X)$) BKL-Riesz (MLE)
DM IPW AIPW DM IPW AIPW DM IPW AIPW DM IPW AIPW DM IPW AIPW DM IPW AIPW
MSE 0.00 1.44 0.01 0.39 0.49 0.19 0.39 1.38 0.08 0.38 1.50 0.10 0.40 1.52 0.10 0.39 3.79 0.23
CR 1.00 0.84 0.98 0.06 0.98 0.87 0.12 0.80 0.89 0.08 0.73 0.77 0.06 0.68 0.81 0.06 0.32 0.60
9 Experiments
We evaluate generalized Riesz regression as a building block for debiased machine learning, focusing on average treatment effect (ATE) estimation. Across all experiments, we compare three ways of estimating the ATE Riesz representer (bias-correction term) introduced in Section 3: SQ-Riesz (squared-loss objective), UKL-Riesz (unnormalized-KL objective), and BKL-Riesz (binary-KL objective). In the ATE setting, BKL-Riesz coincides with estimating the propensity score by Bernoulli likelihood (logistic MLE) and then plugging it into the closed-form ATE Riesz representer; we therefore refer to it as “BKL-Riesz = MLE.”
Given an estimate of the outcome regression $\widehat{\gamma}$ and an estimate of the Riesz representer $\widehat{\alpha}$, we report three ATE estimators:
  • DM: the plug-in direct method based only on $\widehat{\gamma}$,
  • IPW: the IPW estimator based only on $\widehat{\alpha}$,
  • AIPW: the Neyman-orthogonal (doubly robust) estimator combining $\widehat{\gamma}$ and $\widehat{\alpha}$ as in Section 2.
We quantify accuracy by the mean squared error (MSE) of the ATE estimate and quantify uncertainty by the empirical coverage ratio (CR) of nominal $95$
9.1 Experiments with synthetic dataset
Design
The covariates are three-dimensional, $K=3$, and we fix the sample size at $n=3000$. In each Monte Carlo replication, we generate covariates $X_{i}\in{\mathbb{R}}^{3}$ from a multivariate normal distribution $\mathcal{N}(0,I_{3})$ and construct a nonlinear propensity score model with polynomial and interaction terms as
$e_{0}(X_{i})=\frac{1}{1+\exp\bigl(-h(X_{i})\bigr)},$
where
$h(X_{i})=\sum_{j=1}^{3}a_{j}X_{i,j}+\sum_{j=1}^{3}b_{j}X_{i,j}^{2}+c_{1}X_{i,1}X_{i,2}+c_{2}X_{i,2}X_{i,3}+c_{3}X_{i,1}X_{i,3}.$
The coefficients $a_{j}$, $b_{j}$, and $c_{j}$ are independently drawn from $\mathcal{N}(0,0.5)$. Given these propensity scores, the treatment assignment $D_{i}$ is sampled accordingly. We then generate the outcome as
$Y_{i}=1.0+\left(\sum^{3}_{j=1}X_{i,j}\widetilde{a}_{j}\right)^{2}+1/\left(1+\exp\left(-\left(\sum^{3}_{j=1}X^{2}_{i,j}\widetilde{b}_{j}\right)\right)\right)+\tau_{0}D_{i}+\varepsilon_{i},$
where $\varepsilon_{i}\sim\mathcal{N}(0,1)$ and $\tau_{0}=5.0$.
Estimators and implementation
We estimate the Riesz representer using the following variants, matched to Table 2.
  • SQ-Riesz (Linear) and SQ-Riesz (Logit): squared-loss generalized Riesz regression with two different link specifications for the Riesz-representer model.
  • UKL-Riesz (${\bm{\phi}}(Z)$) and UKL-Riesz (${\bm{\phi}}(X)$): UKL generalized Riesz regression with a log-type link, comparing two feature sets. Here $X=(D,Z)$ and ${\bm{\phi}}(Z)$ uses only $Z$, while ${\bm{\phi}}(X)$ uses the full regressor (allowing treatment-dependent features).
  • BKL-Riesz (MLE): propensity-score MLE (Bernoulli likelihood) followed by plugging $\widehat{e}(Z)$ into the ATE Riesz representer.
For the Riesz representer and regression models, we separately use a neural network with one hidden layer consisting of $100$ nodes. To avoid relying on the Donsker condition, we estimate all nuisance functions using two-fold cross fitting. In each replication, we split the sample into two folds, estimate the nuisance functions on one fold, evaluate the corresponding scores on the other fold, and then swap the roles of the folds. The final estimators aggregate the two cross-fitted scores.
This experiment does not guarantee automatic Neyman orthogonalization, since we use cross fitting and do not use the same basis functions for outcome modeling. However, this implementation is standard in debiased machine learning; therefore, we adopt it.
We repeat the simulation 100 times. The “True” columns in Table 2 report infeasible oracle performance using the true nuisance functions.
Results
Table 2 highlights three robust patterns. First, oracle baselines separate estimation error from intrinsic variance. The oracle AIPW estimator is close to the efficiency benchmark (MSE =0.01=0.01) and achieves near-nominal coverage (CR =0.98=0.98). In contrast, even with the true propensity score, oracle IPW remains noisy (MSE =1.44=1.44) and undercovers (CR =0.84=0.84), reflecting the well-known finite-sample instability of pure weighting in challenging overlap regimes.
Second, the plug-in DM estimator is not reliable for inference in this design. Across feasible implementations, DM has moderate MSE (about 0.380.380.400.40) but extremely poor coverage (CR =0.06=0.060.120.12). This indicates that the outcome regression learner, while not catastrophically inaccurate in MSE, does not deliver a reliable uncertainty estimate when used without orthogonalization, and the resulting Wald intervals are severely miscalibrated.
Third, how we fit the Riesz representer matters substantially for IPW, and AIPW mitigates (but does not eliminate) this sensitivity. For IPW, SQ-Riesz (Linear) is the best-performing option in Table 2 (MSE =0.49=0.49) and yields near-nominal coverage (CR =0.98=0.98). In contrast, IPW based on UKL-Riesz has larger MSE (about $1.50$) and noticeably lower coverage (CR =0.68=0.680.730.73), while BKL-Riesz (= MLE) performs worst (MSE =3.79=3.79, CR =0.32=0.32), consistent with propensity-score MLE producing more extreme effective weights in this design.
The AIPW estimator is uniformly more stable than IPW and DM in terms of MSE, but calibration still depends on the Riesz-representer fit. SQ-Riesz (Logit) attains the best AIPW MSE (0.08) with CR 0.89. UKL-Riesz achieves similarly small AIPW MSE (0.10) but exhibits undercoverage (CR =0.77=0.770.810.81). BKL-Riesz (= MLE) improves substantially over its IPW counterpart (AIPW MSE =0.23=0.23), yet its coverage remains poor (CR =0.60=0.60). Overall, directly fitting the Riesz representer via generalized Riesz regression can materially improve finite-sample performance relative to the MLE plug-in baseline, and the combination of objective and link specification plays a first-order role, especially for IPW and for the calibration of AIPW intervals.
Table 3: Experimental results using the the semi-synthetic IHDP dataset. We report the mean squared error (MSE) and the empirical coverage ratio (CR) of nominal 9595% confidence intervals over 1000 replications for the direct method (DM), inverse probability weighting (IPW), and augmented IPW (AIPW) estimators. Nuisance functions are estimated either by a neural network with one hidden layer of size 100 or by an RKHS regression with 100 Gaussian basis functions. The columns correspond to SQ-Riesz regression (SQ-Riesz), UKL-Riesz regression (UKL-Riesz), and BKL-Riesz regression (BKL-Riesz (MLE)). BKL-Riesz (MLE) implies BKL-Riesz regression is essentially equivalent to MLE of logistic models for the propensity score.
Neural network
SQ-Riesz UKL-Riesz BKL-Riesz (MLE)
DM IPW AIPW DM IPW AIPW DM IPW AIPW
MSE 1.52 6.82 0.31 1.55 2.84 0.32 1.58 3.00 0.43
CR 0.03 0.41 1.00 0.03 0.73 0.94 0.01 0.61 0.90
RKHS
SQ-Riesz UKL-Riesz BKL-Riesz (MLE)
DM IPW AIPW DM IPW AIPW DM IPW AIPW
MSE 19.98 3.56 19.97 2.59 1.78 4.45 2.48 1.22 2.32
CR 0.00 0.00 0.00 0.48 0.93 0.88 0.39 0.81 0.84
9.2 Experiments with semi synthetic datasets
We next evaluate the same family of estimators on the semi-synthetic IHDP benchmark, following Chernozhukov et al. (2022a). We use the standard setting “A” in the npci package and generate 1000 replications. Each replication contains $n=747$ observations with a binary treatment, an outcome, and $p=25$ covariates. The estimand is the ATE.
We report DM, IPW, and AIPW for each Riesz-representer estimator (SQ-Riesz, UKL-Riesz, and BKL-Riesz (= MLE)). We consider two nuisance-learner families:
  • a feedforward neural network with one hidden layer of 100 units,
  • an RKHS learner with 100 Gaussian basis functions (with tuning by cross validation).
For each configuration, we compute the MSE of the ATE estimate and the empirical coverage ratio (CR) of nominal 9595% Wald-type confidence intervals across the 1000 replications; CR close to $0.95$ indicates well-calibrated uncertainty quantification. Results appear in Table 3.
Two findings stand out. With neural networks, AIPW is consistently accurate (MSE around 0.310.310.430.43) and well calibrated for UKL-Riesz and BKL-Riesz (CR =0.94=0.94 and $0.90$), while SQ-Riesz yields overly conservative AIPW intervals (CR =1.00=1.00). In contrast, DM has very low coverage (CR near zero) and IPW exhibits large error, especially for SQ-Riesz (MSE =6.82=6.82), reinforcing that orthogonalization is essential in this benchmark.
With RKHS, performance becomes much more sensitive to the particular objective: SQ-Riesz deteriorates sharply (MSE around $20$ with CR =0=0 for both DM and AIPW), whereas UKL-Riesz and BKL-Riesz remain substantially more stable. In particular, UKL-Riesz attains strong IPW calibration under RKHS (CR =0.93=0.93) with comparatively low MSE (1.78), while BKL-Riesz provides a competitive alternative (IPW MSE =1.22=1.22 with CR =0.81=0.81). These results underscore that, in finite samples, the interaction between the Riesz-representer objective and the nuisance-function learner can be decisive, and that UKL-type objectives can offer noticeably more robust behavior than squared-loss fitting in this semi-synthetic setting.
10 Conclusion
This paper develops a unified perspective on estimating the Riesz representer, namely, the bias correction term that appears in Neyman orthogonal scores for a broad class of causal and structural parameters. We formulate Riesz representer estimation as fitting a model to the unknown representer under a Bregman divergence, which yields an empirical risk minimization objective that depends only on observed data. This generalized Riesz regression recovers Riesz regression and least squares importance fitting under squared loss, it recovers KL based tailored loss minimization and its dual entropy balancing weights under a KL type loss, and it connects logistic likelihood based propensity modeling with classification based density ratio estimation through a binary KL criterion. By pairing the loss with an appropriate link function, we make explicit a dual characterization that delivers automatic covariate balancing or moment matching, which clarifies when popular balancing schemes arise as primal or dual solutions. We provide convergence rate results for kernel methods and neural networks, including minimax optimality under standard smoothness classes, and we show how the framework instantiates in ATE, AME, APE, and covariate shift adaptation. Our experiments suggest that directly estimating the bias correction term can be competitive with common propensity score based baselines and can be stable across divergence choices when combined with cross fitting. Overall, the proposed framework bridges density ratio estimation and causal inference, and it offers a single set of tools for designing, analyzing, and implementing Riesz representer estimators, while motivating extensions such as nearest neighbor and score matching based constructions.
Appendix A Related Work
A.1 Asymptotically Efficient Estimation
Early history.
The construction of asymptotically efficient estimators is a classical problem in statistics, machine learning, economics, epidemiology, and related fields. In this problem, we consider semiparametric models with a low-dimensional parameter of interest and additional nuisance parameters. Our interest lies in obtaining $\sqrt{n}$-consistent and asymptotically normal estimators of the parameter of interest. The difficulty stems from the estimation error of nuisance parameters, whose convergence rates are typically slower than, or at best comparable to, $\sqrt{n}$. Reducing the influence of nuisance estimation error has been investigated in many studies, including Levit (1976), Ibragimov & Khas’minskii (1981), Pfanzagl & Wefelmeyer (1982), Klaassen (1987), Robinson (1988), Newey (1994), van der Vaart (1998), Bickel et al. (1998), Ai & Chen (2012), and Chernozhukov et al. (2018).
In the construction of asymptotically efficient estimators, we aim to develop estimators that are $\sqrt{n}$-consistent and asymptotically normal, with asymptotic variances that attain the asymptotic efficiency bounds. Asymptotic efficiency bounds are called H’ajek–Le Cam efficiency bounds, or semiparametric efficiency bounds when we consider semiparametric models (Hájek, 1970; Le Cam, 1972, 1986). They share the same motivation as the Cram’er–Rao lower bound. While the Cram’er–Rao lower bound provides a lower bound for unbiased estimators, asymptotic efficiency bounds provide lower bounds for asymptotically unbiased estimators, called regular estimators. It is known that efficient estimators are regular and asymptotically linear (RAL) with the efficient influence function. Therefore, the construction of asymptotically efficient estimators is equivalent to the construction of RAL estimators (van der Vaart, 1998).
There are three main approaches to constructing efficient estimators, one-step bias correction, estimating equation methods, and TMLE (Schuler et al., 2018; van der Vaart, 2002; van der Laan, 2006; van der Laan & Rose, 2018). In many cases, these approaches yield estimators that are asymptotically equivalent. However, their finite sample behavior may differ. Another related line of work is post-selection inference with high-dimensional control variables (Belloni et al., 2011, 2014, 2016).
Debiased machine learning.
Chernozhukov et al. (2018) proposes debiased machine learning as a framework for constructing asymptotically efficient estimators. This framework formulates the problem through Neyman orthogonality, which is closely connected to the semiparametric efficient influence function (Newey, 1994).
Riesz representer.
The Riesz representer appears as part of the Neyman orthogonal score. It is often called the bias-correction term in one-step bias correction and the clever covariate in TMLE. The terminology “Riesz representer” appears in the debiased machine learning formulation based on Neyman orthogonality (Chernozhukov et al., 2022b).
A.2 ATE Estimation
Randomized controlled trials are the gold standard for causal inference, where treatments are assigned while maintaining balance between treatment groups. However, they are not always feasible, and we aim to estimate causal effects from observational data, where imbalance between treatment groups often arises. To correct this imbalance, propensity scores or balancing weights have been proposed.
In ATE estimation, the Riesz representer corresponds to inverse propensity weights. Accurate estimation of the propensity score is therefore central to ATE estimation. A standard choice is maximum likelihood estimation, but many alternative approaches have been studied. Riesz regression provides an end-to-end approach to estimating the Riesz representer and can be applied to tasks beyond ATE estimation (Chernozhukov et al., 2021, 2022a; Lee & Schuler, 2025). Another promising approach is covariate balancing. The propensity score is also known as the coarsest balancing score (Rosenbaum & Rubin, 1983), and propensity score estimation via covariate balancing has been extensively studied (Li et al., 2017; Imai & Ratkovic, 2013a; Hainmueller, 2012; Zubizarreta, 2015; Tarr & Imai, 2025; Athey et al., 2018). As discussed in this study and related works (Bruns-Smith & Feller, 2022; Bruns-Smith et al., 2025; Ben-Michael et al., 2021; Zhao, 2019), Riesz regression and covariate balancing are dual to each other, in the sense that they correspond to essentially the same optimization problem.
Covariate balancing.
Covariate balancing is a popular approach for propensity score or balancing weight estimation. The propensity score is a balancing score, and based on this property, existing studies propose estimating the propensity score or the weights so that weighted covariate moments match between treated and control groups. Imai & Ratkovic (2013b) proposes estimating the propensity score by matching first moments, and Hazlett (2020) extends this idea to higher-moment matching by mapping covariates into a high-dimensional space via basis functions. On the other hand, methods that do not directly specify a propensity score model have also been proposed. Such methods are called empirical balancing and include entropy balancing (Hainmueller, 2012) and stable weights (Zubizarreta, 2015). These two approaches may appear different, but Zhao (2019) and Bruns-Smith et al. (2025) show that they are essentially equivalent through a duality relationship.
A.3 Density Ratio Estimation
A parallel line of work is density ratio estimation, which has been extensively studied in machine learning. We refer to the ratio between two densities as the density ratio. Density ratio estimation has attracted considerable attention as an essential task in various machine learning problems, such as regression under covariate shift (Shimodaira, 2000; Kato et al., 2024a), learning with noisy labels (Liu & Tao, 2016; Fang et al., 2020), anomaly detection (Smola et al., 2009; Hido et al., 2008; Abe & Sugiyama, 2019), two-sample testing (Keziou & Leoni-Aubin, 2005; Kanamori et al., 2010; Sugiyama et al., 2011a), change-point detection (Kawahara & Sugiyama, 2009), causal inference (Uehara et al., 2020), and recommendation systems (Togashi et al., 2021). While the density ratio can be estimated by separately estimating each density, such an approach may magnify estimation errors by compounding two independent estimations. To address this issue, end-to-end, direct density ratio estimation methods have been studied, including moment matching (Huang et al., 2007; Gretton et al., 2009), probabilistic classification (Qin, 1998; Cheng & Chu, 2004), density matching (Nguyen et al., 2010), density ratio fitting (Kanamori et al., 2009), and PU learning (Kato et al., 2019). It is also known that when complicated models such as neural networks are used for this task, the loss function can diverge in finite samples (Kiryo et al., 2017). Therefore, density ratio estimation methods with neural networks have been investigated (Kato & Teshima, 2021; Rhodes et al., 2020).
Density ratios are used not only for semiparametric analysis but also in tasks such as anomaly detection (Hido et al., 2008). As discussed in this study and in existing work such as Uehara et al. (2020) and Lin et al. (2023), density ratio estimation is closely related to propensity score estimation. In particular, this study shows that the formulations of Riesz regression and LSIF in density ratio estimation are essentially the same. While Riesz regression applies to more general problems, the LSIF literature provides a range of theoretical and empirical results. One important extension is to generalize LSIF via Bregman divergence minimization (Sugiyama, 2006), and this study is strongly inspired by that work.
In causal inference, Uehara et al. (2020) investigates efficient ATE estimation and policy learning under covariate shift. Kato et al. (2024b) applies this approach to adaptive experimental design, and Kato et al. (2025) extends the framework to a PU learning setup.
Appendix B Proof of the Automatic Covariate Balancing Property
B.1 Constrained Optimization Problem
From the Riesz representation theorem, the following equation holds:
${\mathbb{E}}\big[m(W,(\partial g)\circ\alpha)\big]={\mathbb{E}}\big[\alpha_{0}(X)\partial g\big(\alpha(X)\big)\big].$
Therefore, we can consider an algorithm that estimates $\alpha_{0}$ so that its estimator $\widehat{\alpha}$ satisfies
${\mathbb{E}}\Big[m(W,(\partial g)\circ\widehat{\alpha})\Big]\approx{\mathbb{E}}\Big[\widehat{\alpha}(X)\partial g\big(\widehat{\alpha}(X)\big)\Big],$
where we replace $\alpha_{0}$ with $\widehat{\alpha}$. Then, it holds that
$\min_{\bm{\alpha}\in{\mathbb{R}}^{n}}$ $\ \ \ \frac{1}{n}\sum^{n}_{i=1}g(\alpha_{i})$
subjectto $\ \ \ \left|\frac{1}{n}\sum^{n}_{i=1}\Big(\alpha_{i}\partial g\big(\alpha(X_{i})\big)-m(W_{i},(\partial g)\circ\phi_{j})\Big)\right|\leq\lambda\qquad j=1,\dots,p.$
Linearity for the basis functions.
Next, we consider the case where
$\partial g\big(\alpha_{{\bm{\beta}}}(X_{i})\big)=\sum^{p}_{j=1}{\bm{\beta}}_{j}\widetilde{g}\left(X_{i},\phi_{j}(X_{i})\right).$
We consider the following constrained optimization problem:
$\min_{\bm{\alpha}\in{\mathbb{R}}^{n}}$ $\ \ \ \frac{1}{n}\sum^{n}_{i=1}g(\alpha_{i})$
subjectto $\ \ \ \left|\frac{1}{n}\sum^{n}_{i=1}\Big(\alpha_{i}\widetilde{g}\left(X_{i},\phi_{j}(X_{i})\right)-m(W_{i},(\partial g)\circ\phi_{j})\Big)\right|\leq\lambda\qquad j=1,\dots,p.$
B.2 Dual Formulation
Using Lagrange multipliers ${\bm{\beta}}_{j}\in{\mathbb{R}}$ ($j=1,2,\dots,p$), the constrained optimization problem can be written as
$\min_{{\bm{\alpha}}\in{\mathcal{A}}^{n}}\sup_{{\bm{\beta}}\in{\mathbb{R}}^{p}}$ $\ \ \ \left\{\frac{1}{n}\sum^{n}_{i=1}g(\alpha_{i})+\sum^{p}_{j=1}{\bm{\beta}}_{j}\left(\frac{1}{n}\sum^{n}_{i=1}\Big(m(W_{i},(\partial g)\circ\phi_{j})-\alpha_{i}\widetilde{g}\left(X_{i},\phi_{j}(X_{i})\right)\Big)-\operatorname{sign}({\bm{\beta}}_{j})\lambda\right)\right\}.$
The dual problem of the above constrained problem is
$\max_{{\bm{\beta}}\in{\mathbb{R}}^{p}}\inf_{{\bm{\alpha}}\in{\mathcal{A}}^{n}}$ $\ \ \ \left\{\frac{1}{n}\sum^{n}_{i=1}g(\alpha_{i})+\sum^{p}_{j=1}\left({\bm{\beta}}_{j}\frac{1}{n}\sum^{n}_{i=1}\Big(m(W_{i},(\partial g)\circ\phi_{j})-\alpha_{i}\widetilde{g}\left(X_{i},\phi_{j}(X_{i})\right)\Big)-|{\bm{\beta}}_{j}|\lambda\right)\right\}.$
Let $\alpha_{{\bm{\beta}}}(X_{i})={\bm{\phi}}(X_{i})^{\top}{\bm{\beta}}$. Recall that the empirical Bregman divergence objective is given by
$\widehat{\text{BD}}_{g}(\alpha_{{\bm{\beta}}})\coloneqq\frac{1}{n}\sum^{n}_{i=1}\Big(-g(\alpha_{{\bm{\beta}}}(X_{i}))+\partial g(\alpha_{{\bm{\beta}}}(X_{i}))\alpha_{{\bm{\beta}}}(X_{i})-m\big(W_{i},\partial g(\alpha_{{\bm{\beta}}}(X_{i}))\big)\Big).$
Let $\alpha_{i}=\alpha_{{\bm{\beta}}}(X_{i})$. Then the objective can be written as
$\max_{{\bm{\beta}}\in{\mathbb{R}}^{p}}\inf_{{\bm{\alpha}}\in{\mathcal{A}}^{n}}$ $\ \ \ \frac{1}{n}\sum^{n}_{i=1}\left(g(\alpha_{i})+\sum^{p}_{j=1}\left({\bm{\beta}}_{j}\frac{1}{n}\sum^{n}_{i=1}\Big(m(W_{i},(\partial g)\circ\phi_{j})-\alpha_{i}\widetilde{g}\left(X_{i},\phi_{j}(X_{i})\right)\Big)-|{\bm{\beta}}_{j}|\lambda\right)\right).$
From $\partial g\big(\alpha_{{\bm{\beta}}}(X_{i})\big)=\sum^{p}_{j=1}{\bm{\beta}}_{j}\widetilde{g}\left(X_{i},\phi_{j}(X_{i})\right)$, we have
$\max_{{\bm{\beta}}\in{\mathbb{R}}^{p}}\inf_{{\bm{\alpha}}\in{\mathcal{A}}^{n}}$ $\ \ \ \left\{\frac{1}{n}\sum^{n}_{i=1}\Big(g(\alpha_{i})-\alpha_{i}\partial g\big(\alpha_{{\bm{\beta}}}(X_{i})\big)+m\big(W_{i},(\partial g)\circ\alpha_{{\bm{\beta}}}\big)\Big)+\lambda\|{\bm{\beta}}\|_{1}\right\}.$
Consider the problem
$\inf_{{\bm{\alpha}}\in{\mathcal{A}}^{n}}\ \ \ \left\{\frac{1}{n}\sum^{n}_{i=1}\Big(g(\alpha_{i})-\alpha_{i}\partial g\big(\alpha_{{\bm{\beta}}}(X_{i})\big)+m\big(W_{i},(\partial g)\circ\alpha_{{\bm{\beta}}}\big)\Big)+\lambda\|{\bm{\beta}}\|_{1}\right\}.$
Since $g$ is twice differentiable and strictly convex, the infimum is attained when
$\alpha_{i}=\alpha_{{\bm{\beta}}}(X_{i})={\bm{\phi}}(X_{i})^{\top}{\bm{\beta}},\qquad i=1,\dots,n.$
Substituting $\alpha_{i}=\alpha_{{\bm{\beta}}}(X_{i})={\bm{\phi}}(X_{i})^{\top}{\bm{\beta}}$, we obtain
$\max_{{\bm{\beta}}\in{\mathbb{R}}^{p}}$ $\ \ \ \left\{\frac{1}{n}\sum^{n}_{i=1}\Big(g(\alpha_{{\bm{\beta}}}(X_{i}))-\alpha_{i}\partial g\big(\alpha_{{\bm{\beta}}}(X_{i})\big)+m\big(W_{i},(\partial g)\circ\alpha_{{\bm{\beta}}}\big)\Big)-\lambda\|{\bm{\beta}}\|_{1}\right\}.$
This is equivalent to
$\min_{{\bm{\beta}}\in{\mathbb{R}}^{p}}$ $\ \ \ \left\{\frac{1}{n}\sum^{n}_{i=1}\Big(-g(\alpha_{i})+\alpha_{i}\partial g\big(\alpha_{{\bm{\beta}}}(X_{i})\big)-m\big(W_{i},(\partial g)\circ\alpha_{{\bm{\beta}}}\big)\Big)+\lambda\|{\bm{\beta}}\|_{1}\right\}.$
Appendix C Preliminary for the Convergence Rate Analysis
This section introduces notions that are useful for the theoretical analysis.
C.1 Rademacher complexity
Let $\sigma_{1},\dots,\sigma_{n}$ be $n$ independent Rademacher random variables; that is, independent random variables for which $P(\sigma_{i}=1)=P(\sigma_{i}=-1)=1/2$. Let us define
$\mathfrak{R}_{n}f\coloneqq\frac{1}{n}\sum^{n}_{i=1}\sigma_{i}f(W_{i}).$
Additionally, given a class $\mathcal{F}$}, we define
$\mathfrak{R}_{n}{\mathcal{F}}\coloneqq\sup_{f\in{\mathcal{F}}}\mathfrak{R}_{n}f.$
Then, we define the Rademacher average as ${\mathbb{E}}[\mathfrak{R}_{n}{\mathcal{F}}]$ and the empirical Rademacher average as ${\mathbb{E}}_{\sigma}[\mathfrak{R}_{n}{\mathcal{F}}\mid X_{1},\dots,X_{n}]$.
C.2 Local Rademacher complexity bound
Let $\mathcal{F}$} be a class of functions that map $\mathcal{X}$} into $[a,b]$. For $f\in{\mathcal{F}}$, let us define
$Pf\coloneqq{\mathbb{E}}[f(W)],$
$P_{n}f\coloneqq\frac{1}{n}\sum^{n}_{i=1}f(W_{i}).$
We introduce the following result about the Rademacher complexity.
Proposition C.1 (From Theorem 2.1 in Bartlett et al. (2005)).
Let $\mathcal{F}$} be a class of functions that map $\mathcal{X}$} into $[a,b]$. Assume that there is some $r>0$ such that for every $f\in{\mathcal{F}}$, $\text{Var}(f(W))\leq r$. Then, for every $z>0$, with probability at least $1-\exp(-z)$, it holds that
$\sup_{f\in{\mathcal{F}}}\Big(Pf-P_{n}f\Big)\leq\inf_{\alpha>0}\left\{2(1+\alpha){\mathbb{E}}[\mathfrak{R}_{n}f]+\sqrt{\frac{2rx}{n}}+(b-a)\left(\frac{1}{3}+\frac{1}{\alpha}\right)\frac{z}{n}\right\}.$
C.3 Bracketing entropy
We define the bracketing entropy. For a more detailed definition, see Definition 2.2 in van de Geer (2000).
Definition C.1.
Bracketing entropy. Given a class of functions $\mathcal{F}$}, the logarithm of the smallest number of balls in a norm $\|\cdot\|_{2,P}$ of radius $\delta>0$ needed to cover $\mathcal{F}$} is called the $\delta$-entropy with bracketing of $\mathcal{F}$} under the $L_{2}(P)$ metric, denoted by $H_{B}(\delta,{\mathcal{F}},P)$.
C.4 Talagrand’s concentration inequality
We introduce Talagrand’s lemma.
Proposition C.2 (Talagrand’s Lemma).
[…]
$\mathfrak{R}_{n}(\phi\circ{\mathcal{F}})\leq L\mathfrak{R}_{n}({\mathcal{F}}).$
Appendix D Basic inequalities
D.1 Strong convexity
Lemma D.1 ($L_{2}$ distance bound from Lemma 4 in Kato & Teshima (2021)).
If $\inf_{\alpha\in(-\infty),\infty}g^{\prime\prime}(\alpha)>0$, then there exists $\mu>0$ such that for all $\alpha\in{\mathcal{A}}$,
$\|\alpha-\alpha_{0}\|^{2}_{2}\leq\frac{2}{\mu}\Big(\text{BD}_{g}(\alpha)-\text{BD}_{g}(\alpha_{0})\Big)$
holds.
From the strong convexity and Lemma D.1, we have
$\frac{\mu}{2}\|\widehat{\alpha}-\alpha_{0}\|^{2}_{2}\leq\text{BD}_{g}(\widehat{\alpha})-\text{BD}_{g}\big(\alpha_{0}\big).$
Recall that we have defined an estimator $\widehat{r}$ as follows:
$\widehat{\alpha}\coloneqq\operatorname*{arg\,min}_{\alpha\in{\mathcal{A}}}\widehat{\text{BD}}_{g}(\alpha)+\lambda J(\alpha),$
where $J\left(h\right)$ is some regularization term.
D.2 Preliminary
Proposition D.2.
The estimator $\widehat{r}$ satisfies the following inequality:
$\widehat{\text{BD}}_{g}(\widehat{\alpha})+\lambda J(\widehat{\alpha})\leq\widehat{\text{BD}}_{g}(\alpha^{*})+\lambda J(\alpha^{*}),$
where recall that
$\widehat{\text{BD}}_{g}(\alpha)\coloneqq\frac{1}{n}\sum^{n}_{i=1}\Big(-g(\alpha(X_{i}))+\partial g(\alpha(X_{i}))\alpha(X_{i})-\partial g(\alpha(1,X_{i}))-\partial g(\alpha(0,X_{i}))\Big).$
Let $Z\in{\mathcal{Z}}$ be a random variable with a space $\mathcal{Z}$}, and $\{Z_{i}\}^{n}_{i=1}$ be its realizations. For a function $f\colon{\mathcal{Z}}\to{\mathbb{R}}$ and $X$ following $P$, let us denote the sample mean as
$\widehat{{\mathbb{E}}}[f(Z)]\coloneqq\frac{1}{n}\sum^{n}_{i=1}f(Z_{i}).$
We also denote $\widehat{{\mathbb{E}}}[f(Z)]-{\mathbb{E}}[f(Z)]=(\widehat{{\mathbb{E}}}-{\mathbb{E}})f(Z)$
D.3 Risk bound
Recall that
$\widehat{\text{BD}}_{g}(\alpha)=\frac{1}{n}\sum^{n}_{i=1}\Big(-g(\alpha(X_{i}))+\partial g(\alpha(X_{i}))\alpha(X_{i})-\partial g(\alpha(1,X_{i}))-\partial g(\alpha(0,X_{i}))\Big).$
Let us define
$L(h,D,X)\coloneqq-g(\alpha(X))+\partial g(\alpha(X))\alpha(X)-\partial g(\alpha(1,X))-\partial g(\alpha(0,X)),$
and we can write
$\widehat{\text{BD}}_{g}(\alpha)=\widehat{{\mathbb{E}}}\big[L(h,D,X)\big]$
Then, from Proposition D.2, we have
$\widehat{{\mathbb{E}}}\big[L(\alpha^{*},D,X)\big]-\widehat{{\mathbb{E}}}\big[L(\widehat{\alpha},D,X)\big]+\lambda J(\widehat{\alpha})-\lambda J(\alpha^{*})\geq 0.$
Throughout the proof, we use the following basic inequalities that hold for $\widehat{\alpha}$.
Proposition D.3.
The estimator $\widehat{r}$ satisfies the following inequality:
$\frac{\mu}{2}\left\|\widehat{\alpha}(X)-\alpha_{0}(X)\right\|^{2}_{L_{2}(P_{0})}$
[…]
Proof of Proposition D.2 is trivial. We prove Proposition D.3 below.
Proof.
From the strong convexity and Lemma D.1, we have
$\frac{\mu}{2}\|\widehat{\alpha}-\alpha_{0}\|^{2}_{2}\leq\text{BD}_{g}(\widehat{\alpha})-\text{BD}_{g}\big(\alpha_{0}\big)={\mathbb{E}}\left[L(\widehat{\alpha},D,X)-L(\alpha_{0},D,X)\right].$
From Proposition D.2, we have
$\frac{\mu}{2}\left\|\widehat{\alpha}(X)-\alpha_{0}(X)\right\|^{2}_{L_{2}(P_{0})}$
$\leq{\mathbb{E}}\left[L(\widehat{\alpha},D,X)-L(\alpha_{0},D,X)\right]$
$={\mathbb{E}}\left[L(\widehat{\alpha},D,X)-L(\alpha_{0},D,X)\right]$
$\ \ \ \ \ \ -\widehat{{\mathbb{E}}}\left[L(\widehat{\alpha},D,X)-L(\alpha_{0},D,X)\right]$
$\ \ \ \ \ \ +\widehat{{\mathbb{E}}}\left[L(\widehat{\alpha},D,X)-L(\alpha_{0},D,X)\right]$
$\leq{\mathbb{E}}\left[L(\widehat{\alpha},D,X)-L(\alpha_{0},D,X)\right]$
$\ \ \ \ \ \ -\widehat{{\mathbb{E}}}\left[L(\widehat{\alpha},D,X)-L(\alpha_{0},D,X)\right]$
$\ \ \ \ \ \ +\widehat{{\mathbb{E}}}\left[L(\widehat{\alpha},D,X)-L(\alpha_{0},D,X)\right]$
$\ \ \ \ \ \ -\widehat{{\mathbb{E}}}\left[L(\widehat{\alpha},D,X)-L(\alpha^{*},D,X)\right]+\lambda J(\widehat{\alpha})-\lambda J(\alpha_{0}).$
Appendix E Proof of Theorem 6.1
We show Theorem 6.1 by bounding
$\left({\mathbb{E}}-\widehat{{\mathbb{E}}}\right)\left[L(\widehat{\alpha},D,X)-L(\alpha_{0},D,X)\right],$ (8)
in Proposition D.3. We can bound this term by using the empirical-process arguments.
Note that since $\alpha_{0}\in{\mathcal{A}}$, it holds that $\alpha^{*}=\alpha_{0}$, which implies that
E.1 Preliminary
We introduce the following propositions from van de Geer (2000), Kanamori et al. (2012) and Kato & Teshima (2021).
Definition E.1 (Derived function class and bracketing entropy (from Definition 4 in Kato & Teshima (2021))).
Given a real-valued function class $\mathcal{F}$}, define $\ell\circ{\mathcal{F}}:=\{\ell\circ f\colon f\in{\mathcal{F}}\}$. By extension, we define $I:\ell\circ{\mathcal{A}}\to[1,\infty)$ by $I(\ell\circ h)=I(\alpha)$ and $\ell\circ{\mathcal{A}}_{M}:=\{\ell\circ\alpha\colon\alpha\in{\mathcal{A}}_{M}\}$. Note that, as a result, $\ell\circ{\mathcal{A}}_{M}$ coincides with $\{\ell\circ\alpha\in\ell\circ{\mathcal{A}}:I(\ell\circ h)\leq M\}$.
Proposition E.1.
Let $\ell:{\mathbb{R}}\to{\mathbb{R}}$ be a $v$-Lipschitz continuous function. Let $H_{B}\big(\delta,{\mathcal{F}},\|\cdot\|_{L_{2}(P_{0})}\big)$ denote the bracketing entropy of $\mathcal{F}$} with respect to a distribution $P$. Then, for any distribution $P$, any $\gamma>0$, any $M\geq 1$, and any $\delta>0$, we have
$H_{B}\big(\delta,\ell\circ{\mathcal{A}},\|\cdot\|_{L_{2}(P_{0})}\big)$ $\leq\frac{(s+1)(2v)^{\gamma}}{\gamma}\left(\frac{M}{\delta}\right)^{\gamma}.$
Moreover, there exists $M>0$ such that for any $M\geq 1$ and any distribution $P$,
$\sup_{\ell\circ\alpha\in\ell\circ{\mathcal{A}}_{M}}\|\ell\circ\alpha-\ell\circ\alpha^{*}\|_{L_{2}(P_{0})}$ $\leq c_{0}vM,$
$\sup_{\stackrel{{\scriptstyle\ell\circ\alpha\in\ell\circ{\mathcal{A}}_{M}}}{{\|\ell\circ\alpha-\ell\circ\alpha^{*}\|_{L_{2}(P_{0})}\leq\delta}}}\|\ell\circ\alpha-\ell\circ\alpha^{*}\|_{\infty}$ $\leq c_{0}vM,\quad\text{for all }\delta>0.$
Proposition E.2 (Lemma 5.13 in van de Geer (2000), Proposition 1 in Kanamori et al. (2012)).
Let ${\mathcal{F}}\subset L^{2}(P)$ be a function class and the map $I(f)$ be a complexity measure of $f\in{\mathcal{F}}$, where $I$ is a non-negative function on $\mathcal{F}$} and $I(f_{0})<\infty$ for a fixed $f_{0}\in{\mathcal{F}}$. We now define ${\mathcal{F}}_{M}=\{f\in{\mathcal{F}}:I(f)\leq M\}$ satisfying ${\mathcal{F}}=\bigcup_{M\geq 1}{\mathcal{F}}_{M}$. Suppose that there exist $c_{0}>0$ and $0<\gamma<2$ such that
$\sup_{f\in{\mathcal{F}}_{M}}\|f-f_{0}\|\leq c_{0}M,\ \sup_{\stackrel{{\scriptstyle f\in{\mathcal{F}}_{M}}}{{\|f-f_{0}\|_{L^{2}(P)}\leq\delta}}}\|f-f_{0}\|_{\infty}\leq c_{0}M,\quad\text{for all }\delta>0,$
and that $H_{B}(\delta,{\mathcal{F}}_{M},P)=O\left((M/\delta)^{\gamma}\right)$. Then, we have
$\sup_{f\in{\mathcal{F}}}\frac{\left|\int(f-f_{0})d(P-P_{n})\right|}{D(f)}=O_{p}(1),\ (n\to\infty),$
where $D(f)$ is defined by
$D(f)=\max{\frac{\|f-f_{0}\|_{L^{2}(P)}^{1-\gamma/2}I(f)^{\gamma/2}}{\sqrt{n}}}{\frac{I(f)}{n^{2/(2+\gamma)}}}.$
Proposition E.3.
Let $g\colon{\mathcal{K}}\to{\mathbb{R}}$ be twice continuously differentiable and strictly convex for the space $\mathcal{K}$} of $\alpha_{0}$, and suppose that there exists $M>0$ such that
$|g^{\prime\prime}(t)|\leq M\quad\text{for all }t\in{\mathbb{R}}.$
Let $\zeta^{-1}\colon{\mathbb{R}}\to{\mathbb{R}}$ be continuously differentiable and globally Lipschitz, that is, there exists $L_{\zeta}>0$ such that
$|\zeta^{-1}(s)-\zeta^{-1}(t)|\leq L_{\zeta}|s-t|\quad\text{for all }s,t\in{\mathbb{R}}.$
Assume also that $\zeta^{-1}(0)$ is finite, and define
$a_{0}:=|\zeta^{-1}(0)|,\qquad a_{1}:=L_{\zeta},$
so that
$|\zeta^{-1}(u)|\leq a_{0}+a_{1}|u|\quad\text{for all }u\in{\mathbb{R}}.$
Let $h$ be a bounded real-valued function on the domain of $(D,X)$, and write
$\|h\|_{\infty}:=\sup_{d,x}|\alpha(d,x)|.$
Let $L$ be a linear functional acting on bounded functions, such that for some constant $C_{L}>0$,
$|L(f)|\leq C_{L}\bigl(1+\|f\|_{\infty}\bigr)\quad\text{for all bounded }f.$
Define
$L(\zeta^{-1}\circ f)$ $=g\bigl(\zeta^{-1}\circ f(D,X)\bigr)+\partial g\bigl(\zeta^{-1}\circ f(D,X)\bigr)\zeta^{-1}\circ\alpha(D,X)$
$\quad-\partial g\bigl(\zeta^{-1}\circ f(1,X)\bigr)-\partial g\bigl(\zeta^{-1}\circ f(0,X)\bigr).$
Then there exists a constant $C>0$ (depending only on $g$, $\zeta^{-1}$ and $C_{L}$) such that
$|L(\zeta^{-1}\circ f)|\leq C\bigl(1+\|f\|_{\infty}^{2}\bigr).$
E.2 Upper bound using the empirical-process arguments
From Propositions E.1E.3, we obtain the following result.
Proposition E.4.
Under the conditions of Theorem 6.1, for any $0<\gamma<2$, we have
$d\left({\mathbb{E}}-\widehat{{\mathbb{E}}}\right)\left[L(\widehat{\alpha},D,X)-L(\alpha_{0},D,X)\right]$
$=O_{p}\left(\max\left\{\frac{\|\widehat{\alpha}-\alpha^{*}\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{\alpha}\right\|_{{\mathcal{A}}}\right)^{1+\gamma/2}}{\sqrt{n}},\frac{\left(1+\left\|\widehat{\alpha}\right\|_{{\mathcal{A}}}\right)^{2}}{n^{2/(2+\gamma)}}\right\}\right),$
as $n\to\infty$.
E.3 Proof of Theorem 6.1
We prove Theorem 6.1 following the arguments in Kanamori et al. (2012).
Proof.
From Proposition D.3 and $\alpha_{0}\in\cal\alpha^{\text{RKHS}}$, we have
$\left\|\widehat{\alpha}(X)-\alpha_{0}(X)\right\|^{2}_{L_{2}(P_{0})}+\lambda\|\widehat{\alpha}\|^{2}_{{\mathcal{A}}}$
$\leq\left({\mathbb{E}}-\widehat{{\mathbb{E}}}\right)\left[L(\widehat{\alpha},D,X)-L(\alpha_{0},D,X)\right]+\lambda\|f_{0}\|^{2}_{{\mathcal{A}}}.$
From Proposition E.4, we have
$\left\|\widehat{\alpha}(X)-\alpha_{0}(X)\right\|^{2}_{L_{2}(P_{0})}+\lambda\|\widehat{f}\|^{2}_{{\mathcal{A}}}$
$=O_{p}\left(\max\left\{\frac{\|\widehat{\alpha}-\alpha_{0}\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{f}\right\|_{{\mathcal{A}}}\right)^{1+\gamma/2}}{\sqrt{n}},\frac{\left(1+\left\|\widehat{\alpha}\right\|_{{\mathcal{A}}}\right)^{2}}{n^{2/(2+\gamma)}}\right\}\right)+\lambda\|r_{0}\|^{2}_{{\mathcal{A}}}.$
We consider the following three possibilities:
$\left\|\widehat{\alpha}(X)-\alpha_{0}(X)\right\|^{2}_{L_{2}(P_{0})}+\lambda\|\widehat{f}\|^{2}_{{\mathcal{A}}}=O_{p}(\lambda),$ (9)
$\left\|\widehat{\alpha}(X)-\alpha_{0}(X)\right\|^{2}_{L_{2}(P_{0})}+\lambda\|\widehat{f}\|^{2}_{{\mathcal{A}}}=O_{p}\left(\frac{\|\widehat{f}-f_{0}|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{f}\right\|_{{\mathcal{A}}}\right)^{1+\gamma/2}}{\sqrt{n}}\right),$ (10)
$\left\|\widehat{\alpha}(X)-\alpha_{0}(X)\right\|^{2}_{L_{2}(P_{0})}+\lambda\|\widehat{f}\|^{2}_{{\mathcal{A}}}=O_{p}\left(\frac{\left(1+\left\|\widehat{f}\right\|_{{\mathcal{A}}}\right)^{2}}{n^{2/(2+\gamma)}}\right).$ (11)
The above inequalities are analyzed as follows:
Case (9).
We have
$\left\|\widehat{\alpha}(X)-\alpha_{0}(X)\right\|^{2}_{L_{2}(P_{0})}=O_{p}(\lambda),$
$\lambda\|\widehat{f}\|^{2}_{{\mathcal{A}}}=O_{p}(\lambda).$
Therefore, we have $\left\|\widehat{\alpha}(X)-\alpha_{0}(X)\right\|_{P_{0}}=O_{p}(\lambda^{1/2})$ and $\|\widehat{r}\|_{{\mathcal{A}}}=O_{p}(1)$.
Case (10).
We have
$\left\|\widehat{\alpha}(X)-\alpha_{0}(X)\right\|^{2}_{L_{2}(P_{0})}=O_{p}\left(\frac{\|\widehat{f}-f_{0})\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{f}\right\|_{{\mathcal{F}}}\right)^{1+\gamma/2}}{\sqrt{n}}\right),$
$\lambda\|\widehat{f}\|^{2}_{{\mathcal{A}}}=O_{p}\left(\frac{\|\widehat{f}-f_{0})\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{f}\right\|_{{\mathcal{F}}}\right)^{1+\gamma/2}}{\sqrt{n}}\right).$
From the first inequality, we have
$\left\|\widehat{\alpha}(X)-\alpha_{0}(X)\right\|_{P_{0}}=\sum_{d\in\{1,0\}}O_{p}\left(\frac{\left(1+\left\|\widehat{f}\right\|_{{\mathcal{F}}}\right)^{1+\gamma/2}}{n^{1/(2+\gamma)}}\right).$
By using this result, from the second inequality, we have
$\lambda\|\widehat{f}\|^{2}_{{\mathcal{A}}}$ $=O_{p}\left(\frac{\|\widehat{f}-f_{0})\|_{L^{2}(P_{0})}^{1-\gamma/2}\left(1+\left\|\widehat{f}\right\|_{{\mathcal{F}}}\right)^{1+\gamma/2}}{\sqrt{n}}\right)$
$=O_{p}\left(\left(\frac{1+\left\|\widehat{f}\right\|_{{\mathcal{F}}}}{n^{1/(2+\gamma)}}\right)^{1-\gamma/2}\frac{\left(1+\left\|\widehat{f}\right\|_{{\mathcal{F}}}\right)^{1+\gamma/2}}{\sqrt{n}}\right)$
$=O_{p}\left(\frac{\left(1+\left\|\widehat{f}\right\|_{{\mathcal{F}}}\right)^{2}}{n^{2/(2+\gamma)}}\right).$
This implies that
$\|\widehat{f}\|_{{\mathcal{A}}}$ $=O_{p}\left(\frac{\left(1+\left\|\widehat{f}\right\|_{{\mathcal{F}}}\right)^{2}}{\lambda^{1/2}n^{2/(2+\gamma)}}\right)=o_{p}(1).$
Therefore, the following inequity is obtained.
$\left\|\widehat{\alpha}(X)-\alpha_{0}(X)\right\|_{P_{0}}=O_{p}\left(\frac{1}{n^{1/(2+\gamma)}}\right)=O_{p}(\lambda^{1/2}).$
Case 11.
We have
$\left\|\widehat{\alpha}(X)-\alpha_{0}(X)\right\|^{2}_{L_{2}(P_{0})}=O_{p}\left(\frac{\left(1+\left\|\widehat{f}\right\|_{{\mathcal{F}}}\right)^{2}}{n^{2/(2+\gamma)}}\right),$
$\lambda\|\widehat{f}\|^{2}_{{\mathcal{A}}}=O_{p}\left(\frac{\left(1+\left\|\widehat{f}\right\|_{{\mathcal{F}}}\right)^{2}}{n^{2/(2+\gamma)}}\right).$
As well as the argument in (10), we have $\|\widehat{r}\|_{{\mathcal{A}}}=o_{p}(1)$. Therefore, we have
$\left\|\widehat{\alpha}(X)-\alpha_{0}(X)\right\|_{P_{0}}=O_{p}\left(\frac{1}{n^{1/(2+\gamma)}}\right)=O_{p}(\lambda^{1/2}).$
Appendix F Proof of Theorem 6.2
Our proof procedure mainly follows those in Kato & Teshima (2021) and Zheng et al. (2022). In particular, we are inspired by the proof in Zheng et al. (2022).
We prove Theorem 6.2 by proving the following lemma:
Lemma F.1.
Suppose that Assumption 6.3 holds. For any $n\geq\text{Pdim}({\mathcal{F}}^{\text{FNN}})$, there exists a constant $C>0$ depending on $(\mu,\sigma,M)$ such that for any $\gamma>0$, with probability at least $1-\exp(-\gamma)$, it holds that
$\Big\|\widehat{f}-f_{0}\Big\|_{2}\leq C\left(\sqrt{\frac{\text{Pdim}({\mathcal{F}}^{\text{FNN}})\log(n)}{n}}+\big\|f^{*}-f_{0}\big\|_{2}+\sqrt{\frac{\gamma}{n}}\right).$
As shown in Zheng et al. (2022), we can bound $\text{Pdim}({\mathcal{F}}^{\text{FNN}})\log(n)$ by specifying neural networks and obtain Theorem 6.2.
F.1 Proof of Lemma F.1
We prove Lemma F.1 by bounding (8) in Proposition D.3.
To bound (8), we show several auxiliary results. Define
$\widehat{{\mathcal{F}}}^{f^{*},u}$ $\coloneqq\{f\in{\mathcal{F}}^{\text{FNN}}\colon\frac{1}{n}\sum^{n}_{i=1}(f(X_{i})-f^{*}(X_{i}))^{2}\leq u\},$
$\overline{{\mathcal{G}}}^{f^{*},u}$ $\coloneqq\Big\{(f-f^{*})\colon f\in\widehat{{\mathcal{F}}}^{f^{*},u}\Big\},$
$\kappa^{u}_{n}(u)$ $\coloneqq{\mathbb{E}}_{\sigma}\left[\mathfrak{R}_{n}\overline{{\mathcal{G}}}^{f^{*},u}\right],$
$u^{\dagger}$ $\coloneqq\inf\big\{u\geq 0\colon\kappa^{u}_{n}(s)\leq s^{2}\ \ \ \forall s\geq u\big\}.$
Here, we show the following two lemmas:
Lemma F.2 (Corresponding to (26) in Zheng et al. (2022)).
Suppose that the conditions in Lemma F.1 hold. Then, for any $z>0$, with probability $1-\exp(-z)$ it holds that
$\widehat{{\mathbb{E}}}\left[L(\widehat{\alpha},D,X)-L(\alpha_{0},D,X)\right]$
$\leq C\left(\|f^{*}(X)-f_{0}(X)\|^{2}_{2}+\|f^{*}(X)-f_{0}(X)\|_{2}\sqrt{\frac{z}{n}}+\frac{16Mz}{3n}\right).$
Lemma F.3 (Corresponding to (29) in Zheng et al. (2022)).
Suppose that the conditions in Lemma F.1 hold. If there exists $u_{0}>0$ such that
$\|\widehat{f}(X)-f^{*}(X)\|_{2}\leq u_{0},$
then it holds that
$\left({\mathbb{E}}-\widehat{{\mathbb{E}}}\right)\left[L(\widehat{\alpha},D,X)-L(\alpha_{0},D,X)\right]$
$\leq C\left({\mathbb{E}}_{\sigma}\left[\mathfrak{R}_{n}\overline{{\mathcal{G}}}^{f^{*},u_{0}}\right]+u_{0}\sqrt{\frac{z}{n}}+\frac{Mz}{n}\right).$
Additionally, we use the following three propositions directly from Zheng et al. (2022).
Proposition F.4 (From (32) in Zheng et al. (2022)).
Let $u>0$ be a positive value such that
$\|f-f_{0}\|_{2}\leq u$
for all $f\in{\mathcal{F}}$. Then, for every $z>0$, with probability at least $1-2\exp(-z)$, it holds that
$\sqrt{\frac{1}{n}\sum^{n}_{i=1}\big(f(X_{i})-f_{0}(X_{i})\big)^{2}}\leq 2u.$
Proposition F.5 (Corresponding to (36) in Step 3 of Zheng et al. (2022)).
Suppose that the conditions in Lemma F.1 hold. Then, there exists a universal constant $C>0$ such that
$u^{\dagger}\leq CM\sqrt{\frac{\text{Pdim}({\mathcal{F}}^{\text{FNN}})\log(n)}{n}}.$
Proposition F.6 (Upper bound of the Rademacher complexity).
Suppose that the conditions in Lemma F.1 hold. If $n\geq\text{Pdim}({\mathcal{F}}^{\text{FNN}})$, $u_{0}\geq 1/n$, and $n\geq(2eM)^{2}$, we have
${\mathbb{E}}_{\sigma}\left[\mathfrak{R}_{n}\overline{{\mathcal{G}}}^{f^{*},u_{0}}\right]\leq Cr_{0}\sqrt{\frac{\text{Pdim}({\mathcal{F}}^{\text{FNN}})\log n}{n}}.$
Then, we prove Lemma F.1 as follows:
Proof of Lemma F.1.
If there exists $u_{0}>0$ such that
$\|\widehat{f}(X)-f^{*}(X)\|_{2}\leq u_{0},$
then from (8) and Lemmas F.2 and F.3, for every $z>0$, there exists a constant $C>0$ independent $n$ such that
$\left\|\widehat{\alpha}(X)-\alpha_{0}(X)\right\|^{2}_{L_{2}(P_{0})}$
$\leq C\left(\|f^{*}-f_{0}\|_{2}\sqrt{\frac{z}{n}}+\frac{16Mz}{3n}+u_{0}\sqrt{\frac{\text{Pdim}({\mathcal{F}}^{\text{FNN}})\log n}{n}}+u_{0}\sqrt{\frac{z}{n}}+\frac{Mz}{n}\right).$ (12)
This result implies that if $\sqrt{\text{Pdim}({\mathcal{F}}^{\text{FNN}})}$, then there exists $n_{0}$ such that for all $n>n_{0}$, there exists $u_{1}<u_{0}$ such that
$\left\|\widehat{\alpha}(X)-\alpha_{0}(X)\right\|^{2}_{L_{2}(P_{0})}\leq u_{1}.$
For any $z>0$, define $\overline{u}$ as
$\overline{u}_{z}\geq\max\left\{\sqrt{\log(n)/n},4\sqrt{3}M\sqrt{z/n},u^{\dagger}\right\}.$
Define a subspace of ${\mathcal{F}}^{\text{FNN}}$ as
${\mathcal{S}}^{\text{FNN}}(f_{0},\overline{u}_{z}\coloneqq\big\{f\in{\mathcal{F}}^{\text{FNN}}\colon\|f-f_{0}\|\leq\overline{u}_{z}\big\}.$
Define
$\ell\coloneqq\lfloor\log_{2}(2M/\sqrt{\log(n)/n})\rfloor.$
Using the definition of subspaces, we divide ${\mathcal{F}}^{\text{FNN}}$ into the following $\ell+1$ subspaces:
$\overline{{\mathcal{S}}}^{\text{FNN}}_{0}\coloneqq$ ${\mathcal{S}}^{\text{FNN}}(f_{0},\overline{u}),$
$\overline{{\mathcal{S}}}^{\text{FNN}}_{1}\coloneqq$ ${\mathcal{S}}^{\text{FNN}}(f_{0},\overline{u})\backslash{\mathcal{S}}^{\text{FNN}}(f_{0},\overline{u}),$
$\quad\vdots$
$\overline{{\mathcal{S}}}^{\text{FNN}}_{\ell}\coloneqq$ ${\mathcal{S}}^{\text{FNN}}(f_{0},2^{\ell}\overline{u})\backslash{\mathcal{S}}^{\text{FNN}}(f_{0},2^{\ell-1}\overline{u}).$
Since $\overline{u}_{z}>u^{\dagger}$, from the definition of $u^{\dagger}$, we have
$\overline{u}^{2}_{z}\leq\kappa^{u}_{n}(\overline{u}).$
If there exists $j\leq\ell$ such that $\widehat{f}\in\overline{{\mathcal{S}}}^{\text{FNN}}_{j}$, then from (F.1), for every $z>0$, with probability at least $1-8\exp(-z)$, there exists a constant $C>0$ independent of $n$ such that
$\left\|\widehat{\alpha}(X)-\alpha_{0}(X)\right\|^{2}_{2}$
$\leq C\left(2^{\ell-1}\overline{u}\left(\sqrt{\frac{\text{Pdim}({\mathcal{F}}^{\text{FNN}})\log(n)}{n}}+\sqrt{\frac{z}{n}}\right)+\|f^{*}-f_{0}\|^{2}_{2}+\|f^{*}-f_{0}\|_{2}\sqrt{\frac{z}{n}}+\frac{Mz}{n}\right).$ (13)
Additionally, if
$C\left(\sqrt{\frac{\text{Pdim}({\mathcal{F}}^{\text{FNN}})\log(n)}{n}}+\sqrt{\frac{z}{n}}\right)\leq\frac{1}{8}2^{j}\overline{u},$ (14)
$C\left(\|f^{*}-f_{0}\|^{2}_{2}+\|f^{*}-f_{0}\|_{2}\sqrt{\frac{z}{n}}+\frac{Mz}{n}\right)\leq\frac{1}{8}2^{2j}\overline{u}^{2}$ (15)
hold, then
$\left\|\widehat{\alpha}(X)-\alpha_{0}(X)\right\|_{2}\leq 2^{j-1}\overline{u}.$ (16)
Here, to obtain (16), we used $\overline{u}\geq\max\left\{\sqrt{\log(n)/n},4\sqrt{3}M\sqrt{z/n},u^{\dagger}\right\}$, (F.1), (14), and (15).
From Proposition F.5, it holds that
$u^{\dagger}\leq CM\sqrt{\frac{\text{Pdim}({\mathcal{F}}^{\text{FNN}})\log(n)}{n}}.$
Therefore, we can choose $\overline{u}$ as
$\overline{u}\coloneqq C\left(\sqrt{\frac{\text{Pdim}({\mathcal{F}}^{\text{FNN}})\log(n)}{n}}+\sqrt{\log(n)/n}+4\sqrt{3}M\sqrt{z/n}\right),$
where $C>0$ is a constant independent of $n$. ∎
F.2 Proof of Lemma F.2
From Proposition C.1, we have
$\widehat{{\mathbb{E}}}\left[L(\widehat{\alpha},D,X)-L(\alpha_{0},D,X)\right]$
$\leq{\mathbb{E}}\left[L(\widehat{\alpha},D,X)-L(\alpha_{0},D,X)\right]+\sqrt{2}C\|f^{*}(X)-f_{0}(X)\|\sqrt{\frac{z}{n}}+\frac{16C_{1}Mz}{3n}.$
This is a direct consequence of Proposition C.1. Note that $\alpha^{*}$ and $\alpha_{0}$ are fixed, and it is enough to apply the standard law of large numbers; that is, we do not have to consider the uniform law of large numbers. However, we can still apply Proposition C.1, which is a general than the standard law of large numbers, with ignoring the Rademacher complexity part.
We have
$\widehat{{\mathbb{E}}}\left[L(\widehat{\alpha},D,X)-L(\alpha_{0},D,X)\right]$
$\leq{\mathbb{E}}\left[L(\widehat{\alpha},D,X)-L(\alpha_{0},D,X)\right]$
$\ \ \ \ \ +\sqrt{2}C_{1}\|f^{*}-f_{0}\|\sqrt{\frac{z}{n}}+\frac{16C_{2}Mz}{3n}+\sqrt{2}C_{2}\|f^{*}-f_{0}\|\sqrt{\frac{z}{n}}+\frac{16C_{2}Mz}{3n}$
$\leq C\left(\|f^{*}-f_{0}\|^{2}_{2}+\|f^{*}-f_{0}\|\sqrt{\frac{z}{n}}+\frac{16CMz}{3n}\right).$
F.3 Proof of Lemma F.3
Let $g\coloneqq(f-f^{*})^{2}$. From the definition of FNNs, we have
$g\leq 4M^{2}$
Additionally, we assumed that $\|\widehat{f}-f^{*}\|_{2}\leq u_{0}$ holds. Then, it holds that $\text{Var}_{P_{0}}(g)\leq 4M^{2}u^{2}_{0}$.
Here, we note that the followings hold for all $f$ ($r$):
$L(\alpha)-L(\alpha^{*})\leq C\Big|f(X)-f^{*}(X)\Big|,$
where $C>0$ is some constant
Then, from Proposition C.1, for every $z>0$, with probability at least $1-\exp(-z)$, it holds that
$\left({\mathbb{E}}-\widehat{{\mathbb{E}}}\right)\left[L(\widehat{\alpha},D,X)-L(\alpha_{0},D,X)\right]$
$\leq C\left({\mathbb{E}}_{\sigma}\left[\mathfrak{R}_{n}\overline{{\mathcal{G}}}^{f^{*},u_{0}}\right]+r_{0}\sqrt{\frac{z}{n}}+\frac{Mz}{n}\right).$
Appendix G Riesz Regression and Density Ratios
As explained in the main text and in Kato (2025d), the Riesz representer is closely connected to density ratio estimation. In particular, for ATE, the Riesz representer can be expressed in terms of two density ratios relative to the marginal covariate distribution, which leads to a decomposition of the squared loss objective into two LSIF problems.
Riesz representer and density ratio.
Let $p_{Z}$ denote the marginal density of $Z$ and $p_{Z\mid D=d}$ the conditional density of $Z$ given $D=d$. Let $\kappa_{d}\coloneqq P_{0}(D=d)$. By Bayes’ rule,
$p_{Z\mid D=d}(z)=\frac{p_{Z}(z)P_{0}(D=d\mid Z=z)}{P_{0}(D=d)}=\frac{p_{Z}(z)e_{0}(z)^{d}(1-e_{0}(z))^{1-d}}{\kappa_{d}},$
where $e_{0}(z)=P_{0}(D=1\mid Z=z)$.
Define the density ratios with respect to the marginal distribution of $Z$ by
$r_{1}(z)\coloneqq\frac{p_{Z}(z)}{p_{Z\mid D=1}(z)},\qquad r_{0}(z)\coloneqq\frac{p_{Z}(z)}{p_{Z\mid D=0}(z)}.$
From the expression above,
$r_{1}(z)=\frac{\kappa_{1}}{e_{0}(z)},\qquad r_{0}(z)=\frac{\kappa_{0}}{1-e_{0}(z)}.$
Therefore, the ATE Riesz representer can be written as
$\alpha^{\text{ATE}}_{0}(D,Z)=\frac{\mathbbm{1}[D=1]}{e_{0}(Z)}-\frac{\mathbbm{1}[D=0]}{1-e_{0}(Z)}=\mathbbm{1}[D=1]\frac{r_{1}(Z)}{\kappa_{1}}-\mathbbm{1}[D=0]\frac{r_{0}(Z)}{\kappa_{0}}.$
Equivalently, estimating $\alpha^{\text{ATE}}_{0}$ reduces to estimating the pair $(r_{1},r_{0})$, which compare the marginal covariate distribution to the treated and control covariate distributions.
Squared loss objective and decomposition into two LSIF problems.
We next connect this representation to LSIF, a density ratio estimation method proposed in Kanamori et al. (2009). Let $g^{\text{SQ}}(u)\coloneqq(u-1)^{2}$ be the squared loss. The corresponding population squared loss Bregman objective can be written as
$\text{BD}_{g^{\text{SQ}}}(\alpha)={\mathbb{E}}\Big[-2\big(\alpha(1,Z)-\alpha(0,Z)\big)+\alpha(D,Z)^{2}\Big],$
where $\alpha(d,Z)$ denotes the value of the representer evaluated at treatment status $d$ and covariates $Z$. Under the parameterization
$\alpha(D,Z)=\mathbbm{1}[D=1]\frac{r_{1}(Z)}{\kappa_{1}}-\mathbbm{1}[D=0]\frac{r_{0}(Z)}{\kappa_{0}},$
we have $\alpha(1,Z)=r_{1}(Z)/\kappa_{1}$ and $\alpha(0,Z)=-r_{0}(Z)/\kappa_{0}$, hence $\alpha(1,Z)-\alpha(0,Z)=r_{1}(Z)/\kappa_{1}+r_{0}(Z)/\kappa_{0}$. Substituting this into $\text{BD}_{g^{\text{SQ}}}(\alpha)$ and using the law of total expectation yields
$\text{BD}_{g^{\text{SQ}}}(\alpha)$ $=-2{\mathbb{E}}\left[\frac{r_{1}(Z)}{\kappa_{1}}+\frac{r_{0}(Z)}{\kappa_{0}}\right]+{\mathbb{E}}\left[\alpha(D,Z)^{2}\right].$ (17)
Moreover,
${\mathbb{E}}\big[\alpha(D,Z)^{2}\big]=\kappa_{1}{\mathbb{E}}\left[\left(\frac{r_{1}(Z)}{\kappa_{1}}\right)^{2}\Bigm|D=1\right]+\kappa_{0}{\mathbb{E}}\left[\left(\frac{r_{0}(Z)}{\kappa_{0}}\right)^{2}\Bigm|D=0\right].$
Rewriting (17) in terms of expectations with respect to $p_{Z}$ and $p_{Z\mid D=d}$ and dropping constants gives
$\text{BD}_{g^{\text{SQ}}}(\alpha)$ $\coloneqq-2{\mathbb{E}}_{Z}\left[r_{1}(Z)\right]+{\mathbb{E}}_{Z\mid D=1}\left[r_{1}(Z)^{2}\right]-2{\mathbb{E}}_{Z}\left[r_{0}(Z)\right]+{\mathbb{E}}_{Z\mid D=0}\left[r_{0}(Z)^{2}\right].$ (18)
Minimizing this objective is exactly LSIF, and in our setting it coincides with SQ-Riesz regression for ATE estimation.
Furthermore, if $r_{1}(\cdot)$ and $r_{0}(\cdot)$ are treated as independent functions, minimizing $\text{BD}_{g^{\text{SQ}}}(\alpha)$ over $(r_{1},r_{0})$ separates into two independent LSIF type problems
$r_{1}^{*}$ $=\operatorname*{arg\,min}_{r_{1}}\left\{-2{\mathbb{E}}_{Z}[r_{1}(Z)]+{\mathbb{E}}_{Z\mid D=1}[r_{1}(Z)^{2}]\right\},$
$r_{0}^{*}$ $=\operatorname*{arg\,min}_{r_{0}}\left\{-2{\mathbb{E}}_{Z}[r_{0}(Z)]+{\mathbb{E}}_{Z\mid D=0}[r_{0}(Z)^{2}]\right\},$
where ${\mathbb{E}}_{Z}$ and ${\mathbb{E}}_{Z\mid D=d}$ denote expectations under $P_{0}(Z)$ and $P_{0}(Z\mid D=d)$. At the sample level, with ${\mathcal{G}}_{1}$ and ${\mathcal{G}}_{0}$ defined as in the Introduction, the empirical LSIF objectives are
$\widehat{R}_{1}(r_{1})$ $\coloneqq-\frac{2}{n}\sum_{i=1}^{n}r_{1}(Z_{i})+\frac{1}{|{\mathcal{G}}_{1}|}\sum_{i\in{\mathcal{G}}_{1}}r_{1}(Z_{i})^{2},$ (19)
$\widehat{R}_{0}(r_{0})$ $\coloneqq-\frac{2}{n}\sum_{i=1}^{n}r_{0}(Z_{i})+\frac{1}{|{\mathcal{G}}_{0}|}\sum_{i\in{\mathcal{G}}_{0}}r_{0}(Z_{i})^{2}.$ (20)
Appendix H Extensions
H.1 Nearest Neighbor Matching
Following this study, Kato (2025c) shows that nearest neighbor matching for ATE estimation can be interpreted as a special case of SQ-Riesz regression, that is, Riesz regression or LSIF. The key step is to express the ATE Riesz representer $\alpha^{\text{ATE}}_{0}(D,Z)$ in terms of density ratios with respect to the marginal covariate distribution, and to approximate these density ratios via nearest neighbor cells, following the density ratio interpretation in Lin et al. (2023).
NN matching ATE estimator.
Let
$J_{M}(i)\ \subset\ \{1,\ldots,n\}$
be the index set of the $M$ nearest neighbors of $X_{i}$ among the units with $D_{j}=1-D_{i}$. We define estimators $Y(d)$ as
$\widehat{Y}_{i}(0)$ $\coloneqq\begin{cases}Y_{i},&\text{if}\ \ D_{i}=0\\ \frac{1}{M}\sum_{j\in J_{M}(i)}Y_{j},&\text{if}\ \ D_{i}=1\end{cases},$
$\widehat{Y}_{i}(1)$ $\coloneqq\begin{cases}\frac{1}{M}\sum_{j\in J_{M}(i)}Y_{j},&\text{if}\ \ D_{i}=0\\ Y_{i},&\text{if}\ \ D_{i}=1\end{cases}.$
Then, the NN matching ATE estimator is given by
$\widehat{\theta}_{M}\coloneqq\frac{1}{n}\sum_{i=1}^{n}\left(\widehat{Y}_{i}(1)-\widehat{Y}_{i}(0)\right).$
Introduce the matched-times count (the number of times unit $i$ is used as a match by units in the opposite group) as
$K_{M}(i)\coloneqq\sum_{j=1,\ D_{j}=1-D_{i}}^{n}\mathbbm{1}\big[i\in J_{M}(j)\big].$
Then, $\widehat{\theta}_{M}$ can be written as follows:
$\widehat{\theta}_{M}=\frac{1}{n}\left(\sum_{i:D_{i}=1}\left(1+\frac{K_{M}(i)}{M}\right)Y_{i}-\sum_{i:D_{i}=0}\left(1+\frac{K_{M}(i)}{M}\right)Y_{i}\right)=\frac{1}{n}\sum_{i=1}^{n}(2D_{i}-1)\left(1+\frac{K_{M}(i)}{M}\right)Y_{i}.$
Nearest neighbor matching as density ratio estimation.
Lin et al. (2023) first shows that nearest neighbor matching can be interpreted as a method for density ratio estimation. Let $X,Z\in{\mathcal{X}}$ be independent whose pdfs are $p_{1}(x)$ and $p_{0}(z)$. We assume that $p_{1}(x),p_{0}(x)>0$ for all $x\in{\mathcal{X}}$. We observe i.i.d. samples $\{X_{i}\}_{i=1}^{N_{0}}$ and $\{Z_{j}\}_{j=1}^{N_{1}}$ and aim to estimate the density ratio
$r^{\dagger}_{0}(x)\coloneqq\frac{p_{1}(x)}{p_{0}(x)}.$
For $M\in\{1,\dots,N_{0}\}$ and $z\in\mathbb{R}^{d}$, let ${\mathcal{X}}_{(M)}(z)$ be the $M$-th nearest neighbor of $z$ in $\{X_{i}\}_{i=1}^{N_{0}}$ under a given metric $\|\cdot\|$. Define the catchment area of $x$ as
$A_{M}(x)\coloneqq\bigl\{z:\|x-z\|\leq\|{\mathcal{X}}_{(M)}(z)-z\|\bigr\},$
and the matched-times count as
$K_{M}(x)\coloneqq\sum_{j=1}^{N_{1}}\mathbbm{1}\big(Z_{j}\in A_{M}(x)\big).$
Lin et al. (2023) proposes the one-step estimator
$\widehat{r}^{\dagger}_{M}(x)=\frac{N_{0}}{N_{1}}\frac{K_{M}(x)}{M},$
which corresponds to nearest neighbor matching in ATE estimation.
Using this result, Lin et al. (2023) explains that nearest neighbor matching corresponds to the estimation of the density ratio $r_{0}$ defined above. They also show that their method is computationally efficient and rate-optimal for Lipschitz densities.
Nearest neighbor matching as LSIF
We next show that the density ratio estimator of Lin et al. (2023) is a variant of LSIF. Therefore, since we have already discussed that Riesz regression and LSIF are essentially the same, NN matching can also be interpreted as a special case of Riesz regression.
Let us consider the following density ratio model:
$r(1,z)=\phi(z)\beta,$
where $\phi(\cdot)$ is a basis function defined as
$\phi(z)\coloneqq\phi_{c}(z)=\mathbbm{1}\big[z\in A_{M}(c)\big],$
and recall that $A_{M}(c)\coloneqq\bigl\{z:\|c-z\|\leq\|{\mathcal{X}}_{(M)}(z)-z\|\bigr\}$.
For $z=c=X_{i}$, we define an estimator of the density ratio as
$\widehat{r}(1,c)=\phi(c)\widehat{\beta}$
with the estimated parameter defined as
$\widehat{\beta}\coloneqq\operatorname*{arg\,min}_{\beta\in{\mathbb{R}}}\left\{\frac{1}{2\sum^{n}_{i=1}\mathbbm{1}[D_{i}=1]}\Big(\phi_{c}(Z_{i})\beta\Big)^{2}-\frac{1}{n}\phi_{c}(Z_{i})\beta\right\},$
This estimation corresponds to LSIF with the kernel function.
This estimator is equivalent to
$\widehat{\beta}\coloneqq\operatorname*{arg\,min}_{\beta\in\mathbb{R}}\left\{\frac{1}{2}\beta^{\top}\widehat{H}\beta-\beta^{\top}\widehat{h}+\frac{\lambda}{2}\|\beta\|_{2}^{2}\right\}=\left(\widehat{H}+\lambda I\right)^{-1}\widehat{h}$,
where
$\widehat{H}$ $\coloneqq\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}\left[D_{i}=1\right]\phi_{c}(Z_{i})^{2}=\frac{M}{n},$
$\widehat{h}$ $\coloneqq\frac{1}{n}\sum_{i=1}^{n}\phi_{c}(Z_{i})=\frac{1}{n}\big(M+K_{M}(i)\big),$
$\phi_{c}(c)$ $=1.$
Here, we have
$\widehat{H}$ $=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}\left[D_{i}=1\right]\phi_{c}(Z_{i})=\frac{M}{n},$
$\widehat{h}$ $=\frac{1}{n}\sum_{i=1}^{n}\phi_{c}(Z_{i})=\frac{1}{n}\sum_{i=1}^{n}\big(\mathbbm{1}\left[D_{i}=1\right]\phi_{c}(Z_{i})+\mathbbm{1}\left[D_{i}=0\right]\phi_{c}(X_{i})\big)=\frac{1}{n}\big(M+K_{M}(i)\big),$
$\phi_{c}(c)$ $=1,$
where we recall that
$K_{M}(i)\coloneqq\sum_{j=1,\ D_{j}=1-D_{i}}^{n}\mathbbm{1}\left[i\in J_{M}(j)\right].$
Therefore, when $\lambda=0$, the estimator $\widehat{r}_{1}(c)$ is given by
$\widehat{r}(1,c)=\widehat{r}(1,Z_{i})=1+\frac{K_{M}(i)}{M}.$
Similarly, we can estimate $\widehat{r}(0,c)=\phi_{c}(c)\widehat{\beta}$ by solving an empirical version of the following problem:
$\widehat{\beta}\coloneqq\operatorname*{arg\,min}_{\beta\in{\mathbb{R}}}\left\{\frac{1}{2\sum^{n}_{i=1}\mathbbm{1}[D_{i}=0]}\Big(\phi_{c}(Z_{i})\beta\Big)^{2}-\frac{1}{n}\phi_{c}(Z_{i})\beta\right\},$
Then, the estimator is given by
$\widehat{r}(0,c)=\widehat{r}(0,Z_{i})=1+\frac{K_{M}(i)}{M}.$
Using these estimators, we construct the following inverse propensity score estimator for the ATE:
$\widehat{\theta}_{M}=\frac{1}{n}\left(\sum_{i:D_{i}=1}\left(1+\frac{K_{M}(i)}{M}\right)Y_{i}-\sum_{i:D_{i}=0}\left(1+\frac{K_{M}(i)}{M}\right)Y_{i}\right).$
This estimator is equivalent to an ATE estimator proposed in Lin et al. (2023), which is shown to be equal to the NN matching estimator of Abadie & Imbens (2006).
Thus, NN matching estimator is a special case of SQ-Riesz regression (LSIF) with a particular choice of a basis function.
H.2 Causal Tree / Causal Forest
Causal trees and causal forests estimate the conditional average treatment effect (CATE) by constructing a partition of the covariate space and estimating a local ATE within each cell, as in Wager & Athey (2018). We emphasize that this procedure implicitly constructs an estimator of the corresponding Riesz representer. In particular, once a partition is fixed, the leafwise CATE estimator can be rewritten as an inverse probability weighting type estimator, with weights that coincide with a leafwise Riesz representer estimator.
Leafwise CATE as a Riesz representer plug in.
Let $\Pi=\{\ell\}$ be a partition of the covariate space $\mathcal{Z}$} produced by a causal tree, and let $\ell(z)\in\Pi$ denote the leaf containing $z\in{\mathcal{Z}}$. For a leaf $\ell$, define $n_{\ell}\coloneqq\sum_{i=1}^{n}\mathbbm{1}[Z_{i}\in\ell]$, $n_{1,\ell}\coloneqq\sum_{i=1}^{n}\mathbbm{1}[D_{i}=1,Z_{i}\in\ell]$, and $n_{0,\ell}\coloneqq\sum_{i=1}^{n}\mathbbm{1}[D_{i}=0,Z_{i}\in\ell]$. The CATE estimator obtained by a causal tree is the leafwise difference in means
$\widehat{\theta}(z)\coloneqq\frac{1}{n_{1,\ell(z)}}\sum_{i:D_{i}=1,Z_{i}\in\ell(z)}Y_{i}-\frac{1}{n_{0,\ell(z)}}\sum_{i:D_{i}=0,Z_{i}\in\ell(z)}Y_{i}.$
This estimator admits the weighted representation
$\widehat{\theta}(z)=\frac{1}{n}\sum_{i=1}^{n}\widehat{\alpha}(D_{i},Z_{i};z)Y_{i},$
where
$\widehat{\alpha}(D,Z;z)\coloneqq\mathbbm{1}[Z\in\ell(z)]\left(\frac{D}{\widehat{\pi}_{1,\ell(z)}}-\frac{1-D}{\widehat{\pi}_{0,\ell(z)}}\right)\times\frac{1}{\widehat{p}_{\ell(z)}},\qquad\widehat{\pi}_{d,\ell}\coloneqq\frac{n_{d,\ell}}{n_{\ell}},\qquad\widehat{p}_{\ell}\coloneqq\frac{n_{\ell}}{n}.$
Hence $\widehat{\theta}(z)$ is an inverse probability weighting type estimator with a weight function $\widehat{\alpha}(\cdot,\cdot;z)$. This weight function is a plug-in estimator of the leafwise Riesz representer for the local ATE
$\theta(\ell)\coloneqq{\mathbb{E}}\big[Y(1)-Y(0)\mid Z\in\ell\big],$
[…]
Connection to SQ-Riesz regression and adaptive nearest neighbors.
The expression above shows that a causal tree is a histogram-type estimator of the Riesz representer, where the feature dictionary is given by leaf indicators $\{\mathbbm{1}[Z\in\ell]\}_{\ell\in\Pi}$. This is directly analogous to the nearest neighbor histogram model in the previous subsection, except that the partition is learned from the data rather than fixed a priori. From this viewpoint, the splitting criterion in a causal tree can be interpreted as choosing an adaptive partition that reduces the error of the induced leafwise Riesz representer approximation, and hence reduces the error of the resulting local CATE estimator.
A causal forest averages many such trees, built on subsamples and random feature choices, and therefore produces weights that average the leafwise Riesz representer estimators across trees. Equivalently, causal forests produce an adaptive nearest neighbor type representation for CATE, where the neighborhood structure is learned via the random partitions. This clarifies why causal trees and causal forests fit naturally into the same squared loss Bregman divergence, namely SQ-Riesz, perspective as nearest neighbor matching, with the main difference being that causal forests learn the partition adaptively to target CATE estimation accuracy.
H.3 AME Estimation by Score Matching
A subsequent work Kato (2025e) shows that, for derivative-type linear functionals, the Riesz representer can be estimated via score matching. This principle also underlies score-based diffusion models (Song & Ermon, 2020; Song et al., 2021). This viewpoint is useful for AME and APE estimation, and for mitigating overfitting in flexible Riesz representer models, because score matching objectives introduce smoothing through derivatives or noise perturbations.
Score matching identity for AME.
Recall the AME example in Section 2, where
$m^{\text{AME}}(W,\gamma)=\partial_{d}\gamma(D,Z),\qquad\alpha^{\text{AME}}_{0}(D,Z)=-\partial_{d}\log f_{0}(D,Z),$
with $f_{0}$ denoting the joint density of $X=(D,Z)$. Let $s_{0,d}(x)\coloneqq\partial_{d}\log f_{0}(x)$ be the $d$th component of the score. Consider a sufficiently smooth candidate function $\alpha(x)$ such that integration by parts is valid and boundary terms vanish. Then,
$\mathbb{E}\left[\partial_{d}\alpha(X)\right]=\int\partial_{d}\alpha(x)f_{0}(x){\mathrm{d}}x=-\int\alpha(x)\partial_{d}f_{0}(x){\mathrm{d}}x=-\mathbb{E}\left[\alpha(X)s_{0,d}(X)\right].$
Therefore, the squared loss Bregman objective for AME can be rewritten as
$\mathbb{E}\left[\alpha(X)^{2}-2\partial_{d}\alpha(X)\right]=\mathbb{E}\left[\alpha(X)^{2}+2\alpha(X)s_{0,d}(X)\right]=\mathbb{E}\left[\big(\alpha(X)+s_{0,d}(X)\big)^{2}\right]-\mathbb{E}\left[s_{0,d}(X)^{2}\right].$
The last term is constant in $\alpha$. Hence minimizing $\mathbb{E}\left[\alpha(X)^{2}-2\partial_{d}\alpha(X)\right]$ is equivalent to minimizing $\mathbb{E}\left[\big(\alpha(X)-\alpha^{\text{AME}}_{0}(X)\big)^{2}\right]$, and the population minimizer is $\alpha^{\text{AME}}_{0}=-s_{0,d}$. This is a coordinatewise form of the classical score matching principle and shows that, for derivative-type $m$, our squared loss Bregman risk coincides with an $L_{2}$ score matching risk for the Riesz representer.
Denoising score matching via diffusion.
In high dimensions, directly learning the score $x\mapsto\nabla_{x}\log f_{0}(x)$ can be unstable. Score-based diffusion models address this issue by learning scores of noise-perturbed distributions via denoising score matching (Song et al., 2021). Let $T$ be a noise index, continuous or discrete, and generate noisy covariates by
$X_{T}\coloneqq X+\sigma(T)Z,\qquad Z\sim\mathcal{N}(0,I),$
independent of $X\sim f_{0}$. Let $p_{T}$ denote the density of $X_{T}$. A time-dependent score model $s_{\theta}(\cdot,T)$ is trained by minimizing the denoising objective
${\mathbb{E}}\Big[\big\|\sigma(T)s_{\theta}(X_{T},T)+Z\big\|^{2}\Big],$
which is equivalent, up to an additive constant, to matching $s_{\theta}(\cdot,T)$ to the true score $\nabla_{x}\log p_{T}(x)$ under an $L_{2}$ risk. Once $s_{\theta}$ is trained, we can recover an estimator of the original score $\nabla_{x}\log f_{0}(x)$ by evaluating at small noise levels and then extracting the relevant component to estimate
$\alpha_{0}^{\text{AME}}(x)=-\partial_{d}\log f_{0}(x).$
Operationally, this replaces the derivative term $\partial_{d}\alpha(X)$ in the score matching objective with a denoising criterion that learns a smoothed score field. This smoothing can mitigate overfitting in high-capacity models and can be combined with flexible neural architectures through automatic differentiation.
H.4 Riesz Representer Estimation via Infinitesimal Classification
Next, following Kato (2025e), we introduce Riesz representer estimation via infinitesimal classification, which also reduces to score matching. This approach applies to a broader range of applications, not only to AME estimation.
Density ratio estimation via infinitesimal classification
We first review density ratio estimation via infinitesimal classification, proposed in Choi et al. (2022). Let $p_{0}(x)$ and $p_{1}(x)$ be two probability density functions such that $p_{0}(x)>0$ holds for all $x\in{\mathcal{X}}$. For $x\in{\mathcal{X}}$, the density ratio is defined as
$r_{0}(x)\coloneqq\frac{p_{0}(x)}{p_{1}(x)}.$
We aim to estimate $r_{0}$.
We define a continuum of bridge densities $\{p_{t}\}_{t\in[0,1]}$ through a simple sampling procedure. Let $p_{t}(x)$ be the probability density function of the random variable
$X_{t}=\beta^{(1)}_{t}X_{0}+\beta^{(2)}_{t}X_{1},$
where $\beta^{(1)}_{\cdot},\beta^{(2)}_{\cdot}\colon[0,1]\to[0,1]$ are $C^{2}$ and monotonic, and satisfy the boundary conditions $\beta^{(1)}_{0}=1$, $\beta^{(2)}_{0}=0$, $\beta^{(1)}_{1}=0$, and $\beta^{(2)}_{1}=1$. Using $\frac{p_{(t-1)/T}(x)}{p_{t/T}(x)}$ as an intermediate density ratio, we decompose the density ratio into a product of density ratios as
$r_{0}(x)=\prod^{T}_{t=1}\frac{p_{(t-1)/T}(x)}{p_{t/T}(x)}.$
We can choose $\beta^{(1)}_{t}$ and $\beta^{(2)}_{t}$ so that the density ratio can be trained stably. For example, DRE-$\infty$ proposes using $\beta^{(1)}_{t}=1-t$ and $\beta^{(2)}_{t}=t$ in some applications.
In practice, when optimizing objectives that integrate over $t$, we sample $t$ jointly with $(X_{0},X_{1})$. Specifically, for each stochastic gradient step we draw a mini batch $\{(X_{0,i},X_{1,i})\}_{i=1}^{B}$ with $X_{0,i}\sim p_{0}$ and $X_{1,i}\sim p_{1}$ independently, and we draw times $\{t_{i}\}_{i=1}^{B}$ i.i.d. from a reference density $q(t)$ on $[0,1]$. We then form $X_{t_{i},i}=\beta^{(1)}(t_{i})X_{0,i}+\beta^{(2)}(t_{i})X_{1,i}$ and approximate time integrals using importance weights. For example, an integral term of the form $\int_{0}^{1}{\mathbb{E}}_{X_{t}\sim p_{t}}[h(X_{t},t)]{\mathrm{d}}t$ is estimated by
$\int_{0}^{1}{\mathbb{E}}_{X_{t}\sim p_{t}}[h(X_{t},t)]{\mathrm{d}}t\approx\frac{1}{B}\sum_{i=1}^{B}\frac{h(X_{t_{i},i},t_{i})}{q(t_{i})}.$
Endpoint expectations, such as ${\mathbb{E}}_{X_{0}\sim p_{0}}[\cdot]$ and ${\mathbb{E}}_{X_{1}\sim p_{1}}[\cdot]$, are approximated by sample averages over $\{X_{0,i}\}$ and $\{X_{1,i}\}$, respectively. All derivatives with respect to $t$ that appear in the objective, such as $\partial_{t}(\lambda(t)s_{\beta}(X_{t},t))$, can be computed by automatic differentiation through the explicit dependence of $X_{t}$ on $t$ via $\beta^{(1)}(t)$ and $\beta^{(2)}(t)$.
By taking the logarithm, we have
$\log\big(r_{0}(x)\big)=\sum^{T}_{t=1}\log\frac{p_{(t-1)/T}(x)}{p_{t/T}(x)}.$
Then, as $T\to\infty$, the following holds (Choi et al., 2022; Chen et al., 2025):
$\log r_{0}(x)=\log\left(\frac{p_{0}(x)}{p_{1}(x)}\right)=\sum^{T}_{t=1}\log\left(\frac{p_{(t-1)/T}(x)}{p_{t/T}(x)}\right)=\int^{0}_{1}\partial_{t}\log p_{t}(x){\mathrm{d}}t.$
Let $s^{\text{time}}_{\beta}(x,t)$ be a time score model that approximates the time score $\partial_{t}\log p_{t}(x)$. We train $s^{\text{time}}_{\beta}(x,t)$ by minimizing the following time score matching loss (Choi et al., 2022; Chen et al., 2025):
${\mathcal{R}}^{\dagger}\left(s^{\text{time}}_{\beta}\right)\coloneqq\int^{1}_{0}{\mathbb{E}}_{X_{t}\sim p_{t}(x)}\left[\lambda(t)\Big(\partial_{t}\log p_{t}(X_{t})-s^{\text{time}}_{\beta}(X_{t},t)\Big)^{2}\right]{\mathrm{d}}t,$
where $\lambda\colon[0,1]\to{\mathbb{R}}_{+}$ is a positive weighting function. Although $\log p_{t}(x)$ is unknown in practice, the following alternative objective has been proposed, which is equivalent to ${\mathcal{R}}^{\dagger}\left(s^{\text{time}}_{\beta}\right)$ up to a constant term that is irrelevant for optimization:
${\mathcal{R}}\left(s^{\text{time}}_{\beta}\right)$ $\coloneqq{\mathbb{E}}_{X_{0}\sim p_{0}(x)}\Big[\lambda(0)s^{\text{time}}_{\beta}(X_{0},0)\Big]-{\mathbb{E}}_{X_{1}\sim p_{1}(x)}\big[\lambda(1)s^{\text{time}}_{\beta}(X_{1},1)\big]$
$\ \ \ +\int^{1}_{0}{\mathbb{E}}_{X_{t}\sim p_{t}(x)}\left[\partial_{t}\left(\lambda(t)s^{\text{time}}_{\beta}(X_{t},t)\right)+\frac{1}{2}\lambda(t)s^{\text{time}}_{\beta}(X_{t},t)^{2}\right]{\mathrm{d}}t,$
To generate a sample from $p_{t}$, we proceed as follows. First, draw two independent endpoint samples
$X_{0}\sim p_{0},\qquad X_{1}\sim p_{1},$
independently across draws and independent of each other. Second, for a given time $t\in[0,1]$, construct the bridge sample by the deterministic map
$X_{t}\coloneqq\beta^{(1)}(t)X_{0}+\beta^{(2)}(t)X_{1}.$
We define $p_{t}$ as the probability law of $X_{t}$ induced by this procedure, that is, $p_{t}$ is the pushforward of the product measure $p_{0}\otimes p_{1}$ through the map $(x_{0},x_{1})\mapsto\beta^{(1)}(t)x_{0}+\beta^{(2)}(t)x_{1}$. With this definition, expectations under $p_{t}$ can be evaluated by Monte Carlo as
${\mathbb{E}}_{X_{t}\sim p_{t}}\bigl[f(X_{t},t)\bigr]={\mathbb{E}}\bigl[f(\beta^{(1)}(t)X_{0}+\beta^{(2)}(t)X_{1},t)\bigr],$
where the outer expectation is taken over $(X_{0},X_{1})\sim p_{0}\otimes p_{1}$.
Riesz representer estimation via infinitesimal classification
Kato (2025e) extends density ratio estimation via infinitesimal classification to Riesz representer estimation. In this subsection, we introduce an example of the method for APE estimation. For implementations in other applications, see Kato (2025e).
In APE estimation, the Riesz representer is given by
$\alpha^{\text{APE}}(X)\coloneqq\frac{p_{1}(X)-p_{-1}(X)}{p_{0}(X)}.$
By using intermediate density ratios, we have
$\frac{p_{1}(x)}{p_{0}(x)}$ $=\prod^{T}_{t=1}\frac{p_{t/T}(x)}{p_{(t-1)/T}(x)},$
$\frac{p_{-1}(x)}{p_{0}(x)}$ $=\prod^{T}_{t=1}\frac{p_{-t/T}(x)}{p_{-(t-1)/T}(x)}.$
Then, we can approximate the density ratio as
$\log\frac{p_{1}(x)}{p_{0}(x)}$ $=\sum^{T}_{t=1}\log\frac{p_{t/T}(x)}{p_{(t-1)/T}(x)}\to\int^{1}_{0}\partial_{t}\log p_{t}(x){\mathrm{d}}t\quad(T\to\infty),$
$\log\frac{p_{-1}(x)}{p_{0}(x)}$ $=\sum^{T}_{t=1}\log\frac{p_{-t/T}(x)}{p_{-(t-1)/T}(x)}\to\int^{-1}_{0}\partial_{t}\log p_{t}(x){\mathrm{d}}t\quad(T\to\infty).$
We define a random variable $X_{t}$ as
$X_{t}\coloneqq\begin{cases}\beta^{(1)}_{t}X_{1}+\beta^{(2)}_{t}X_{0}&\text{if}\ \ \ t\geq 0\\ \beta^{(1)}_{t}X_{-1}+\beta^{(2)}_{t}X_{0}&\text{if}\ \ \ t<0\end{cases},$
where $\beta^{(1)}_{t},\beta^{(2)}_{t}\colon[-1,1]\to[0,1]$ are of class $C^{2}$ and monotonic, with $\beta^{(1)}_{t}$ increasing and $\beta^{(2)}_{t}$ decreasing for $t\geq 0$, $\beta^{(1)}_{t}$ decreasing and $\beta^{(2)}_{t}$ increasing for $t<0$, and satisfying the boundary conditions: $\beta^{(1)}_{0}=0$, $\beta^{(2)}_{0}=1$, $\beta^{(1)}_{-1}=1$, $\beta^{(2)}_{-1}=0$, $\beta^{(1)}_{1}=1$, and $\beta^{(2)}_{1}=0$.
Let $p_{t}(x)$ be the probability density function. Let $s^{\text{time}}_{\beta}(x,t)$ be a time score model that approximates the time score $\partial_{t}\log p_{t}(x)$. We train the score model by minimizing
${\mathcal{R}}^{\text{APE}\dagger}\left(s^{\text{time}}_{\beta}\right)\coloneqq\int^{1}_{-1}{\mathbb{E}}_{X_{t}\sim p_{t}(x)}\left[\lambda(t)\Big(\partial_{t}\log p_{t}(X_{t})-s^{\text{time}}_{\beta}(X_{t},t)\Big)^{2}\right]{\mathrm{d}}t,$
where $\lambda\colon[-1,1]\to{\mathbb{R}}_{+}$ is a positive weighting function. Since $\partial_{t}\log p_{t}(x)$ is unknown, we minimize the following risk:
${\mathcal{R}}^{\text{APE}}\left(s^{\text{time}}_{\beta}\right)$ $\coloneqq{\mathbb{E}}_{X_{-1}\sim p_{-1}(x)}\Big[\lambda(-1)s^{\text{time}}_{\beta}(X_{-1},-1)\Big]-{\mathbb{E}}_{X_{1}\sim p_{1}(x)}\big[\lambda(1)s^{\text{time}}_{\beta}(X_{1},1)\big]$
$\ \ \ +\int^{1}_{-1}{\mathbb{E}}_{X_{t}\sim p_{t}(x)}\left[\partial_{t}\left(\lambda(t)s^{\text{time}}_{\beta}(X_{t},t)\right)+\frac{1}{2}\lambda(t)s^{\text{time}}_{\beta}(X_{t},t)^{2}\right]{\mathrm{d}}t.$